### v2

# Using iterated local alignment to aggregate trajectory data into a traffic flow map

Abstract
Desire line maps are widely deployed for traffic flow analysis by virtue of their ease of interpretation and computation. They can be considered to be simplified traffic flow maps, whereas the computational challenges in aggregating small scale traffic flows prevent the wider dissemination of high resolution flow maps. Vehicle trajectories are a promising data source to solve this challenging problem. The solution begins with the alignment (or map matching) of the trajectories to the road network. However even the state-of-the-art map matching implementations produce sub-optimal results with small misalignments. While these misalignments are negligible for large scale flow aggregation in desire line maps, they pose substantial obstacles for small scale flow aggregation in high resolution maps. To remove these remaining misalignments, we introduce innovative local alignment algorithms, where we infer road segments to serve as local reference segments, and proceed to align nearby road segments to them. With each local alignment iteration, the misalignments of the trajectories with each other and with the road network are reduced, and so converge closer to a minimal flow map. By analysing a set of empirical trajectories collected in Hannover, Germany, we confirm that our minimal flow map has high levels of spatial resolution, accuracy and coverage.
Keywords: Desire lines, floating car data FCD, GPS, map matching, route finding
1 Introduction
One of the fundamental quantities in transport planning is a traffic flow map, i.e. a map of the traffic flow levels on the road segments in a road network (Ortúzar and Willumsen, 2011). Whilst traffic flow maps are a rich source of information about vehicle mobility patterns, they are costly in terms of time and resources to compute for any reasonably sized road network. To alleviate this cost burden, most approaches restrict the spatial coverage and resolution of the flow map. One of the most common is to place sensors at fixed locations in the road network, whose results are then visualised as a traffic count map. Another are trip intent/recall questionnaires to inform large scale properties such as trajectory origins and destinations, which can be visualised as a desire line flow map/spider diagram (Tobler, 1987). Both of these are simplified flow maps, since the detailed mobility patterns outside of the sensor locations or origin/destination pairs are not known. These unknown patterns can be inferred from other data sources, such as route assignment models (Evans, 1976). While these model-assigned routes are highly detailed, the trade-off is that they are not guaranteed to correspond closely to empirical mobility patterns.
Thus an ideal data source combines the empirical information of road sensor counts or questionnaires, with the small scale details of model-assigned trajectories. This gap in the market can be filled by trajectory data. Due to the prevalence of geolocation-enabled devices, such as vehicle navigation guides and mobile telephones, trajectory data can be acquired with low marginal cost whilst at the same time offering extensive spatial coverage and resolution of empirical mobility patterns (Herrera et al., 2010, Andrienko and Andrienko, 2013). Throughout, we employ the term ‘trajectory data’ rather than Floating Car Data (FCD) or Floating Mobile Data (FMD), as our analysis is not restricted to trajectory data from cars or from mobile phones, so these terms are equivalent for our purposes. An example is the 1183 trajectories are collected from a GPS-enabled mobile phone, from December 2017 to March 2019 by a single driver in Hannover, Germany, with an overall average sampling rate about 1 point per second (Zourlidou et al., 2022). They are plotted as the green circles in Figure 1.
Refer to caption(a)
Refer to caption(b)
Refer to caption(c)
Figure 1: Trajectories in Hannover, Germany. (a) City level. (b) Neighbourhood level, zoom of black rectangle. (c) Small neighbourhood level, zoom of blue rectangle. Trajectory ID = 7 (orange circles), 315 (purple diamonds), others (green circles).
In Figure 1(a), at the city level, the points (green circles) appear to be aligned to the road network. If we zoom in on the small black rectangle in the central region, then in Figure 1(b) at the neighbourhood level, we observe that the points deviate from the road network. This deviation is clearer in the closer zoom in Figure 1(c). Moreover, if we focus on the orange circles (Trajectory ID = 7) and purple diamonds (ID = 315), then we observe that the vehicle location between the recorded points is unknown. These maps illustrate the errors in trajectories. These are broadly classified as ‘measurement errors’ where the recorded coordinates are not the true locations, and ‘sampling errors’ where the information about the trajectory is lost in between recorded coordinates (Ortúzar and Willumsen, 2011, Saki and Hagen, 2022).
Our goal is to produce a traffic flow map from these noisy trajectories, which can be utilised at any scale, from the city/regional level to the individual road segment level. This requires us to minimise the errors in trajectories. Our approach is composed of two stages. The first stage is to align the trajectories to the road network, which is known as map matching (Quddus et al., 2007, Chao et al., 2020). It produces a route, which is a connected sequence of road segments in the road network, that is consistent with the trajectory. Our contribution is an improvement to standard map matching by adding post hoc route finding. In common with many open source transport planning tools, we employ APIs based on the OpenStreetMap (OSM) network. While we are able to improvement the overall alignment to the road network, these map matched routes inherit incompressible misalignments, ranging from several centimetres to several metres, from the OSM road network. These small misalignments prevent the accurate aggregation of traffic flows at this scale.
The second stage is to resolve these misalignments, and our contribution here is the proposed local alignment of map matched routes. In contrast to global alignment approaches, we locally infer which road segments should serve as a local reference, and then proceed to align other nearby road segments to it. To accomplish this, we introduce several novel algorithms which employ a mix of advanced statistical and geospatial methods. These include ‘node snapping’ where the nearby boundary points of road segments are combined via statistical clustering, and ‘line blending’ where road segments, which are near to each other but do not share overlapping road sub-segments, are aligned to maximise their overlapping sub-segments. Inputting these locally aligned routes into a flow aggregation API leads to a more accurate flow map. Iterating these local alignments in turn leads to a minimal flow map.
The outline of the paper is as follows. In Section 2, we describe the computation of map matched routes using off-the-shelf map matching and route finding APIs. In Section 3, we describe the local alignment of the map matched routes and their aggregation into a flow map. In Section 4, we demonstrate that the locally aligned flow map from the Hannover trajectories is well-aligned to the OSM road network, and has a high level of accuracy and spatial coverage of estimated traffic flows compared to reference traffic flows. We then discuss some software implementation issues, and some concluding remarks.
2 Alignment of trajectories with map matching and route finding
We introduce some mathematical notation to state precisely the problem of map matching. We represent the road network by a graph $\mathcal{N}=\mathcal{N}(\mathcal{E},\mathcal{V})$, where the $n_{\mathcal{E}}$ edges =$\mathcal{E}=\{{\bm{e}}_{1},\dots,{\bm{e}}_{n_{\mathcal{E}}}\}$ of this graph are the road segments and the $n_{\mathcal{V}}$ nodes/vertices $\mathcal{V}=\{{\bm{v}}_{1},\dots,{\bm{v}}_{n_{\mathcal{V}}}\}$ indicate that two (or more) different road segments are accessible to/from each other at this node point. These nodes ${\bm{v}}_{i}$ are single points. Each road segment is composed of a sequence of connected piecewise linear segments, so ${\bm{e}}_{i}=\{{\bm{e}}_{i,1},\dots,{\bm{e}}_{i,n_{i}}\}$ is an ordered sequence of $n_{i}$ points ${\bm{e}}_{i,j},j=1,\dots,n_{i}$. We refer to this as a ‘linestring’ geometry, following the Open Geospatial Consortium terminology (OGC, 2010). We set $\mathcal{N}$ to be the OSM network which is a freely available road network with global coverage.
We denote a single trajectory $G=\{{\bm{g}}_{1},\dots,{\bm{g}}_{n_{G}}\}$ as a temporally ordered sequence of $n_{G}$ points ${\bm{g}}_{i},i=1,\dots,n_{G}$. This is known as a ‘multipoint’ geometry (OGC, 2010). Whilst some authors require that each point ${\bm{g}}_{i}$ be accompanied by their timestamp $t_{i}$ to be considered a trajectory, this is not strictly required since ${\bm{g}}_{i}$ are ordered according to the timestamps, even if the timestamps themselves are not recorded in the trajectory. Due to measurement error, the points of a trajectory $G$ are not necessarily coincident with the road network $\mathcal{N}$, and due to sampling error, there is no information about the vehicle in between the points of $G$.
We represent the output of a map matching algorithm $M$ from an empirical trajectory $G$ as $M(G)=\{{\bm{m}}_{1},\dots,{\bm{m}}_{n_{M}}\}$ which is an ordered, connected sequence of $n_{M}$ edges. The goal is that the map matched route follows closely the road network $\mathcal{N}(\mathcal{V},\mathcal{E})$. Whilst it is straightforward to ensure that all boundary points of the segments $\{{\bm{m}}_{1},\dots,{\bm{m}}_{n_{M}}\}$ coincide with the nodes $\mathcal{V}$ of the road network graph $\mathcal{N}$, it is more challenging to ensure that the segments $\{{\bm{m}}_{1},\dots,{\bm{m}}_{n_{M}}\}$ themselves coincide with the edges $\mathcal{E}$. In comparison to an empirical trajectory $G$, all boundary points of the edges of a map matched route $M(G)$ are aligned to the road network (reduced measurement error) and the vehicle position is estimated by the linestring connecting the boundary points of the road segment (reduced sampling error).
There is a large body of research on this difficult problem of map matching. We focus on the popular class of Hidden Markov Models (HMM) map matching algorithms. HMM methods iteratively build the map matched route by selecting the most likely next segment to connect to the current route using a probabilistic model. According to a review of map matching algorithms (Chao et al., 2020), HMM is a state-transition method. The other three classes are similarity, candidate-evolving and scoring methods. Further details of alternative map matching algorithms are found in Quddus et al. (2007), Chao et al. (2020). We leave this discussion here since the improvements offered by our proposed methods are valid for any map matching algorithm, and concentrate on HMM map matching due to its accuracy and computational efficiency. Even if we restrict ourselves to HMM algorithms on the OSM road network, there are many off-the-shelf map matching APIs available. We focus on the Valhalla routing engine (https://valhalla.github.io/valhalla), which includes its highly recommended map matching API (Saki and Hagen, 2022).
Figure 2 displays the $n=1147$ map matched routes by the Valhalla map matching API. We discard 36 trajectories (3.04%) from our original data set of 1183 trajectories. In Figure 2(a), the map matched routes (blue lines) overall are well-aligned to the road network. For the orange map matched route, all its segments are aligned to the road network, whereas the purple route appears to be displaced by several metres from the road centreline. The measurement and sampling errors of the map matched routes $M(G)$ are reduced in comparison those for the empirical trajectories $G$, though these errors remain sizeable at the road segment level in Figure 2(b).
Refer to caption(a) Refer to caption(b)
Figure 2: Map matched routes, using only Valhalla map matching API. (a) Neighbourhood level. (b) Small neighbourhood level, zoom of black rectangle. Trajectory ID = 7 (orange), 315 (purple), others (blue).
Our first contribution is to better align the edges of the map matched routes $M(G)$ to the road network edges $\mathcal{E}$. We propose a post hoc adjustment of the map matching output by an additional call to a route finding API, as outlined in Algorithm 1. The inputs of ST_ROUTE are the empirical trajectory $G$, the map matching API $M$, the route finding API $R$, and the number of waypoints ${\bm{n}}_{W}$ for the route finding. Since we do not have an a priori single optimal value for the number of waypoints, we consider a range of $w$ values ${\bm{n}}_{W}=(n_{W,1},\dots,n_{W,w})$. In Step 1, we compute the initial map matched route $M(G)$ from the empirical trajectory $G$ by calling the map matching API $M$. This initial map matched route $M(G)=\{{\bm{m}}_{1},\dots,{\bm{m}}_{n_{M}}\}$ is a linestring with $n_{M}$ edges, with $n_{M}+1$ points. In Steps 2–5, we loop over the $w$ number of waypoints in ${\bm{n}}_{W}$. In Steps 3–4, we take a sample of $n_{W,i}$ waypoints, where the first waypoint ${\bm{w}}_{1}$ is the start point of ${\bm{m}}_{1}$ and the $n_{W,i}$th way point is the end point of ${\bm{m}}_{n_{M}}$, and the intermediate waypoints ${\bm{w}}_{2},\dots,{\bm{w}}_{n_{W,i}-1}$ are sampled from the start points of the edges ${\bm{m}}_{2},\dots,{\bm{m}}_{n_{M}}$. In Step 5, we call the route finding API $R$ with the waypoints ${\bm{w}}_{1},\dots,{\bm{w}}_{n_{W,i}}$. The result is the map matched route $M_{i}^{*}(G)=R({\bm{w}}_{1},\dots,{\bm{w}}_{n_{W,i}})=\{{\bm{m}}_{1}^{*},\dots,{\bm{m}}_{n_{M^{*}}}\}$, which is a route of $n_{M^{*}}$ connected edges. In Step 6, we select the route with the smallest dynamic time warping (DTW) normalised distance between the routes $M_{i}^{*}(G)$ and the empirical trajectory $G$, for $i=1,\dots w$. The DTW distance is based on the lengths of all the distortions of $M_{i}^{*}(G)$ to achieve a maximal alignment between $M_{i}^{*}(G)$ and $G$ (Sakoe and Chiba, 1978, Giorgino, 2009).
Algorithm 1 ST_ROUTE – Map matched route
1:Input: $G$ trajectory, $M$ map matching API, $R$ route finding API, ${\bm{n}}_{W}$waypoints
2:Output: $M^{*}(G)$ map matched route
3:Compute initial map matched route $M(G)$ from empirical trajectory $G$
4:for $i:=1$ to $w$ do
5:     Set $n_{W}:=n_{W,i}$
6:     Sample $n_{W}$ waypoints ${\bm{w}}_{1},\dots,{\bm{w}}_{n_{W}}$ from $M(G)$, with ${\bm{w}}_{1}$ := $\operatorname{Start}(M)$, ${\bm{w}}_{n_{W}}:=\operatorname{End}(M)$
7:     Compute map matched route $M^{*}_{i}(G):=R({\bm{w}}_{1},\dots,{\bm{w}}_{n_{W}})$
8:Select minimal route $M^{*}(G):=\operatorname{argmin}_{i\in\{1\dots w\}}\,\operatorname{DTW}(M^{*}_{i}(G),G)$
We set $R$ to be the Valhalla Odin turn-by-turn route finding API to be consistent with our choice of $M$ as the Valhalla Meili map matching API. We set the number of waypoints as ${\bm{n}}_{W}=3,13,23,33,43,63,83$. Figure 3 displays the results $M^{*}(G_{1}),\dots,M^{*}(G_{1147})$ from ST_ROUTE. For the trajectory ID = 7 (orange) $n_{W}=83$, and ID = 315 (purple) $n_{W}=23$ give the minimal DTW route. The overall impression of the map matched routes $M^{*}$ in Figure 3(a) is that the misalignment has been reduced, especially in the purple line since it now aligns more closely to the road centreline. In Figure 3(b), at the level of road segments, whilst the map matched routes $M^{*}$ tend to be contained inside the road segments, they are not exactly coincident with each other. This is in part because the road network graph $\mathcal{N}$ itself contains small misalignments, and so they are propagated into any map matching or route finding algorithm based on it. These small alignments lead to a lack of overlapping sub-segments, which in turn lead to inaccurate flow aggregation.
Refer to caption(a) Refer to caption(b)
Figure 3: Map matched routes, using Valhalla map matching and route finding APIs. (a) Neighbourhood level. (b) Small neighbourhood level, zoom of black rectangle. Trajectory ID = 7 (orange), 315 (purple), others (blue).
3 Local alignment of road segments for flow map aggregation
Our goal is to resolve the remaining misalignments from the map matching/route finding in the previous section, so we are able to aggregate accurately traffic flows on road segments. We aim to achieve this by local, internal alignment between the map matched routes. By internal alignment, we mean that we align the routes with each other, rather than to the external road network graph. Since an external reference road network is not required as an explicit input, our proposal can be deployed in more cases, e.g. when the quality of the road network graph is insufficient, or when the alignment to the road network graph is computationally intensive. By local alignment, we mean that we focus on aligning sub-segments of the routes, rather than complete routes.
[…]
We begin by illustrating the difference in flow aggregation between the $M$ and $M^{*}$ map matched routes. For the routes $M$ from Figure 2(b) and $M^{*}$ routes from Figure 3(b), the flow maps are given below in Figure 4(a–b) respectively. In these maps, the colour (purple to orange) and width of the road segments is proportional to the traffic flow. We observe that there are fewer, wider linestring segments in Figure 4(b) than in Figure 4(a).
Refer to caption(a) Refer to caption(b)
Figure 4: Flow maps for map matched routes. (a) Map matched routes with Valhalla map matching only. (b) Map matched routes with Valhalla map matching and route finding. Colour (purple to orange) and width of road segments is proportional to traffic flow.
The flow aggregation in Figure 4 was carried out using the overline function in the R package stplanr, which we refer to as ST_OVERLINE_PLANR (Lovelace and Ellison, 2018). Starting with the map matched routes =$\mathcal{M}^{*}=\{M^{*}({G_{1}}),\dots,M^{*}(G_{n})\}$, then the flow map is $\mathcal{F}={\tt ST\_OVERLINE\_PLANR}(\mathcal{M}^{*})$. This flow aggregation involves the search for all road segments from the routes which exactly equal to each other. These exactly equal road segments are reduced to a single common segment, and the associated traffic flow is the number of exactly equal segments. Since it relies on exactly equal road segments, then small misalignments are sufficient to make the flow aggregations inaccurate.
Our goal of producing a minimal flow map relies on resolving the crucial problem of how to aggregate similar, but not exactly overlapping, road segments. Many solutions have been offered, such as edge bundling (Zhou et al., 2013) and rasterisation (Wood et al., 2010, Morgan and Lovelace, 2021). Edge bundling consists of clustering trajectory linestrings and replacing all cluster members with a single representative linestring. These (and subsequent) authors conclude that it performs poorly when applied to noisy trajectories at the road segment level, and remains mostly suited to coarser aggregations, such as desire lines. Rasterisation relies on converting the vectorial flow map into a raster matrix, and aggregating the flows within the same raster pixel neighbourhood. Whilst this is indeed able to improve flow aggregations, it depends highly on the raster pixel neighbourhood size, and the rasterisation of the vectorial flow map leads to a loss of spatial resolution. We propose an alternative aggregation which does not lose resolution. Due to the complexity of this aggregation, it is divided into several algorithms, so that after the application of each algorithm, we progress further towards a minimal flow map.
We develop our novel algorithms within the R statistical analysis environment, to take advantage of its integrated access to advanced statistical and geospatial analysis methods. Whilst R is not a bona fide GIS (Geographical Informations System), its geospatial functionalities conform to the OGC standards (OGC, 2010) via the package sf (Pebesma, 2018), and is a viable option for research in transport geospatial data analysis (Necula, 2015, Lovelace et al., 2019).
3.1 Node snapping with hierarchical clustering
The proposed algorithm is ST_SNAPNODE, where the boundary points of the traffic flow linestrings are snapped to each other. Since the former are also nodes of the flow map, this gives the name to the algorithm. We focus on snapping these nodes since the linestring misalignments are in part caused by the existence of nodes which are close to each other but not exactly equal.
Since we are searching for points which are close to together, then this is well-suited to statistical clustering. There are many statistical clustering algorithms available, and we focus on hierarchical clustering (Gordon, 1999). A naive implementation where we consider all boundary points of all flow linestrings in a 1-pass complete linkage clustering is computationally intensive for any reasonable number of routes (Müllner, 2013). To resolve this computational bottleneck, we approximate the 1-pass complete linkage clustering by a nested 2-pass clustering. We begin with an efficient single linkage clustering of the boundary points of all linestrings in the R package fastcluster (Müllner, 2013). Since single linkage can result in chain-like clusters, we compute a subsequent complete linkage clustering to break these potential chains. In this nested approach, the complete linkage distance matrix is calculated only within each single linkage cluster, and so we are less likely to reach computational limits.
Algorithm 2 is a description of ST_SNAPNODE. The inputs are the $n_{\mathcal{F}}$ traffic flow linestrings =$\mathcal{F}=\{{\bm{f}}_{1},\dots,{\bm{f}}_{n_{\mathcal{F}}}\}$ and the snap tolerance $\varepsilon_{S}$. In Step 1, we extract the boundary points of all flow linestrings. In Steps 2–3, we compute a single linkage clustering on all boundary points, and cut the dendrogram at height $\varepsilon_{S}$, resulting in $C^{\prime}$ clusters. In Steps 4–7, within each of these single linkage clusters, we compute a complete linkage clustering, and cut the dendrogram at height $\varepsilon_{S}$. This divides each single linkage cluster into $C^{\prime\prime}$ clusters, where all cluster members are at most $\varepsilon_{S}$ distance from each other. In Steps 8–11, we compute the weighted spatial centroid in each of the $C^{\prime\prime}$ complete linkage clusters, weighted by the corresponding traffic flow values. In Step 12, we renumber the cluster labels from the above nested clusterings to approximate a 1-pass complete linkage clustering into $C^{*}$ clusters. In Steps 13–16, within each of these $C^{*}$ clusters, for all points of the corresponding flow linestrings which are closer than $\varepsilon_{S}$ distance to the centroid, we snap them to the centroid. In Steps 17–18, we collate and sort all the $n_{\mathcal{F}^{*}}$ snapped linestrings =$\mathcal{F}^{*}=\{{\bm{f}}_{1}^{*},\dots,{\bm{f}}_{n_{\mathcal{F}^{*}}}^{*}\}$. In comparison to the unsnapped linestrings $\mathcal{F}$, there are fewer snapped linestrings $\mathcal{F}^{*}$ which have higher flow values, are more interconnected, and share more exactly overlapping sub-segments.
Algorithm 2 ST_SNAPNODE – Snap node clustering of linestrings
1:Input: $\mathcal{F}$ flow linestrings, $\varepsilon_{S}$ snap tolerance
2:Output: $\mathcal{F}^{*}$ snap node clustered flow linestrings
3:Extract boundary points $B(\mathcal{F}):=\{\operatorname{Start}({\bm{e}}_{1}),\operatorname{End}({\bm{e}}_{1}),\dots,\operatorname{Start}({\bm{e}}_{n_{\mathcal{F}}}),\operatorname{End}({\bm{e}}_{n_{\mathcal{F}}})\}$
4:Compute hierarchical clustering with single linkage on $B(\mathcal{F})$
5:Cut single linkage dendrogram at height $\varepsilon_{S}$ to compute $C^{\prime}$ cluster labels
6:for $i:=1$ to $C^{\prime}$ do
7:     Extract $i$th cluster of boundary points $B_{i}^{\prime}:=\{{\bm{b}}_{1}^{\prime},\dots,{\bm{b}}_{n_{B^{\prime}}}^{\prime}\}$
8:     Compute hierarchical clustering with complete linkage on $B_{i}^{\prime}$
9:     Cut complete linkage dendrogram at height $\varepsilon_{S}$ to compute $C^{\prime\prime}$ cluster labels
10:     for $j:=1$ to $C^{\prime\prime}$ do
11:         Extract $j$th cluster of boundary points $B_{j}^{\prime\prime}:=\{{\bm{b}}_{1}^{\prime\prime},\dots,{\bm{b}}_{n_{B^{\prime\prime}}}^{\prime\prime}\}$
12:         Extract corresponding flows $\{{\bm{f}}_{1}^{\prime\prime},\dots,{\bm{f}}_{n_{B^{\prime\prime}}}^{\prime\prime}\}$ from $\mathcal{F}$
13:         Compute ${\bm{b}}_{j}^{*}$ := weighted centroid of $\{{\bm{f}}_{1}^{\prime\prime},\dots,{\bm{f}}_{n_{B^{\prime\prime}}}^{\prime\prime}\}$, weights := $f_{1}^{\prime\prime},\dots,f_{n_{B^{\prime\prime}}}^{\prime\prime}$
14:Renumber collated complete linkage cluster labels to unique $C^{*}$ labels
15:for $i:=1$ to $C^{*}$ do
16:     Extract corresponding flow linestrings $\mathcal{F}_{i}^{*}:=\{{\bm{f}}_{1}^{*},\dots,{\bm{f}}_{n_{B^{*}}}^{*}\}$ from $\mathcal{F}$
17:     Snap points of $\{{\bm{f}}_{1}^{*},\dots,{\bm{f}}_{n_{B^{*}}}^{*}\}$ within $\varepsilon_{S}$ distance of cluster centroid ${\bm{b}}^{*}_{i}$ to ${\bm{b}}^{*}_{i}$
18:     $\mathcal{F}_{i}^{*}$ := rejoin snapped and unsnapped points into linestrings
19:Collate snapped linestrings $\{\mathcal{F}_{1}^{*},\dots\mathcal{F}_{C^{\prime*}}^{*}\}$ into :=$\mathcal{F}^{*}:=\{{\bm{f}}_{1}^{*},\dots,{\bm{f}}_{n_{\mathcal{F}^{*}}}^{*}\}$
20:Sort $\mathcal{F}^{*}$ by descending flow and length
An illustration of Algorithm 2 is given in Figure 5. In Figure 5(a) is the flow map before node snapping, where we observe that the boundary points (black solid circles) of some of the purple linestrings are closer to each than the snapping tolerance $\varepsilon_{S}=4$ m. Figure 5(b) shows an update flow map after node snapping, and applying ST_OVERLINE_PLANR. There are fewer thin purple road segments, since by snapping their intersection nodes, they share more exact sub-segments, and these shared sub-segments are merged during the flow aggregation to result in wide orange road segments.
Refer to caption(a) Refer to caption(b)
Figure 5: Flow maps for unsnapped and node snapped segments, with tolerance $\varepsilon_{S}=4~{}\mathrm{m}$. (a) Unsnapped. (b) Node snapped. Colour (purple to orange) and width of road segments is proportional to traffic flow.
3.2 Node splitting to add missing intersection nodes
Node snapping is focused near the boundary points of the flow linestrings. So it does not address misalignments far from the boundary points. Moreover there remain some linestrings which do indeed intersect but whose intersection is not correctly computed by ST_OVERLINE_PLANR. The solution to both of these problems is the explicit addition of the missing intersection nodes to these linestrings.
ST_SPLITNODE is the collation of a couple of algorithms which explicitly add the nodes at the intersections in the interior of linestrings, and splits these linestrings into simple linestring segments at these added nodes. This gives the name to this method. The inputs to Algorithm 3 are the linestrings $\mathcal{F}$ and the node splitting $S$. For computational stability and efficiency, our method draws on two existing algorithms: in Steps 1–3, to_spatial_subdivision ($S=$ ‘subdivision’) in the sfnetworks package (van der Meer et al., 2023), and in Steps 4–5, geos_unary_union ($S=$ ‘unary’) in the geos package (Dunnington and Pebesma, 2023). The first option tends to find fewer intersection nodes, but this can be an advantage for our proposed line blending (to be introduced in the next subsection) since many small linestrings with similar flow values do not provide a clear prioritisation of this blending. It is also similar to the intersection computation performed by ST_OVERLINE_PLANR. The second option finds more missing intersection nodes, and leads to more comprehensive flow aggregation. We require both types of node splitting to compute a minimal flow map.
Algorithm 3 ST_SPLITNODE – Split nodes at interior intersections of linestrings
1:Input: $\mathcal{F}$ flow linestrings, $S$ node splitting type
2:Output: $\mathcal{F}^{*}$ node split flow linestrings
3:Initialise local network $\mathcal{N}^{*}$ with linestrings $\mathcal{F}$
4:if $S==$ ‘subdivision’ then
5:     $\mathcal{F}^{*}:=$ sfnetworks::to_spatial_subdivision$(\mathcal{N}^{*})$
6:else if $S==$ ‘unary’ then
7:     $\mathcal{F}^{*}:=$ geos::geos_unary_union$(\mathcal{N}^{*})$
In Figure 6(a) is the flow map without any node splitting or node snapping. This map has missing intersection nodes, and intersections nodes which are close to each other. In Figure 6(b) is the flow map after node splitting ($S=$ ‘unary’) to add the missing intersections nodes, and then node snapping (with snap tolerance $\varepsilon_{S}=4$ m). The result is that there are fewer road segments with higher flow values. Due to the combined action of ST_SPLITNODE and ST_SNAPNODE, Figure 6(b) is an improvement over Figures 5(a–b) and 6(a).
Refer to caption(a) Refer to caption(b)
Figure 6: Flow maps for unsplit/unsnapped and node split/node snapped segments, with tolerance $\varepsilon_{S}=4~{}\mathrm{m}$. (a) Unsplit and unsnapped. (b) Node split and node snapped. Colour (purple to orange) and width of road segments is proportional to traffic flow.
3.3 Line blending to align similar linestrings with local reference
So far we have focused on improving the alignment of linestrings induced by resolving inconsistencies at their intersections. We now focus on aligning linestrings more generally. For this, we require a comparison relationship to establish an order of alignment of nearby linestrings. For two linestrings ${\bm{f}}_{1}$ and ${\bm{f}}_{2}$ from a flow map $\mathcal{F}$, we define that ${\bm{f}}_{2}$ is a candidate to be aligned to the reference ${\bm{f}}_{1}$ at threshold $\varepsilon\geq 0$ if
${\bm{f}}_{2}\subseteq{\tt ST\_BUFFER}({\bm{f}}_{1},\varepsilon),\ f_{2}\leq f_{1}.$ (1)
This relation is insensitive to any local complexities in ${\bm{f}}_{2}$ as long as they are all contained within the buffer zone around ${\bm{f}}_{1}$. The buffer zone we employ has flat edges, e.g. for the R package sf, this corresponds to ST_BUFFER(endCapStyle="FLAT"), so the buffer zone ends at the boundary points of ${\bm{f}}_{1}$. This avoids erroneously considering neighbouring segments of ${\bm{f}}_{1}$, which are connected to ${\bm{f}}_{1}$ as a part of a longer sequence of road segments, to be candidates to be aligned to ${\bm{f}}_{1}$. The condition on the flow values means that we place a higher priority on linestrings with higher flows. Since we can define a local reference linestring, a global road network graph is no longer required to align the candidate linestrings.
Let ${\bm{f}}_{\rm ref}$ be a reference linestring from $\mathcal{F}$. The set of $m$ linestrings from \$\mathcal{F}\backslash\{{\bm{f}}_{\rm ref}\}$ which satisfy Equation (1) are the candidate linestrings $\mathcal{F}_{\rm cand}=\{{\bm{f}}_{{\rm cand},1},\dots,{\bm{f}}_{{\rm cand},n_{\mathcal{F}_{\rm cand}}}\}$, with the convention that if $n_{\mathcal{F}_{\rm cand}}=0$ then $\mathcal{F}_{\rm cand}$ is the empty set. We call our approach ‘line blending’, since we will blend the candidate linestrings onto the reference linestring. The inputs in Algorithm 4 are the reference linestring ${\bm{f}}_{\rm ref}$, the set of $m$ candidate linestrings $\mathcal{F}_{\rm cand}$, and the blend tolerance $\varepsilon$. The output are the modified reference and $m$ modified candidate flow linestrings, all with added interior points for accurate calculation of exactly equal linestring segments. In Step 1, we initialise a local network graph $\mathcal{N}^{*}$ with the reference linestring. In Steps 2–3, we extract all points of the candidate linestrings, and use the network_blend function in the sfnetworks package (which we denote as ST_NETWORK_BLEND) to blend efficiently these points into the reference linestring (van der Meer et al., 2023). This network blending requires a blend threshold, which we set to be $\varepsilon$. ST_NETWORK_BLEND projects the candidate linestrings onto the reference linestring, and explicitly adds them to the network, thereby creating new edges in $\mathcal{N}^{*}$. The result is a local network graph with more, shorter edges and with nodes at the projected candidate points, and whose union is the reference linestring. In Steps 4–5, we extract the $m$ blended candidate linestrings with these added interior points by applying the shortest path search between the start and end point of each blended candidate linestring along the network graph $\mathcal{N}^{*}$. Step 6 is the equivalent for the blended reference linestring. Step 7 involves collating the blended candidate linestrings.
Algorithm 4 ST_LINEBLEND – Blend candidate linestrings onto reference linestring
1:Input: ${\bm{f}}_{\rm ref}$ reference, $\mathcal{F}_{\rm cand}$ candidate flow linestrings, $\varepsilon$ blend tolerance
2:Output: $({\bm{f}}^{*}_{\rm ref},\mathcal{F}^{*}_{\rm cand})$ blended reference and $m$ candidate flow linestrings
3:Initialise local network graph $\mathcal{N}^{*}$ with ${\bm{f}}_{\rm ref}$
4:Extract all points $G_{\rm cand}$ of candidate linestrings and flow values $f_{\rm cand}$ from $\mathcal{F}_{\rm cand}$
5:Update network by blending candidate points $\mathcal{N}^{*}:={\tt ST\_NETWORK\_BLEND}(\mathcal{N}^{*},G_{\rm cand},\varepsilon)$
6:for $i:=1$ to $n_{\mathcal{F}_{\rm cand}}$ do
7:     ${\bm{f}}^{*}_{{\rm cand},i}$ := shortest path from $\operatorname{Start}({\bm{f}}_{{\rm cand},i})$ to $\operatorname{End}({\bm{f}}_{{\rm cand},i})$ along network $\mathcal{N}^{*}$
8:${\bm{f}}^{*}_{{\rm ref}}$ := shortest path from $\operatorname{Start}({\bm{f}}_{{\rm ref}})$ to $\operatorname{End}({\bm{f}}_{{\rm ref}})$ along network $\mathcal{N}^{*}$
9:Collate $\mathcal{F}^{*}_{\rm cand}:=\{{\bm{f}}_{{\rm cand},1}^{*},\dots,{\bm{f}}_{{\rm cand},n_{\mathcal{F}_{\rm cand}}}^{*}\}$
The result from ${\tt ST\_LINEBLEND}$ is the modified reference linestring and the projected candidate linestrings with added interior points. These added interior points resolve a limitation of the flow aggregation of ST_OVERLINE_PLANR (Morgan and Lovelace, 2021). If there are many candidate linestrings, then this may lead to many sub-segments in the projected linestrings, each with their own flow values. So we assign the weighted mean flow, weighted by the sub-segment length, to all sub-segments. This single flow value takes into account the contribution of each candidate linestring to the flow along the reference linestring. Moreover, we take the rounded value of this weighted mean flow to expedite the aggregation computations and to reduce visual clutter of the flow map.
In Figure 7(a), we illustrate this line blending with the linestrings $ABC$ with flow $f_{ABC}=7$ in blue, and $AD$ with flow $f_{AD}=2$ in orange. As these two linestrings do not share a common sub-segment, so flow aggregation does not modify them. Since $AD\subset{\tt ST\_BUFFER}(ABC,4~{}\mathrm{m})$ (the pale blue rectangle), and $f_{AD}<f_{ABC}$, then $ABC$ is the reference ${\bm{f}}_{\rm ref}$ and $AD$ is the candidate linestring ${\bm{f}}_{\rm cand}$, according to Equation (1). Line blending involves projecting $AD$ to $ABC$, so $D$ is projected to $D^{\prime}$ (which lies exactly on the reference linestring), and $B$ is added explicitly to this projected linestring. We also add $D^{\prime}$ to the reference linestring. The reference linestring becomes ${\bm{f}}^{*}_{\rm ref}=(7,ABD^{\prime}C)$, and the candidate ${\bm{f}}^{*}_{\rm cand}=(2,ABD^{\prime})$. Now ${\bm{f}}^{*}_{\rm ref},{\bm{f}}^{*}_{\rm cand}$ share the sub-segment $ABD^{\prime}$ exactly, and ST_OVERLINE_PLANR gives the aggregated flows as $(9,ABD^{\prime})$ and $(2,D^{\prime}C)$. We take the rounded value of the weighted mean of these two flow linestrings. The result of ST_LINEBLEND with ST_OVERLINE_PLANR is a single linestring $(9,ABD^{\prime}C)$, as shown in Figure 7(b).
Refer to caption(a) Refer to caption(b)
Figure 7: Line blending removes small misalignments and leads to correct flow aggregation. (a) Before line blending and flow aggregation. Reference linestring $(7,ABC)$ in blue, candidate linestring $(2,AD)$ in orange. (b) After blending candidate into reference linestring, with tolerance $\varepsilon=4$ m, and flow aggregation. Blended linestring is $(9,ABD^{\prime}C)$.
More complex situations arise when other linestrings touch the candidate linestring, but are not candidates themselves for blending. Since line blending projects the candidates to the reference linestring, then we have to also project these other linestrings to avoid leaving a gap in the updated flow map. This procedure ST_SNAP_CAND_TOUCH is outlined in Algorithm 5. The inputs are the reference linestring ${\bm{f}}_{\rm ref}$, the $n_{\mathcal{F}_{\rm cand}}$ candidate linestrings $\mathcal{F}_{\rm cand}$, the $n_{\mathcal{F}_{\rm ct}}$ candidate-touching linestrings $\mathcal{F}_{\rm ct}$, and the snap tolerance $\varepsilon_{S}$. In Steps 1–8, we iterate over each candidate-touching linestring. In Steps 2–3, we compute the intersection points between the candidate-touching linestring and the boundary points of the candidate linestrings, and the respective distances. In Steps 4–7, if this intersection point is within $\varepsilon_{S}$ distance to the boundary points of ${\bm{f}}_{\rm ref}$, then we snap the candidate-touching linestring to the closest boundary point. In Step 8, if the intersection point is not within $\varepsilon_{S}$ distance, then we snap it to the nearest point on ${\bm{f}}_{\rm ref}$. These snapping operations are similar to those in Steps 15–16 in ST_SNAPNODE in Algorithm 2, and ensure that we maintain connectivity between $\mathcal{F}_{\rm ct}$ and ${\bm{f}}_{\rm ref}$. In Step 9, we collate these snapped linestrings.
Algorithm 5 ST_SNAP_CAND_TOUCH – Snap candidate-touching linestrings onto reference
1:Input: ${\bm{f}}_{\rm ref}$ reference, $\mathcal{F}_{\rm cand}$ candidate, $\mathcal{F}_{\rm ct}$ candidate-touching flow linestrings, $\varepsilon_{S}$ snap tolerance
2:Output: $\mathcal{F}^{*}_{\rm ct}$ snapped candidate-touching flow linestrings
3:for $i:=1$ to $n_{\mathcal{F}_{\rm ct}}$ do
4:     ${\bm{g}}_{{\rm ct},i}^{*}:={\bm{f}}_{{\rm ct},i}\cap\{\operatorname{Start}(\mathcal{F}_{{\rm cand}}),\operatorname{End}(\mathcal{F}_{{\rm cand}})\}$
5:     $d_{i,\operatorname{Start}}^{*}:={\tt ST\_DIST}({\bm{g}}_{{\rm ct},i}^{*},\operatorname{Start}({\bm{f}}_{\rm ref}))$; $d_{i,\operatorname{End}}^{*}:={\tt ST\_DIST}({\bm{g}}_{{\rm ct},i}^{*},\operatorname{End}({\bm{f}}_{\rm ref}))$
6:     if ($d_{i,\operatorname{Start}}^{*}\leq\varepsilon_{S}$ and $d_{i,\operatorname{Start}}^{*}\leq d_{i,\operatorname{End}}^{*}$then
7:         ${\bm{f}}^{*}_{{\rm ct},i}:=$ snap ${\bm{f}}_{{\rm ct},i}$ to $\operatorname{Start}({\bm{f}}_{\rm ref})$
8:     else if ($d_{i,\operatorname{End}}^{*}\leq\varepsilon_{S}$ and $d_{i,\operatorname{End}}^{*}\leq d_{i,\operatorname{Start}}^{*}$then
9:         ${\bm{f}}^{*}_{{\rm ct},i}:=$ snap ${\bm{f}}_{{\rm ct},i}$ to $\operatorname{End}({\bm{f}}_{\rm ref})$
10:     else ${\bm{f}}^{*}_{{\rm ct},i}:=$ snap ${\bm{f}}_{{\rm ct},i}$ to nearest point of ${\bm{f}}_{\rm ref}$
11:Collate $\mathcal{F}^{*}_{\rm ct}:=\{{\bm{f}}^{*}_{{\rm ct},1},\dots,{\bm{f}}^{*}_{{\rm ct},n_{\mathcal{F}_{\rm ct}}}\}$
Figure 8(a) is an illustration of ST_SNAP_CAND_TOUCH with the reference $ABC$ with flow $f_{ABC}=7$ (blue), the candidate linestring $AD$ with flow $f_{AD}=2$ (orange), and the blending buffer zone with blend tolerance $\varepsilon=4$ m (light blue). The candidate-touching linestring ${\bm{f}}_{\rm ct}=(5,DE)$ (purple) touches the candidate $AD$ at $D$. The orange line is entirely within the pale blue buffer zone around the blue reference line, whereas the purple line extends outside of the buffer zone and so is not a candidate for blending to the reference linestring. If we apply ST_LINEBLEND to blend the candidate linestring, then $D$ is projected to $D^{\prime}$ to lie on the reference linestring, i.e. ${\bm{f}}_{\rm cand}^{*}=(2,ABD^{\prime})$. If we apply ST_SNAP_CAND_TOUCH to the candidate-touching linestring, then as the intersection point $D$ is less than $\varepsilon_{S}=4$ m from the boundary points of ${\bm{f}}_{\rm ref}$, it is projected to the boundary point $C$, i.e. ${\bm{f}}^{*}_{\rm ct}=(5,CE)$. Thus line blending, candidate-touching snapping and flow aggregation yields that ${\bm{f}}^{*}_{\rm ct}=(5,CE)$ remains connected to ${\bm{f}}_{\rm ref}^{*}=(9,ABD^{\prime}C)$.
Refer to caption(a) Refer to caption(b)
Figure 8: Snap candidate-touching linestrings to reference linestring after line blending, with blend tolerance $\varepsilon=4$ m, snap tolerance $\varepsilon_{S}=4$ m. (a) Before line blending and snapping. Reference linestring $(7,ABC)$ in blue, candidate linestring $(2,AD)$ orange, and candidate touching linestring $(5,DE)$ purple. (b) After line blending and snapping. Reference linestring becomes $(9,ABD^{\prime}C)$, and candidate-touching linestring $(5,CE)$.
So far we have only considered line blending where the reference linestring has a single flow. Due to the noisiness of the empirical trajectories, this is insufficient to reach a minimal flow map. So we allow the reference to be a sequence of $k$ connected edges, each with their own flow. If we treat this edge sequence momentarily as a single linestring, with flow equal to the weighted mean flow, weighted by the length of the $k$ edges, then we can apply Equation (1) to search for potential candidate linestrings. Due to computational limitations, we retain that candidates be single edges with single flow values. The generalisation to $k$-edge reference linestrings allows us to blend candidate linestrings which exceed the buffer zones of 1-edge reference linestrings.
We have described the situation for blending candidate and candidate-touching linestrings to a single reference linestring (possibly composed of $k$ edges). The next step to determine the priority for line blending for a set of reference linestrings. We require that a reference linestring cannot be a candidate linestring to another reference linestring, and a candidate linestring is a candidate for one reference linestring only. These ensure that if two linestrings ${\bm{f}}_{2}\subseteq{\tt ST\_BUFFER}({\bm{f}}_{1},\varepsilon)$ and ${\bm{f}}_{1}\subseteq{\tt ST\_BUFFER}({\bm{f}}_{2},\varepsilon)$, then we select only one of them. The flow condition $f_{2}\leq f_{1}$ usually distinguishes the reference linestring, except for where $f_{1}=f_{2}$. In the situation of equal flow values, then we designate the longer linestring to be the reference.
The inputs to ST_LINEBLEND_PRIORITY in Algorithm 6 are the flow linestrings $\mathcal{F}$, the number of connected edges for the reference linestrings $k$, and the blend tolerance $\varepsilon$. The output is a set of non-overlapping reference linestrings $\mathcal{F}_{\rm ref}$, a set of unique candidate linestrings $\mathcal{F}_{\rm cand}$ and a set of (potentially repeated) candidate-touching $\mathcal{F}_{\rm ct}$ linestrings. In Steps 1–2, we set up a network graph from the linestrings $\mathcal{F}$, and then extract $\mathcal{K}$, all simple paths (i.e. all paths composed of connected, unique edges) with $k$-edges. In Steps 3–4, we compute the weighted mean of the $k$ flow values, weighted by the length of individual edges, and assign this to the entire $k$-edge path. The function ${\rm Edges}(\cdot)$ extracts the edges from a $k$-edge path. In Step 5, we sort the flow linestrings, in descending order of their flow values and length. In Steps 6–8, we construct a maximal set of non-overlapping $k$-edge paths $\mathcal{K}^{*}$. We initialise $\mathcal{K}^{*}$ to contain the first path. We iterate through $\mathcal{K}$ and, if this path in $\mathcal{K}$ does not overlap the current $\mathcal{K}^{*}$, then we add it to $\mathcal{K}^{*}$. For $k=1$, we bypass Steps 2–8. In Steps 9–18, we step through the ${\bm{k}}_{i}^{*}\in\mathcal{K}^{*}$, starting from the linestring with the highest flow and length. In Steps 11–13, if the current flow linestring ${\bm{k}}_{i}^{*}$ has no incident edges in the candidate linestrings $\mathcal{F}_{\rm cand}$, then we set it to be a reference linestring. In Step 13, we search for candidate linestrings for ${\bm{k}}_{i}^{*}$, from those linestrings which are not already reference or candidate linestrings, according to Equation (1). In Steps 14–17, if there is at least one candidate linestring, then we update $\mathcal{F}_{{\rm ref}},\mathcal{F}_{\rm cand}$ and $\mathcal{F}_{\rm ct}$. In Step 18, we extract the candidate $\mathcal{F}_{\rm cand}$ and candidate-touching $\mathcal{F}_{\rm ct}$ linestring sets, since the reference linestring set is already computed as $\mathcal{F}_{\rm ref}$.
Algorithm 6 ST_LINEBLEND_PRIORITY – Compute line blending priority
1:Input: $\mathcal{F}$ flow linestrings, $k$edges, $\varepsilon$ blend tolerance
2:Output: $(\mathcal{F}_{\rm ref},\mathcal{F}_{\rm cand},\mathcal{F}_{\rm ct})$ reference, candidate, candidate-touching linestrings
3:Initialise network graph $\mathcal{N}^{*}$ from linestrings $\mathcal{F}$
4:Extract all simple paths of length $k$ edges $\mathcal{K}:=\{{\bm{k}}_{1},\dots,{\bm{k}}_{n_{\mathcal{K}}}\}$ from $\mathcal{N}^{*}$
5:for $i:=1$ to $n_{\mathcal{K}}$ do
6:     $f_{i}$ := weighted mean of $k$ flows, weight := len(Edges(${\bm{k}}_{i}$))
7:Sort $\mathcal{K}$ by descending flow and combined length
8:Initialise $\mathcal{K}^{*}:=\{{\bm{k}}_{1}\}$
9:for $i:=2$ to $n_{\mathcal{K}}$ do
10:     if (${\bm{k}}_{i}\cap\mathcal{K}^{*}=\{\}$then $\mathcal{K}^{*}:=\mathcal{K}^{*}\cup\{{\bm{k}}_{i}\}$
11:Initialise $\mathcal{Q}:=\mathcal{F}_{\rm ref}:=\mathcal{F}_{\rm cand}:=\mathcal{F}_{\rm ct}:=\{\}$
12:for $i:=1$ to $n_{\mathcal{K}^{*}}$ do
13:     if (${\rm Edges}({\bm{k}}_{i}^{*})\notin\mathcal{F}_{\rm cand}$then
14:         Set reference linestring ${\bm{k}}_{{\rm ref}}:={\bm{k}}_{i}^{*}$
15:         Search candidate linestrings $\mathcal{F}_{{\rm cand}}^{*}:=\{{\bm{f}}\in\mathcal{F}\backslash({\rm Edges}(\mathcal{F}_{\rm ref})\cup\{{\bm{k}}_{\rm ref}\}\cup\mathcal{F}_{\rm cand}):{\bm{f}}\subseteq$
16:    ${\tt ST\_BUFFER}({\bm{k}}_{\rm ref},\varepsilon),f\leq f_{\rm ref}\}$
17:         if $\mathcal{F}_{\rm cand}^{*}\neq\{\}$ then
18:              Update $\mathcal{F}_{\rm ref}:=\mathcal{F}_{\rm ref}\cup\{{\bm{k}}_{\rm ref}\},\mathcal{F}_{\rm cand}:=\mathcal{F}_{\rm cand}\cup\mathcal{F}_{\rm cand}^{*}$
19:              Search touches $\mathcal{F}_{{\rm ct}}:=\{{\bm{f}}\in\mathcal{F}\backslash\{{\rm Edges}(\mathcal{F}_{\rm ref})\ \cup\mathcal{F}_{\rm cand}\}:{\tt ST\_TOUCHES}(\mathcal{F}_{{\rm cand}}^{*},{\bm{f}})\}$
20:              Update $\mathcal{Q}:=\mathcal{Q}\cup\{(\{{\bm{k}}_{\rm ref}\},\mathcal{F}_{\rm cand}^{*},\mathcal{F}_{\rm ct}^{*})\}$
21:Extract $\mathcal{F}_{\rm cand}:=\{\mathcal{F}^{*}_{{\rm cand},1},\dots,\mathcal{F}^{*}_{{\rm cand},n_{\mathcal{Q}}}\},\mathcal{F}_{\rm ct}:=\{\mathcal{F}^{*}_{{\rm ct},1},\dots,\mathcal{F}^{*}_{{\rm ct},n_{\mathcal{Q}}}\}$ from $\mathcal{Q}$
The local alignment of road segments, in Algorithm 7, is established by combining Algorithms 26. Its inputs are the flow linestrings $\mathcal{F}$, the split node type $S$, the number of connected edges for the reference linestrings $k$, the blend tolerance $\varepsilon$ and the snap tolerance $\varepsilon_{S}$. In Step 1, we node split the flow linestrings $\mathcal{F}$. Step 2, we node snap the flow linestrings. In Step 3, we set up the priority for the line blending, and in Step 4 we store all the linestrings that will not be modified by the line blending in $\mathcal{F}^{c}$. In Steps 5–8, we iterate over each reference linestring. In Step 6, we update the reference and candidate flow linestrings by blending the candidate linestrings. In Step 7, we apply ST_OVERLINE_PLANR within each of the $k$-edges of the updated reference linestrings. In Step 8, we compute the weighted mean flow, weighted by the length of the corresponding sub-segments, and assign it as the flow value to all sub-segments. The result from Steps 6–8 is a modified $k$-edge reference linestring with $k$ aggregated flows. In Step 9, we snap the candidate-touching linestrings to the reference linestring to maintain connectivity. In Steps 10–11, we collate the modified linestrings from Steps 5–9, with the unmodified linestrings $\mathcal{F}^{c}$ from Step 4, and sort them in descending flow and length.
Algorithm 7 ST_OVERLINE_LINEBLEND – Locally align road segment flows using line blending
1:Input: $\mathcal{F}$ flow linestrings, $S$ split node, $k$edges, $\varepsilon$ blend tolerance, $\varepsilon_{S}$ snap tolerance
2:Output: $\mathcal{F}^{*}$ aggregated aligned flow linestrings
3:$\mathcal{F}:={\tt ST\_SPLITNODE}(\mathcal{F},S)$
4:$\mathcal{F}:={\tt ST\_SNAPNODE}(\mathcal{F},\varepsilon_{S})$
5:$(\mathcal{F}_{\rm ref},\mathcal{F}_{\rm cand},\mathcal{F}_{\rm ct}):={\tt ST\_LINEBLEND\_PRIORITY}(\mathcal{F},k,\varepsilon)$
6:$\Edgessubscriptrefsubscriptcandsubscriptct\mathcal{F}^{c}:=\mathcal{F}\backslash({\rm Edges}(\mathcal{F}_{\rm ref})\cup\mathcal{F}_{\rm cand}\cup\mathcal{F}_{\rm ct})$
7:for $i:=1$ to $n_{\mathcal{F}_{\rm ref}}$ do
8:     $({\bm{f}}_{{\rm ref},i}^{*},\mathcal{F}_{{\rm cand},i}^{*}):={\tt ST\_LINEBLEND}({\bm{f}}_{{\rm ref},i},\mathcal{F}_{{\rm cand},i},\varepsilon)$
9:     ${\bm{f}}_{{\rm ref},i}^{*}:={\tt ST\_OVERLINE\_PLANR}({\bm{f}}_{{\rm ref},i}^{*}\cup\mathcal{F}_{{\rm cand},i}^{*})$
10:     $f^{*}_{{\rm ref},i}$ := weighted mean of ${\bm{f}}^{*}_{{\rm ref},i}$, weight $:=\mathrm{len}({\rm Edges}({\bm{f}}_{{\rm ref},i}^{*}))$
11:     $\mathcal{F}_{{\rm ct},i}^{*}:={\tt ST\_SNAP\_CAND\_TOUCH}({\bm{f}}_{{\rm ref},i}^{*},\mathcal{F}_{{\rm cand},i},\mathcal{F}_{{\rm ct},i},\varepsilon_{S})$
12:$\mathcal{F}^{*}:=\mathcal{F}_{\rm ref}^{*}\cup\mathcal{F}_{\rm ct}^{*}\cup\mathcal{F}^{c}$
13:Sort $\mathcal{F}^{*}$ in descending order of flow and length
We outline the overall workflow ST_OVERLINE to compute a minimal flow map, by combining pre- and post-processing with the iteration of local alignment. For Algorithm 8, the inputs are the map matched routes $\mathcal{M}$, the split node type $S$, the blend tolerance $\varepsilon$, the simplify tolerance $\varepsilon_{D}$, the snap tolerance $\varepsilon_{S}$, the number of connected edges in reference linestrings ${\bm{k}}$, and the maximum number of iterations $j_{\max}$. In Step 1, we apply ST_OVERLINE_PLANR to the map matched routes to produce an initial flow map $\mathcal{F}$. In Step 2, we apply ST_SPLITNODE to this flow map, because applying ST_SPLITNODE on a flow map is more robust than on the map matched routes. In Step 3, we employ the standard ST_SIMPLIFY with simplify tolerance $\varepsilon_{D}$. These simplified linestrings are modified so that all modified segments are at most $\varepsilon_{D}$ distance from the unmodified segments (Ramer, 1972, Douglas and Peucker, 1973). These simplified linestrings usually lead to more overlapping segments, which assist the flow aggregation in Step 4. Steps 5–10 is the iteration of ST_OVERLINE_LINEBLEND for the $k$-edge reference linestrings. For each $k$ in ${\bm{k}}$, we iterate until the maximum number of iterations $j_{\max}$ is reached or two consecutive flow maps are identical. Within each iteration, we search for the $k$-edge reference linestrings, blend the candidate linestrings, and snap the candidate-touching linestrings, and compute the flow aggregation. In Steps 11–14 is some housekeeping in ST_OVERLINE_PRUNE, where we remove pseudo nodes, and replace the incident edges with a concatenated edge with the weighted mean flow, as well as some loops and leaf edges with low flow values.
Algorithm 8 ST_OVERLINE – Compute locally aligned flow map from map matched routes
1:Input: $\mathcal{M}$ map matched routes, $S$ split node, $\varepsilon_{D}$ simplify tolerance, ${\bm{k}}$edges, $j_{\max}$ max #iterations, $\varepsilon$ blend tolerance, $\varepsilon_{S}$ snap tolerance
2:Output: $\mathcal{F}$ aligned road segment flows
3:$\mathcal{F}:={\tt ST\_OVERLINE\_PLANR}(\mathcal{M})$;
4:$\mathcal{F}:={\tt ST\_SPLITNODE}(\mathcal{F},S)$
5:$\mathcal{F}:={\tt ST\_SIMPLIFY}(\mathcal{F},\varepsilon_{D})$
6:$\mathcal{F}:={\tt ST\_OVERLINE\_PLANR}(\mathcal{F})$
7:for $k$ in ${\bm{k}}$ do
8:     $\mathcal{F}_{\rm prev}:=\{\}$; $j:=0$
9:     while $j<j_{\max}$ and $\mathcal{F}_{\rm prev}\neq\mathcal{F}$ do
10:         $\mathcal{F}_{\rm prev}:=\mathcal{F}$; $j:=j+1$
11:         $\mathcal{F}:={\tt ST\_OVERLINE\_LINEBLEND}(\mathcal{F},k,\varepsilon,\varepsilon_{S})$
12:         $\mathcal{F}:={\tt ST\_OVERLINE\_PLANR}(\mathcal{F})$
13:$\mathcal{F}_{\rm prev}:=\{\}$; $j:=0$
14:while $j<j_{\max}$ and $\mathcal{F}_{\rm prev}\neq\mathcal{F}$ do
15:     $\mathcal{F}_{\rm prev}:=\mathcal{F}$; $j:=j+1$
16:     $\mathcal{F}:={\tt ST\_OVERLINE\_PRUNE}(\mathcal{F})$
4 Results
In this section we compute a minimal flow map for the Hannover trajectories. From the complete set of 1183 trajectories, we keep 1177 trajectories with length greater than $100$ m. We input the 1177 trajectories into ST_ROUTE with $M$ as the map matching and $R$ the route finding APIs from the Valhalla routing engine, and ${\bm{n}}_{W}=3,13,23,33,43,63,83$ waypoints. We employ the dockerised version 3.4.0 of the Valhalla routing engine (GIS OPS, 2023). Of these input trajectories, 1147 yield a sufficiently high quality match, where the Hausdorff distance $d_{\operatorname{Haus}}(M^{*}(G),G)<100$ m, and the ratio $\operatorname{len}(M^{*}(G))/\operatorname{len}(G)<1.1$. We continue the analysis with these 1147 map matched routes.
For all iterations, we set $j_{\max}=20$. For the line blending, we begin with 1 iteration of ST_OVERLINE (Steps 1–4) with simplify tolerance $\varepsilon_{D}=1~{}\mathrm{m}$. We follow with ST_OVERLINE (Steps 5–14) with blend and snap tolerances $\varepsilon=\varepsilon_{S}=4~{}\mathrm{m}$, #edges ${\bm{k}}=1,2$, and node split type $S=$ ‘subdivision’. These simplify, blend and snap tolerances were chosen heuristically as a trade-off between being sufficiently large to account for the noisiness of the trajectories and the map matching/route finding APIs, whilst not being too large to obscure separate road segments within region with a dense road network. We set ${\bm{k}}=1,2$ since the search for connected $k$-edges with $k>2$ is too computationally intensive for our setup due to the large number of road segments (13 495). We set $S=$ ‘subdivision’ node splitting, as it provides more stable line blending priority at this early stage. We follow with 1 iteration of ST_OVERLINE (Steps 5–14), with the same tuning parameter choices, except with $S=$ ‘unary’. ‘Unary’ node splitting is usually applied after an iteration of ‘subdivision’ node splitting since the former can now add any missing intersection nodes without adversely affecting the line blending priority. We end with 2 iterations of ST_OVERLINE (Steps 5–14), with $\varepsilon=\varepsilon_{S}=5~{}\mathrm{m},{\bm{k}}=1,2,3,4,S=$ ‘unary’, since the larger values for $\varepsilon,\varepsilon_{S}$ capture any line blending missed at 4 m, and the connected 3-, 4-edge searches become computationally feasible with the lower number of road segments. These 5 iterations of ST_OVERLINE converge to a minimal flow map with 1 413 road segments, i.e. a 89.5% reduction of the 13 495 segments in the initial flow map.
Figure 9 illustrates the results of these iterations. In Figure 9(a) are the trajectories (green circles) and their map matched routes resulting from ST_ROUTE (blue lines). Whilst the map matched routes no longer completely obscure the road network, they remain misaligned to each other and to the road network. So it is not possible to accurately estimate the traffic flow map directly from them. In Figure 9(b) is a flow map resulting from ST_OVERLINE (Steps 1–4) with $S=$ ‘subdivision’. This is similar to the flow map that would be obtained by following Morgan and Lovelace (2021) without rasterisation. Since each linestring is labelled by its flow value, the crowding of the labels indicates that this is unlikely to be a minimal flow map. In Figure 9(c), we complete the line blending iteration ST_OVERLINE (Steps 5–14) with $S=$ ‘subdivision’. The crowding of the flow value labels is reduced so we are progressing to a minimal flow map. In Figure 9(d), we carry out a further 3 line blending iterations ST_OVERLINE (Steps 5–14) with $S=$ ‘unary’ to arrive at a minimal flow map.
Refer to caption(a) Refer to caption(b)
Refer to caption(c) Refer to caption(d)
Figure 9: Iterations to compute locally aligned flow map. (a) Empirical trajectories (green circles), map matched routes from ST_ROUTE (blue lines). (b) Initial flow map from ST_OVERLINE (1–4), $S=$ ‘subdivision’. (c) Intermediate flow map from 1 iteration of ST_OVERLINE (5–14), $S=$ ‘subdivision’. (d) Minimal flow map from 3 further iterations of ST_OVERLINE (5–14), $S=$ ‘unary’. Label is traffic flow. Colour (purple to orange) of road segments is proportional to traffic flow.
4.1 Validation
A visual inspection of the minimal flow map in Figure 9(d) reveals good alignment to the OSM road network in general, though there remain some data artefacts elsewhere. For example, in Figure 10(a), the green trajectories give no indication of the blue wonky route output by ST_ROUTE, and in Figure 10(c), the trajectories do not involve a loop, though ST_ROUTE contains a loop. Since our local alignment does not use an external road network to correct these wonky or extraneous loops in the map matched routes, they are propagated to the flow maps in Figure 10(a, d). Overall, these type of data artefacts are small and infrequent, and are associated with road segments with low flow.
Refer to caption(a) Refer to caption(b)
Refer to caption(a) Refer to caption(b)
Figure 10: Data artefacts in locally aligned road segment flows. (a–b) Wonky road segments. (c–d) Extraneous road segments. (a, c) Trajectories (green circles), map matched routes (blue lines). (b, d) Colour (purple to orange) of road segments is proportional to traffic flow.
For a more quantitative validation of the accuracy of our proposed flow map, we require a gold standard reference flow map. The experimental design of the Hannover trajectories is to serve as a reference data set for learning turning rules at road junctions, rather than as a reference flow map (Zourlidou et al., 2022). So we have to compute a proxy reference flow map. For this, we first compute line transects $\mathcal{T}_{i}=\{{\bm{t}}_{i,1},\dots,{\bm{t}}_{i,n_{\mathcal{T}_{i}}}\}={\tt ST\_TRANSECT}({\bm{f}}_{i},\varepsilon_{T},\delta_{T})$ for a road linestring ${\bm{f}}_{i}\in\mathcal{F}$, where a line transect is an orthogonal line segment (of length $2\varepsilon_{T}$ m) to ${\bm{f}}_{i}$. If $\operatorname{len}({\bm{f}}_{i})>\delta_{T}$ m, then these transects are placed at every $\delta_{T}$ m of ${\bm{f}}_{i}$, or if $\operatorname{len}({\bm{f}}_{i})\leq\delta_{T}$ m then the single transect is subtended from the centroid of ${\bm{f}}_{i}$. The empirical proxy flows are the number of intersection points between the map matched routes from $\mathcal{M}$ which intersect the line transects $f_{i,j}^{\mathrm{emp}}=\#\{{\tt ST\_INTERSECTION}(\mathcal{M},{\bm{t}}_{i,j}):{\bm{t}}_{i,j}\in\mathcal{T}_{i}\},j=1,\dots,n_{\mathcal{T}_{i}}$, and the mean discrepancy over all line transects of ${\bm{f}}_{i}$ is $\bar{f}_{i}^{\mathrm{emp}}=\mathrm{mean}\{f_{i,j}^{\mathrm{emp}},j=1,\dots,n_{\mathcal{T}_{i}}\}$. Our error measure of ${\bm{f}}_{i}$ is the absolute difference $\operatorname{Err}_{i,j}=|f_{i}-\bar{f}_{i}^{\mathrm{emp}}|$ for all line transects ${\bm{t}}_{j},j=1,\dots,n_{\mathcal{T}_{i}}$ which intersect ${\bm{f}}_{i}$. We also compute the relative error $\operatorname{RErr}_{i,j}=\operatorname{Err}_{i,j}/f_{i}$, which is well-defined since $f_{i}\geq 1$ for all ${\bm{f}}_{i}\in\mathcal{F}$.
With the tuning parameters tolerance $\varepsilon_{T}=5$ m and $\delta_{T}=50$ m, then we compute 7 380 line transects. The colour (purple to orange) of the line transects is proportional to the absolute error. In Figure 11(a), the vast majority of the error values are low (purple), with only a few high error values (orange), in/near the solid black rectangle. The zoom of the black rectangle is given in Figure 11(b), along with the road segment flows in blue. The road segment ${\bm{f}}_{11}$ (thick blue segment) with the estimated flow is $f_{11}=238$, and the empirical proxy flow $\bar{f}_{11}^{\mathrm{emp}}=254$, has a high absolute error, i.e. $\operatorname{Err}_{11}=16,\operatorname{RErr}_{11}=0.07$. The road segment ${\bm{f}}_{331}$ with $f_{331}=25,\bar{f}_{331}^{\mathrm{emp}}=12$, has a higher relative error, i.e. $\operatorname{Err}_{331}=13,\operatorname{RErr}_{331}=0.52$. These errors are due to that these road segments, at an earlier iteration of the flow map, comprised shorter sub-segments with higher flows due to the presence of the more than $100$ trajectories passing near the intersection of ${\bm{f}}_{11}$ and ${\bm{f}}_{331}$. Since the nodes of these sub-segments are pseudo nodes, they were removed, and the concatenated sub-segments assigned the single weighted mean flow = 25.
Refer to caption(a) Refer to caption(b)
Figure 11: Validation of line transects of locally aligned flow map. (a) Neighbourhood level. (b) Zoom of black rectangle. Line transects computed with tolerance $\varepsilon_{T}=5$ m, at every $\delta_{T}=50$ m. Label is absolute error (Err). Colour (purple to orange) of line transects is proportional to absolute error. Width of road segments (blue) is proportional to traffic flow.
To supplement the visual examination of these flow estimation errors, Figure 12 shows the bivariate (Err, RErr) histogram plot of the errors. The vast majority (6518 or 88.3%) in the orange hexagonal bin have zero absolute and relative error, and only 39 (0.53%) have Err $>4$, and 364 (4.9%) have RErr $>0.1$, and only 16 (0.21%) have Err $>4$ and RErr $>0.1$ which comprise most of the purple bins.
Figure 12: Histogram plot of absolute and relative errors of line transects of locally aligned flow map. Label in hexagonal bin is number of line transects. Vertical dotted line is Err = 4, horizontal dotted line is RErr = 0.1.
This demonstrates that our flow map is mostly accurate as it usually controls the estimation error. By accuracy, we mean by how close the estimated flows are to the empirical proxy flows. However accuracy is insufficient on its own: if we take the extreme situation of a flow map with a single road segment with zero error, then this has the highest accuracy, but we have no estimated flows outside of this single segment. So we require that the spatial coverage of the flow map be also high. We cannot resolve this question of spatial coverage unambiguously since we do not have a gold standard flow map, though we can verify that all $n=1147$ map matched routes intersect at least one line transect from the estimated flow map. Whilst this calculation does not exclude that there can be some regions of some map matched routes are without nearby line transects, we can further verify with a visual inspection that these regions tend to be small in area and/or comprise a low number of routes. So we claim that our minimal flow map has high levels of accuracy and spatial coverage.
To conclude the validation of our flow maps, the flow map at the city level is illustrated in Figure 13(a). The orange segments with high flows are apparent, whereas these high flows were not apparent from the scatter plot of the trajectories in Figure 1(a). The desire line map (or spider diagram) is in Figure 13(b) and represent the traffic flows between the origin/destination hub nodes (black solid circles) These hub nodes are the centroids from a hierarchical clustering, similar to that in ST_SNAPNODE, of the boundary points of trajectories with the cutoff at 5 000 m. The straight lines indicate the trajectories whose origin/destination are associated with different hub nodes, and the circles indicate the trajectories whose origin/destination are associated with the same hub node. The desire line map is in effect a low resolution map of straight-line flows between the hub nodes only, whereas the high resolution map shows the traffic flows on all road segments.
Refer to caption(a) Refer to caption(b)
Figure 13: High and low resolution flow maps. (a) Locally aligned flow map. (b) Desire line map. Colour (purple to orange) and width of road segments is proportional to traffic flow.
4.2 Software
These analysis algorithms have been developed in R since the complexity of Algorithms 18 require a mix of advanced geospatial and statistical methods. As a compiled language, R can have slower execution times. There are two main computational bottlenecks. The first consists of the map matching/route finding in Algorithm 1. The Valhalla routing engine APIs are available as a web-based service (e.g. https://valhalla1.openstreetmap.de) or as a local dockerised image (e.g. https://github.com/gis-ops/docker-valhalla). We use the latter as it allows for faster computation since these local API requests are not sent to a remote web-based server, and can be parallelised on a stand-alone machine. We conduct a small study of the execution times based on 10 replicates of ST_ROUTE on 10 randomly selected trajectories on an Intel i5 Quad core 3.10 GHz machine running Ubuntu 22.04 and R 4.4.0. Executing ST_ROUTE with a local Valhalla API is around 7.1 times faster than the web-based API, and a parallelisation (with 3 workers) is around 1.8 times faster than a serial computation. This is less than 3 because the dockerised image is not optimised for simultaneous API calls. Combining these together, a parallelised local API achieves around a 12.6 fold speed improvement in comparison to a serial web-based API.
The second bottleneck is the line blending in Algorithm 7. Based on the execution times based on 10 replicates of ST_OVERLINE on a subset of 433 trajectories, 3-worker parallelisation is around 1.8 times faster than a serial computation. This is less than 3 because only the repeated application of line blending (Algorithm 4) is parallelised, whilst the line blending priority (Algorithm 6) remains a serial computation. These speed factors are intended to be illustrative, since execution times, involving remote web servers APIs and parallelisation, are difficult to predict on different internet connections and machines. We tentatively claim that a local Valhalla API reduces the execution time by an order of magnitude, whereas parallelisation reduces it almost linearly by the number of workers.
Our long term goal is to integrate fast implementations of these local alignment algorithms into an industry standard GIS, so GIS practitioners will be able to efficiently create flow maps from empirical trajectories in a familiar environment. Since this is under development, in the mean time, we provide a geopackage and QGIS project with the input trajectories, map matched routes, iterated flow maps and desire lines, as listed in Table 1. The interested reader is able to interactively explore in QGIS the added value of our proposed high resolution minimal flow map flowmap4, in comparison to the input trajectories traj, the map matched routes route, the desire line flow map flowmap_desire, and the flow map computed according to a leading alternative without rasterization Morgan and Lovelace (2021) which is similar to flowmap0.
Layer Description $n$
traj Empirical trajectories 1 147
route Map matched routes ST_ROUTE 1 147
flowmap0 Flow map ST_OVERLINE(1–4), $\varepsilon_{D}=1~{},S=$ ‘subdivision’ 13 495
flowmap1 Flow map ST_OVERLINE(5–14), ${\bm{k}}=1,2,\varepsilon=\varepsilon_{S}=4,S=$ ‘subdivision’ 2 437
flowmap2 Flow map ST_OVERLINE(5–14), ${\bm{k}}=1,2,\varepsilon=\varepsilon_{S}=4,S=$ ‘unary’ 1 953
flowmap3 Flow map ST_OVERLINE(5–14), ${\bm{k}}=1,2,3,4,\varepsilon=\varepsilon_{S}=5,S=$ ‘unary’ 1 867
flowmap4 Flow map ST_OVERLINE(5–14), ${\bm{k}}=1,2,3,4,\varepsilon=\varepsilon_{S}=5,S=$ ‘unary’ 1 413
flowmap_desire Desire lines ST_DESIRELINE, $\varepsilon=5000$ 41
Table 1: Geospatial layers in geopackage. The first column is layer name, the second is description, and the third is number of geospatial features $n$.
5 Conclusion
We have introduced novel analysis algorithms to compute a flow map from empirical trajectories. Our starting point is to focus on aligning segments of the map matched routes rather than the complete routes. We define a spatial relation to set up local reference road segments, which allows us to align other nearby road segments to this local reference segment. This local alignment is the key innovation to computing a minimal flow map that is aligned to the underlying road network. We presented solid evidence for the high level of spatial resolution, accuracy and coverage for our proposed minimal flow map. Since it accurately shows the traffic flow on all road segments at all scales, it provides increased added value in comparison to the empirical trajectories, to the low resolution desire lines map, and to existing high resolution flow map methodologies.

### v5

@@
 Abstract
-Desire line maps are widely deployed for traffic flow analysis by virtue of their ease of interpretation and computation.
-They can be considered to be simplified traffic flow maps, whereas the computational challenges in aggregating small scale traffic flows prevent the wider dissemination of high resolution flow maps.
-Vehicle trajectories are a promising data source to solve this challenging problem.
-The solution begins with the alignment (or map matching) of the trajectories to the road network.
-However even the state-of-the-art map matching implementations produce sub-optimal results with small misalignments.
-While these misalignments are negligible for large scale flow aggregation in desire line maps, they pose substantial obstacles for small scale flow aggregation in high resolution maps.
-To remove these remaining misalignments, we introduce innovative local alignment algorithms, where we infer road segments to serve as local reference segments, and proceed to align nearby road segments to them.
-With each local alignment iteration, the misalignments of the trajectories with each other and with the road network are reduced, and so converge closer to a minimal flow map.
-By analysing a set of empirical trajectories collected in Hannover, Germany, we confirm that our minimal flow map has high levels of spatial resolution, accuracy and coverage.
-Keywords: Desire lines, floating car data FCD, GPS, map matching, route finding
+Vehicle trajectories are a promising GNSS (Global Navigation Satellite System) data source to compute multi-scale traffic flow maps ranging from the city/regional level to the road level.
+The main obstacle is that trajectory data are prone to measurement noise.
+While this is negligible for city level, large-scale flow aggregation, it poses substantial difficulties for road level, small-scale aggregation.
+To overcome these difficulties, we introduce innovative local alignment algorithms, where we infer road segments to serve as local reference segments, and proceed to align nearby road segments to them.
+We deploy these algorithms in an iterative workflow to compute locally aligned flow maps.
+By applying this workflow to synthetic and empirical trajectories, we verify that our locally aligned flow maps provide high levels of accuracy and spatial resolution of flow aggregation at multiple scales for static and interactive maps.
+Keywords: GNSS, GPS, rasterisation, multi-scale, route finding
 1 Introduction
-One of the fundamental quantities in transport planning is a traffic flow map, i.e.
-a map of the traffic flow levels on the road segments in a road network (Ortúzar and Willumsen, 2011).
+One of the fundamental quantities in transport planning is a traffic flow map, i.e., a map of the traffic flow levels on the road segments in a road network (Ortúzar and Willumsen, 2011).
 Whilst traffic flow maps are a rich source of information about vehicle mobility patterns, they are costly in terms of time and resources to compute for any reasonably sized road network.
-To alleviate this cost burden, most approaches restrict the spatial coverage and resolution of the flow map.
-One of the most common is to place sensors at fixed locations in the road network, whose results are then visualised as a traffic count map.
-Another are trip intent/recall questionnaires to inform large scale properties such as trajectory origins and destinations, which can be visualised as a desire line flow map/spider diagram (Tobler, 1987).
+To alleviate this cost burden, most approaches restrict the spatial resolution of the flow map.
+One of the most common is to place sensors at fixed locations in the road network, whose results are then visualized as a traffic count map.
+Another common approach are trip intent/recall questionnaires to inform large-scale properties such as trajectory origins and destinations, whose results are then visualized as a desire line flow map/spider diagram (Tobler, 1987).
 Both of these are simplified flow maps, since the detailed mobility patterns outside of the sensor locations or origin/destination pairs are not known.
@@
 While these model-assigned routes are highly detailed, the trade-off is that they are not guaranteed to correspond closely to empirical mobility patterns.
-Thus an ideal data source combines the empirical information of road sensor counts or questionnaires, with the small scale details of model-assigned trajectories.
-This gap in the market can be filled by trajectory data.
-Due to the prevalence of geolocation-enabled devices, such as vehicle navigation guides and mobile telephones, trajectory data can be acquired with low marginal cost whilst at the same time offering extensive spatial coverage and resolution of empirical mobility patterns (Herrera et al., 2010, Andrienko and Andrienko, 2013).
-Throughout, we employ the term ‘trajectory data’ rather than Floating Car Data (FCD) or Floating Mobile Data (FMD), as our analysis is not restricted to trajectory data from cars or from mobile phones, so these terms are equivalent for our purposes.
-An example is the 1183 trajectories are collected from a GPS-enabled mobile phone, from December 2017 to March 2019 by a single driver in Hannover, Germany, with an overall average sampling rate about 1 point per second (Zourlidou et al., 2022).
-They are plotted as the green circles in Figure 1.
-Refer to caption(a)
-Refer to caption(b)
-Refer to caption(c)
-Figure 1: Trajectories in Hannover, Germany.
-(a) City level.
-(b) Neighbourhood level, zoom of black rectangle.
-(c) Small neighbourhood level, zoom of blue rectangle.
-Trajectory ID = 7 (orange circles), 315 (purple diamonds), others (green circles).
-In Figure 1(a), at the city level, the points (green circles) appear to be aligned to the road network.
-If we zoom in on the small black rectangle in the central region, then in Figure 1(b) at the neighbourhood level, we observe that the points deviate from the road network.
-This deviation is clearer in the closer zoom in Figure 1(c).
-Moreover, if we focus on the orange circles (Trajectory ID = 7) and purple diamonds (ID = 315), then we observe that the vehicle location between the recorded points is unknown.
-These maps illustrate the errors in trajectories.
-These are broadly classified as ‘measurement errors’ where the recorded coordinates are not the true locations, and ‘sampling errors’ where the information about the trajectory is lost in between recorded coordinates (Ortúzar and Willumsen, 2011, Saki and Hagen, 2022).
-Our goal is to produce a traffic flow map from these noisy trajectories, which can be utilised at any scale, from the city/regional level to the individual road segment level.
-This requires us to minimise the errors in trajectories.
-Our approach is composed of two stages.
-The first stage is to align the trajectories to the road network, which is known as map matching (Quddus et al., 2007, Chao et al., 2020).
-It produces a route, which is a connected sequence of road segments in the road network, that is consistent with the trajectory.
-Our contribution is an improvement to standard map matching by adding post hoc route finding.
-In common with many open source transport planning tools, we employ APIs based on the OpenStreetMap (OSM) network.
-While we are able to improvement the overall alignment to the road network, these map matched routes inherit incompressible misalignments, ranging from several centimetres to several metres, from the OSM road network.
-These small misalignments prevent the accurate aggregation of traffic flows at this scale.
-The second stage is to resolve these misalignments, and our contribution here is the proposed local alignment of map matched routes.
-In contrast to global alignment approaches, we locally infer which road segments should serve as a local reference, and then proceed to align other nearby road segments to it.
+Trajectory data, derived from GNSS (Global Navigation Satellite Systems), combine the large-scale information of road sensor counts or questionnaires, with the small-scale details of model-assigned trajectories.
+Due to the prevalence of GNSS-enabled devices, such as vehicle navigation guides and mobile telephones, trajectory data can be acquired with low marginal cost whilst at the same time offering extensive spatial coverage and resolution of empirical mobility patterns (Herrera et al., 2010, Andrienko and Andrienko, 2013).
+Our motivating example of trajectory data are the 1183 trajectories in EPSG 32632 (WGS 84/UTM zone 32N) projected coordinates, collected from a GNSS-enabled mobile phone, from December 2017 to March 2019 by a single driver in Hannover, Germany, with an overall average sampling rate about 1 point per second (Zourlidou et al., 2022).
+In Figure 1(a), at the city level, the green circles appear to be aligned to the road network.
+Figure 1: Trajectory data in Hannover, Germany.
+(a) Empirical trajectories (green circles).
+Klusreide road segment (solid yellow circle).
+(b) Zoom of trajectories and map matched routes (blue lines), with road network (solid black lines), near Klusreide.
+(d) Zoom of 2 m $\times$ 2 m rasterised flow aggregation.
+Colour scale is traffic flow in grid cells.
+If we zoom in on these trajectory points around Klusriede (solid yellow circle), a residential street near the A2 autobahn, in Figure 1(b), then we observe that the points deviate from the road network (solid black lines).
+These noisy trajectories are unsuitable for direct aggregation into a flow map, so the usual first step is to apply a map matching procedure (Quddus et al., 2007, Chao et al., 2020).
+The map matched routes are the blue lines in Figure 1(b).
+These follow the road network more closely.
+Map matching is a highly difficult problem and map matching with exact alignment to the underlying road network in a reasonable time frame is usually not feasible.
+The approach for fast computations taken by the Valhalla routing engine introduces some approximations which do not lead to exact alignment of empirical trajectories with the road network or to each other (Sghaier, 2019).
+Nonetheless the misalignment is reduced to less than 5 m, which is, for many purposes, including most notably navigation, indeed a suitable result.
+For these inexactly aligned trajectories, applying the state-of-the-art traffic flow aggregation API (Morgan and Lovelace, 2021) leads to an inaccurate flow map.
+The authors resolve this by rasterising the flow map to a raster grid: and within each raster grid cell, the flow is the sum of the flows of all trajectories which intersect the grid cell.
+This rasterised flow map is displayed in Figure 1(c).
+The colour scale for the flow ranges from 100 (violet) to 300 (pink) to 600 (orange).
+The 2 m $\times$ 2 m raster grid obscures the road network at this scale.
+The resolution of a rasterised aggregation flow map is fixed by the resolution of the raster grid.
+If we require a higher resolution, then we must recalculate the flow map on a raster grid with a higher resolution.
+Our goal is to produce a traffic flow map from map matched routes, which can be utilised at any scale without recalculation, from the city level to the individual road level.
+In our proposed local alignment approach, we infer which road segments should serve as a local reference, and then proceed to align other nearby road segments to it.
 To accomplish this, we introduce several novel algorithms which employ a mix of advanced statistical and geospatial methods.
-These include ‘node snapping’ where the nearby boundary points of road segments are combined via statistical clustering, and ‘line blending’ where road segments, which are near to each other but do not share overlapping road sub-segments, are aligned to maximise their overlapping sub-segments.
-Inputting these locally aligned routes into a flow aggregation API leads to a more accurate flow map.
-Iterating these local alignments in turn leads to a minimal flow map.
+These include “node snapping” where the nearby boundary points of road segments are combined via statistical clustering, and “line blending” where road segments, which are near each other but do not share overlapping sub-segments, are aligned to maximise their overlapping sub-segments.
+We combine these novel algorithms into an iterative workflow which outputs a locally aligned flow map.
+This flow map is composed of road segments rather than raster grid cells, so its resolution is not limited by the raster grid resolution.
 The outline of the paper is as follows.
-In Section 2, we describe the computation of map matched routes using off-the-shelf map matching and route finding APIs.
-In Section 3, we describe the local alignment of the map matched routes and their aggregation into a flow map.
-In Section 4, we demonstrate that the locally aligned flow map from the Hannover trajectories is well-aligned to the OSM road network, and has a high level of accuracy and spatial coverage of estimated traffic flows compared to reference traffic flows.
-We then discuss some software implementation issues, and some concluding remarks.
-2 Alignment of trajectories with map matching and route finding
-We introduce some mathematical notation to state precisely the problem of map matching.
-We represent the road network by a graph $\mathcal{N}=\mathcal{N}(\mathcal{E},\mathcal{V})$, where the $n_{\mathcal{E}}$ edges =$\mathcal{E}=\{{\bm{e}}_{1},\dots,{\bm{e}}_{n_{\mathcal{E}}}\}$ of this graph are the road segments and the $n_{\mathcal{V}}$ nodes/vertices $\mathcal{V}=\{{\bm{v}}_{1},\dots,{\bm{v}}_{n_{\mathcal{V}}}\}$ indicate that two (or more) different road segments are accessible to/from each other at this node point.
-These nodes ${\bm{v}}_{i}$ are single points.
-Each road segment is composed of a sequence of connected piecewise linear segments, so ${\bm{e}}_{i}=\{{\bm{e}}_{i,1},\dots,{\bm{e}}_{i,n_{i}}\}$ is an ordered sequence of $n_{i}$ points ${\bm{e}}_{i,j},j=1,\dots,n_{i}$.
-We refer to this as a ‘linestring’ geometry, following the Open Geospatial Consortium terminology (OGC, 2010).
-We set $\mathcal{N}$ to be the OSM network which is a freely available road network with global coverage.
+In Section 2, we construct an iterative workflow which computes a locally aligned flow map from noisy trajectories.
+In Section 3, we apply this workflow to synthetic trajectories from Los Angeles, USA and empirical trajectories from Hannover, Germany.
+We verify that, with a suitable tuning parameter choice, the locally aligned flow maps have high resolution and spatial accuracy.
+We then discuss some computational issues and end with some concluding remarks.
+2 Local alignment of route segments for flow aggregation
+We represent the road network by a graph $\mathcal{N}=\mathcal{N}(\mathcal{E},\mathcal{V})$, where the $n_{\mathcal{E}}$ edges =$\mathcal{E}=\{{\bm{e}}_{1},\dots,{\bm{e}}_{n_{\mathcal{E}}}\}$ are the road segments and the $n_{\mathcal{V}}$ nodes/vertices $\mathcal{V}=\{{\bm{v}}_{1},\dots,{\bm{v}}_{n_{\mathcal{V}}}\}$ indicate that different road segments are accessible to/from each other at this node point.
+We set $\mathcal{N}$ to be the OpenStreetMap (OSM) road network.
+Each path in the network is composed of a connected sequence of edges, and we refer to this as a “linestring” geometry, following the Open Geospatial Consortium (OGC) terminology (OGC, 2010).
 We denote a single trajectory $G=\{{\bm{g}}_{1},\dots,{\bm{g}}_{n_{G}}\}$ as a temporally ordered sequence of $n_{G}$ points ${\bm{g}}_{i},i=1,\dots,n_{G}$.
-This is known as a ‘multipoint’ geometry (OGC, 2010).
-Whilst some authors require that each point ${\bm{g}}_{i}$ be accompanied by their timestamp $t_{i}$ to be considered a trajectory, this is not strictly required since ${\bm{g}}_{i}$ are ordered according to the timestamps, even if the timestamps themselves are not recorded in the trajectory.
-Due to measurement error, the points of a trajectory $G$ are not necessarily coincident with the road network $\mathcal{N}$, and due to sampling error, there is no information about the vehicle in between the points of $G$.
-We represent the output of a map matching algorithm $M$ from an empirical trajectory $G$ as $M(G)=\{{\bm{m}}_{1},\dots,{\bm{m}}_{n_{M}}\}$ which is an ordered, connected sequence of $n_{M}$ edges.
-The goal is that the map matched route follows closely the road network $\mathcal{N}(\mathcal{V},\mathcal{E})$.
-Whilst it is straightforward to ensure that all boundary points of the segments $\{{\bm{m}}_{1},\dots,{\bm{m}}_{n_{M}}\}$ coincide with the nodes $\mathcal{V}$ of the road network graph $\mathcal{N}$, it is more challenging to ensure that the segments $\{{\bm{m}}_{1},\dots,{\bm{m}}_{n_{M}}\}$ themselves coincide with the edges $\mathcal{E}$.
-In comparison to an empirical trajectory $G$, all boundary points of the edges of a map matched route $M(G)$ are aligned to the road network (reduced measurement error) and the vehicle position is estimated by the linestring connecting the boundary points of the road segment (reduced sampling error).
+This is a “multipoint” geometry (OGC, 2010).
+The points of a trajectory $G$ are not necessarily coincident with the road network $\mathcal{N}$.
+2.1 Map matching
+We represent the output of a map matching algorithm $M$ from a trajectory $G$ as $M(G)=\{{\bm{m}}_{1},\dots,{\bm{m}}_{n_{M}}\}$ which is a connected sequence of $n_{M}$ edges, i.e.
+a linestring geometry.
+In comparison to a trajectory $G$, all boundary points of the edges of a map matched route $M(G)$ are aligned to the road network.
 There is a large body of research on this difficult problem of map matching.
@@
 According to a review of map matching algorithms (Chao et al., 2020), HMM is a state-transition method.
-The other three classes are similarity, candidate-evolving and scoring methods.
-Further details of alternative map matching algorithms are found in Quddus et al.
+The other three classes are similarity, candidate-evolving, and scoring methods.
+Further details of these alternatives are found in Quddus et al.
 (2007), Chao et al.
 (2020).
-We leave this discussion here since the improvements offered by our proposed methods are valid for any map matching algorithm, and concentrate on HMM map matching due to its accuracy and computational efficiency.
+We leave this discussion here since our proposed methods are valid for any map matching algorithm, and concentrate on HMM map matching due to its accuracy and computational efficiency.
 Even if we restrict ourselves to HMM algorithms on the OSM road network, there are many off-the-shelf map matching APIs available.
 We focus on the Valhalla routing engine (https://valhalla.github.io/valhalla), which includes its highly recommended map matching API (Saki and Hagen, 2022).
-Figure 2 displays the $n=1147$ map matched routes by the Valhalla map matching API.
-We discard 36 trajectories (3.04%) from our original data set of 1183 trajectories.
-In Figure 2(a), the map matched routes (blue lines) overall are well-aligned to the road network.
-For the orange map matched route, all its segments are aligned to the road network, whereas the purple route appears to be displaced by several metres from the road centreline.
-The measurement and sampling errors of the map matched routes $M(G)$ are reduced in comparison those for the empirical trajectories $G$, though these errors remain sizeable at the road segment level in Figure 2(b).
-Refer to caption(a) Refer to caption(b)
-Figure 2: Map matched routes, using only Valhalla map matching API.
-(a) Neighbourhood level.
-(b) Small neighbourhood level, zoom of black rectangle.
-Trajectory ID = 7 (orange), 315 (purple), others (blue).
-Our first contribution is to better align the edges of the map matched routes $M(G)$ to the road network edges $\mathcal{E}$.
-We propose a post hoc adjustment of the map matching output by an additional call to a route finding API, as outlined in Algorithm 1.
-The inputs of ST_ROUTE are the empirical trajectory $G$, the map matching API $M$, the route finding API $R$, and the number of waypoints ${\bm{n}}_{W}$ for the route finding.
+To this off-the-shelf Valhalla map matching, we apply a post hoc modification to improve its suitability as input for flow map aggregation, as outlined in Algorithm 1.
+The inputs of ST_ROUTE are the trajectory $G$, the map matching API $M$, the route finding API $R$, and the number of waypoints ${\bm{n}}_{W}$ for the route finding.
+These waypoints guide the route finding API to pass near them.
 Since we do not have an a priori single optimal value for the number of waypoints, we consider a range of $w$ values ${\bm{n}}_{W}=(n_{W,1},\dots,n_{W,w})$.
-In Step 1, we compute the initial map matched route $M(G)$ from the empirical trajectory $G$ by calling the map matching API $M$.
+In Step 1, we compute the initial map matched route $M(G)$ from the trajectory $G$ by calling the map matching API $M$.
 This initial map matched route $M(G)=\{{\bm{m}}_{1},\dots,{\bm{m}}_{n_{M}}\}$ is a linestring with $n_{M}$ edges, with $n_{M}+1$ points.
-In Steps 2–5, we loop over the $w$ number of waypoints in ${\bm{n}}_{W}$.
+In Steps 2–5, we loop over the number of waypoints in ${\bm{n}}_{W}$.
 In Steps 3–4, we take a sample of $n_{W,i}$ waypoints, where the first waypoint ${\bm{w}}_{1}$ is the start point of ${\bm{m}}_{1}$ and the $n_{W,i}$th waypoint is the end point of ${\bm{m}}_{n_{M}}$, and the intermediate waypoints ${\bm{w}}_{2},\dots,{\bm{w}}_{n_{W,i}-1}$ are sampled from the start points of the edges ${\bm{m}}_{2},\dots,{\bm{m}}_{n_{M}}$.
 In Step 5, we call the route finding API $R$ with the waypoints ${\bm{w}}_{1},\dots,{\bm{w}}_{n_{W,i}}$.
-The result is the map matched route $M_{i}^{*}(G)=R({\bm{w}}_{1},\dots,{\bm{w}}_{n_{W,i}})=\{{\bm{m}}_{1}^{*},\dots,{\bm{m}}_{n_{M^{*}}}\}$, which is a route of $n_{M^{*}}$ connected edges.
-In Step 6, we select the route with the smallest dynamic time warping (DTW) normalised distance between the routes $M_{i}^{*}(G)$ and the empirical trajectory $G$, for $i=1,\dots w$.
-The DTW distance is based on the lengths of all the distortions of $M_{i}^{*}(G)$ to achieve a maximal alignment between $M_{i}^{*}(G)$ and $G$ (Sakoe and Chiba, 1978, Giorgino, 2009).
+The result is the map matched route $R_{i}(G)=R({\bm{w}}_{1},\dots,{\bm{w}}_{n_{W,i}})$.
+In Step 6, we select the route with the smallest dynamic time warping (DTW) normalised distance between the routes $R_{i}(G)$ and the trajectory $G$, for $i=1,\dots w$.
+The DTW distance is based on the lengths of all the distortions of $R_{i}(G)$ to achieve maximal alignment between $R_{i}(G)$ and $G$ (Sakoe and Chiba, 1978, Giorgino, 2009).
 Algorithm 1 ST_ROUTE – Map matched route
 1:Input: $G$ trajectory, $M$ map matching API, $R$ route finding API, ${\bm{n}}_{W}$waypoints
-2:Output: $M^{*}(G)$ map matched route
+2:Output: $R(G)$ map matched route
 3:Compute initial map matched route $M(G)$ from empirical trajectory $G$
@@
 5:     Set $n_{W}:=n_{W,i}$
-6:     Sample $n_{W}$ waypoints ${\bm{w}}_{1},\dots,{\bm{w}}_{n_{W}}$ from $M(G)$, with ${\bm{w}}_{1}$ := $\operatorname{Start}(M)$, ${\bm{w}}_{n_{W}}:=\operatorname{End}(M)$
-7:     Compute map matched route $M^{*}_{i}(G):=R({\bm{w}}_{1},\dots,{\bm{w}}_{n_{W}})$
-8:Select minimal route $M^{*}(G):=\operatorname{argmin}_{i\in\{1\dots w\}}\,\operatorname{DTW}(M^{*}_{i}(G),G)$
+6:     Sample waypoints ${\bm{w}}_{1},\dots,{\bm{w}}_{n_{W}}$ from initial map matched $M(G)$
+7:     Compute map matched routes $R_{i}(G):=R({\bm{w}}_{1},\dots,{\bm{w}}_{n_{W}})$
+8:Select minimal route $R(G):=\operatorname{argmin}_{i\in\{1,\dots,w\}}\,\operatorname{DTW}(R_{i}(G),G)$
 We set $R$ to be the Valhalla Odin turn-by-turn route finding API to be consistent with our choice of $M$ as the Valhalla Meili map matching API.
-We set the number of waypoints as ${\bm{n}}_{W}=3,13,23,33,43,63,83$.
-Figure 3 displays the results $M^{*}(G_{1}),\dots,M^{*}(G_{1147})$ from ST_ROUTE.
-For the trajectory ID = 7 (orange) $n_{W}=83$, and ID = 315 (purple) $n_{W}=23$ give the minimal DTW route.
-The overall impression of the map matched routes $M^{*}$ in Figure 3(a) is that the misalignment has been reduced, especially in the purple line since it now aligns more closely to the road centreline.
-In Figure 3(b), at the level of road segments, whilst the map matched routes $M^{*}$ tend to be contained inside the road segments, they are not exactly coincident with each other.
-This is in part because the road network graph $\mathcal{N}$ itself contains small misalignments, and so they are propagated into any map matching or route finding algorithm based on it.
-These small alignments lead to a lack of overlapping sub-segments, which in turn lead to inaccurate flow aggregation.
-Refer to caption(a) Refer to caption(b)
-Figure 3: Map matched routes, using Valhalla map matching and route finding APIs.
-(a) Neighbourhood level.
-(b) Small neighbourhood level, zoom of black rectangle.
-Trajectory ID = 7 (orange), 315 (purple), others (blue).
-3 Local alignment of road segments for flow map aggregation
-Our goal is to resolve the remaining misalignments from the map matching/route finding in the previous section, so we are able to aggregate accurately traffic flows on road segments.
-We aim to achieve this by local, internal alignment between the map matched routes.
-By internal alignment, we mean that we align the routes with each other, rather than to the external road network graph.
-Since an external reference road network is not required as an explicit input, our proposal can be deployed in more cases, e.g.
-when the quality of the road network graph is insufficient, or when the alignment to the road network graph is computationally intensive.
-By local alignment, we mean that we focus on aligning sub-segments of the routes, rather than complete routes.
-[…]
-We begin by illustrating the difference in flow aggregation between the $M$ and $M^{*}$ map matched routes.
-For the routes $M$ from Figure 2(b) and $M^{*}$ routes from Figure 3(b), the flow maps are given below in Figure 4(a–b) respectively.
-In these maps, the colour (purple to orange) and width of the road segments is proportional to the traffic flow.
-We observe that there are fewer, wider linestring segments in Figure 4(b) than in Figure 4(a).
-Refer to caption(a) Refer to caption(b)
-Figure 4: Flow maps for map matched routes.
-(a) Map matched routes with Valhalla map matching only.
-(b) Map matched routes with Valhalla map matching and route finding.
-Colour (purple to orange) and width of road segments is proportional to traffic flow.
-The flow aggregation in Figure 4 was carried out using the overline function in the R package stplanr, which we refer to as ST_OVERLINE_PLANR (Lovelace and Ellison, 2018).
-Starting with the map matched routes =$\mathcal{M}^{*}=\{M^{*}({G_{1}}),\dots,M^{*}(G_{n})\}$, then the flow map is $\mathcal{F}={\tt ST\_OVERLINE\_PLANR}(\mathcal{M}^{*})$.
-This flow aggregation involves the search for all road segments from the routes which exactly equal to each other.
-These exactly equal road segments are reduced to a single common segment, and the associated traffic flow is the number of exactly equal segments.
-Since it relies on exactly equal road segments, then small misalignments are sufficient to make the flow aggregations inaccurate.
-Our goal of producing a minimal flow map relies on resolving the crucial problem of how to aggregate similar, but not exactly overlapping, road segments.
+We set the number of waypoints as ${\bm{n}}_{W}=3,13,23,33,43,63,83$ since these ${\bm{n}}_{W}$ cover a range of possible values which likely includes an approximately optimal value.
+The map matched Hannover trajectories are displayed as the blue lines in Figure 1.
+2.2 Flow aggregation
+To the road network graph $\mathcal{N}=\mathcal{N}(\mathcal{E},\mathcal{V})$, we add the traffic flows on each of the network edges to form a flow map =$\mathcal{F}=\{(f,{\bm{f}}):{\bm{f}}\in\mathcal{N},f>0\}=\{(f_{1},{\bm{f}}_{1}),\dots,(f_{n_{\mathcal{F}}},{\bm{f}}_{n_{\mathcal{F}}})\}$ composed of $n_{\mathcal{F}}$ flows, where each ${\bm{f}}_{i}$ is composed of edges in $\mathcal{E}$, with traffic flow $f_{i}$, $i=1,\dots,n_{\mathcal{F}}$.
+Our goal is to estimate the traffic flows $f_{1},\dots,f_{n_{\mathcal{F}}}$ and the flow network graph $\mathcal{F}$.
+For exactly aligned trajectories, to compute the flow aggregation, we deploy the overline function in the R package stplanr, which we refer to as ST_OVERLINE_PLANR (Lovelace and Ellison, 2018).
+That is, starting with the map matched routes =$\mathcal{R}=\{R(G_{1}),\dots,R(G_{n})\}$, the flow map is $\mathcal{F}={\tt ST\_OVERLINE\_PLANR}(\mathcal{R})$.
+To compute an accurate flow map from noisy, empirical trajectories relies on resolving the crucial problem of how to aggregate similar, but not exactly overlapping, road segments.
 Many solutions have been offered, such as edge bundling (Zhou et al., 2013) and rasterisation (Wood et al., 2010, Morgan and Lovelace, 2021).
 Edge bundling consists of clustering trajectory linestrings and replacing all cluster members with a single representative linestring.
-These (and subsequent) authors conclude that it performs poorly when applied to noisy trajectories at the road segment level, and remains mostly suited to coarser aggregations, such as desire lines.
-Rasterisation relies on converting the vectorial flow map into a raster matrix, and aggregating the flows within the same raster pixel neighbourhood.
-Whilst this is indeed able to improve flow aggregations, it depends highly on the raster pixel neighbourhood size, and the rasterisation of the vectorial flow map leads to a loss of spatial resolution.
-We propose an alternative aggregation which does not lose resolution.
-Due to the complexity of this aggregation, it is divided into several algorithms, so that after the application of each algorithm, we progress further towards a minimal flow map.
-We develop our novel algorithms within the R statistical analysis environment, to take advantage of its integrated access to advanced statistical and geospatial analysis methods.
-Whilst R is not a bona fide GIS (Geographical Informations System), its geospatial functionalities conform to the OGC standards (OGC, 2010) via the package sf (Pebesma, 2018), and is a viable option for research in transport geospatial data analysis (Necula, 2015, Lovelace et al., 2019).
-3.1 Node snapping with hierarchical clustering
+These (and subsequent) authors conclude that it performs poorly when applied to noisy trajectories at the road level, and remains mostly suited to large-scale aggregations, such as desire line maps (Tobler, 1987).
+We do not consider edge bundling further since it is not suitable for our purposes.
+Rasterisation relies on converting the flow map into a raster matrix, and aggregating the flows within the same raster pixel neighbourhood.
+Whilst rasterisation is able to improve flow aggregation, it depends highly on the raster pixel neighbourhood size, and leads to a loss of spatial resolution.
+We propose an alternative aggregation which solves both of these problems.
+We achieve this goal by local internal alignment between the map matched routes.
+By internal alignment, we mean that we align the routes with each other, rather than to the external road network graph.
+By local alignment, we mean that we align segments of the routes, rather than the complete routes.
+Our proposal is divided into several algorithms, which we develop within the R statistical analysis environment due to its integrated access to advanced statistical and geospatial analysis methods.
+2.3 Node snapping and node splitting
 The proposed algorithm is ST_SNAPNODE, where the boundary points of the traffic flow linestrings are snapped to each other.
@@
 Since we are searching for points which are close to together, then this is well-suited to statistical clustering.
-There are many statistical clustering algorithms available, and we focus on hierarchical clustering (Gordon, 1999).
-A naive implementation where we consider all boundary points of all flow linestrings in a 1-pass complete linkage clustering is computationally intensive for any reasonable number of routes (Müllner, 2013).
-To resolve this computational bottleneck, we approximate the 1-pass complete linkage clustering by a nested 2-pass clustering.
-We begin with an efficient single linkage clustering of the boundary points of all linestrings in the R package fastcluster (Müllner, 2013).
+There are many statistical clustering algorithms available, and we focus on hierarchical clustering.
+To resolve the computational bottleneck of hierarchical clustering, we approximate the 1-pass complete linkage clustering by a nested 2-pass clustering.
+We begin with an efficient single linkage clustering in the R package fastcluster (Müllner, 2013).
 Since single linkage can result in chain-like clusters, we compute a subsequent complete linkage clustering to break these potential chains.
-In this nested approach, the complete linkage distance matrix is calculated only within each single linkage cluster, and so we are less likely to reach computational limits.
+In this nested approach, the complete linkage distance matrix is calculated only within each single linkage cluster, and so it is less likely to reach computational limits.
 Algorithm 2 is a description of ST_SNAPNODE.
-The inputs are the $n_{\mathcal{F}}$ traffic flow linestrings =$\mathcal{F}=\{{\bm{f}}_{1},\dots,{\bm{f}}_{n_{\mathcal{F}}}\}$ and the snap tolerance $\varepsilon_{S}$.
+The inputs are the $n_{\mathcal{F}}$ flow linestrings =$\mathcal{F}=\{(f_{1},{\bm{f}}_{1}),\dots,(f_{n_{\mathcal{F}}},{\bm{f}}_{n_{\mathcal{F}}})\}$ and the snap tolerance $\varepsilon_{S}$.
+As its name suggests, all node points within $\varepsilon_{S}$ distance of each other are snapped together.
 In Step 1, we extract the boundary points of all flow linestrings.
@@
 This divides each single linkage cluster into $C^{\prime\prime}$ clusters, where all cluster members are at most $\varepsilon_{S}$ distance from each other.
-In Steps 8–11, we compute the weighted spatial centroid in each of the $C^{\prime\prime}$ complete linkage clusters, weighted by the corresponding traffic flow values.
+In Steps 8–11, we compute the weighted spatial centroid in each of the $C^{\prime\prime}$ complete linkage clusters, weighted by the traffic flow values.
 In Step 12, we renumber the cluster labels from the above nested clusterings to approximate a 1-pass complete linkage clustering into $C^{*}$ clusters.
 In Steps 13–16, within each of these $C^{*}$ clusters, for all points of the corresponding flow linestrings which are closer than $\varepsilon_{S}$ distance to the centroid, we snap them to the centroid.
-In Steps 17–18, we collate and sort all the $n_{\mathcal{F}^{*}}$ snapped linestrings =$\mathcal{F}^{*}=\{{\bm{f}}_{1}^{*},\dots,{\bm{f}}_{n_{\mathcal{F}^{*}}}^{*}\}$.
+In Steps 17–18, we collate and sort all the $n_{\mathcal{F}^{*}}$ snapped linestrings =$\mathcal{F}^{*}=\{(f_{1}^{*},{\bm{f}}_{1}^{*}),\dots,(f_{n_{\mathcal{F}^{*}}}^{*},{\bm{f}}_{n_{\mathcal{F}^{*}}}^{*})\}$.
 In comparison to the unsnapped linestrings $\mathcal{F}$, there are fewer snapped linestrings $\mathcal{F}^{*}$ which have higher flow values, are more interconnected, and share more exactly overlapping sub-segments.
@@
 2:Output: $\mathcal{F}^{*}$ snap node clustered flow linestrings
-3:Extract boundary points $B(\mathcal{F}):=\{\operatorname{Start}({\bm{e}}_{1}),\operatorname{End}({\bm{e}}_{1}),\dots,\operatorname{Start}({\bm{e}}_{n_{\mathcal{F}}}),\operatorname{End}({\bm{e}}_{n_{\mathcal{F}}})\}$
+3:Extract boundary points $B(\mathcal{F}):=\{\operatorname{Start}({\bm{f}}_{1}),\operatorname{End}({\bm{f}}_{1}),\dots,\operatorname{Start}({\bm{f}}_{n_{\mathcal{F}}}),\operatorname{End}({\bm{f}}_{n_{\mathcal{F}}})\}$
 4:Compute hierarchical clustering with single linkage on $B(\mathcal{F})$
@@
 15:for $i:=1$ to $C^{*}$ do
-16:     Extract corresponding flow linestrings $\mathcal{F}_{i}^{*}:=\{{\bm{f}}_{1}^{*},\dots,{\bm{f}}_{n_{B^{*}}}^{*}\}$ from $\mathcal{F}$
-17:     Snap points of $\{{\bm{f}}_{1}^{*},\dots,{\bm{f}}_{n_{B^{*}}}^{*}\}$ within $\varepsilon_{S}$ distance of cluster centroid ${\bm{b}}^{*}_{i}$ to ${\bm{b}}^{*}_{i}$
+16:     Extract corresponding flow linestrings $\mathcal{F}_{i}^{*}:=\{(f_{1}^{*},{\bm{f}}_{1}^{*}),\dots,(f_{n_{B^{*}}}^{*},{\bm{f}}_{n_{B^{*}}}^{*})\}$ from $\mathcal{F}$
+17:     Snap points of $\{{\bm{f}}_{1}^{*},\dots,{\bm{f}}_{n_{B^{*}}}^{*}\}$ within $\varepsilon_{S}$ distance of cluster centroid to ${\bm{b}}^{*}_{i}$
 18:     $\mathcal{F}_{i}^{*}$ := rejoin snapped and unsnapped points into linestrings
-19:Collate snapped linestrings $\{\mathcal{F}_{1}^{*},\dots\mathcal{F}_{C^{\prime*}}^{*}\}$ into :=$\mathcal{F}^{*}:=\{{\bm{f}}_{1}^{*},\dots,{\bm{f}}_{n_{\mathcal{F}^{*}}}^{*}\}$
+19:Collate snapped linestrings $\{\mathcal{F}_{1}^{*},\dots\mathcal{F}_{C^{\prime*}}^{*}\}$ into :=$\mathcal{F}^{*}:=\{(f_{1}^{*},{\bm{f}}_{1}^{*}),\dots,(f_{n_{\mathcal{F}^{*}}}^{*},{\bm{f}}_{n_{\mathcal{F}^{*}}}^{*})\}$
 20:Sort $\mathcal{F}^{*}$ by descending flow and length
-An illustration of Algorithm 2 is given in Figure 5.
-In Figure 5(a) is the flow map before node snapping, where we observe that the boundary points (black solid circles) of some of the purple linestrings are closer to each than the snapping tolerance $\varepsilon_{S}=4$ m.
-Figure 5(b) shows an update flow map after node snapping, and applying ST_OVERLINE_PLANR.
-There are fewer thin purple road segments, since by snapping their intersection nodes, they share more exact sub-segments, and these shared sub-segments are merged during the flow aggregation to result in wide orange road segments.
-Refer to caption(a) Refer to caption(b)
-Figure 5: Flow maps for unsnapped and node snapped segments, with tolerance $\varepsilon_{S}=4~{}\mathrm{m}$.
-(a) Unsnapped.
-(b) Node snapped.
-Colour (purple to orange) and width of road segments is proportional to traffic flow.
-3.2 Node splitting to add missing intersection nodes
-Node snapping is focused near the boundary points of the flow linestrings.
-So it does not address misalignments far from the boundary points.
-Moreover there remain some linestrings which do indeed intersect but whose intersection is not correctly computed by ST_OVERLINE_PLANR.
+Node snapping is focused near the boundary points of the flow linestrings, so it does not address misalignments far from the boundary points.
+Moreover there remain some linestrings whose intersection is not correctly computed by ST_OVERLINE_PLANR.
 The solution to both of these problems is the explicit addition of the missing intersection nodes to these linestrings.
@@
 This gives the name to this method.
-The inputs to Algorithm 3 are the linestrings $\mathcal{F}$ and the node splitting $S$.
-For computational stability and efficiency, our method draws on two existing algorithms: in Steps 1–3, to_spatial_subdivision ($S=$ ‘subdivision’) in the sfnetworks package (van der Meer et al., 2023), and in Steps 4–5, geos_unary_union ($S=$ ‘unary’) in the geos package (Dunnington and Pebesma, 2023).
-The first option tends to find fewer intersection nodes, but this can be an advantage for our proposed line blending (to be introduced in the next subsection) since many small linestrings with similar flow values do not provide a clear prioritisation of this blending.
-It is also similar to the intersection computation performed by ST_OVERLINE_PLANR.
+The inputs to Algorithm 3 are the flow linestrings $\mathcal{F}$ and the node splitting type $S$.
+Our method deploys for $S=$ “subdivision” in Steps 1–3, to_spatial_subdivision in the sfnetworks package (van der Meer et al., 2023), and for $S=$ “unary” in Steps 4–5, geos_unary_union in the geos package (Dunnington and Pebesma, 2023).
+The first option tends to find fewer intersection nodes and is similar to the intersection computation performed by ST_OVERLINE_PLANR.
 The second option finds more missing intersection nodes, and leads to more comprehensive flow aggregation.
-We require both types of node splitting to compute a minimal flow map.
+We require both types of node splitting to compute a locally aligned flow map.
 Algorithm 3 ST_SPLITNODE – Split nodes at interior intersections of linestrings
@@
 3:Initialise local network $\mathcal{N}^{*}$ with linestrings $\mathcal{F}$
-4:if $S==$ ‘subdivision’ then
+4:if $S==$ “subdivision” then
 5:     $\mathcal{F}^{*}:=$ sfnetworks::to_spatial_subdivision$(\mathcal{N}^{*})$
-6:else if $S==$ ‘unary’ then
+6:else if $S==$ “unary” then
 7:     $\mathcal{F}^{*}:=$ geos::geos_unary_union$(\mathcal{N}^{*})$
-In Figure 6(a) is the flow map without any node splitting or node snapping.
-This map has missing intersection nodes, and intersections nodes which are close to each other.
-In Figure 6(b) is the flow map after node splitting ($S=$ ‘unary’) to add the missing intersections nodes, and then node snapping (with snap tolerance $\varepsilon_{S}=4$ m).
-The result is that there are fewer road segments with higher flow values.
-Due to the combined action of ST_SPLITNODE and ST_SNAPNODE, Figure 6(b) is an improvement over Figures 5(a–b) and 6(a).
-Refer to caption(a) Refer to caption(b)
-Figure 6: Flow maps for unsplit/unsnapped and node split/node snapped segments, with tolerance $\varepsilon_{S}=4~{}\mathrm{m}$.
-(a) Unsplit and unsnapped.
-(b) Node split and node snapped.
-Colour (purple to orange) and width of road segments is proportional to traffic flow.
+In Figure 2(a) are the flow segments without any node splitting or node snapping.
+There are missing intersection nodes, and there are intersection nodes which are close to each other.
+In Figure 2(b) are the flow segments after node splitting ($S=$ “unary”) and then node snapping (with snap tolerance $\varepsilon_{S}=4$ m).
+The result is that there are fewer flow segments with higher flow values.
+Due to the combined action of ST_SPLITNODE and ST_SNAPNODE, Figure 2(b) is an improvement over Figure 2(a).
+Refer to caption Refer to caption
+Figure 2: Unsplit/unsnapped and node split/node snapped flow segments, with snap tolerance $\varepsilon_{S}=4$ m. (a) Unsplit and unsnapped. (b) Node split and node snapped.
 2.4 Line blending to align similar linestrings with local reference
-So far we have focused on improving the alignment of linestrings induced by resolving inconsistencies at their intersections.
+So far we have focused on improving the alignment of linestrings by resolving inconsistencies at their intersections.
 We now focus on aligning linestrings more generally.
 For this, we require a comparison relationship to establish an order of alignment of nearby linestrings.
-For two linestrings ${\bm{f}}_{1}$ and ${\bm{f}}_{2}$ from a flow map $\mathcal{F}$, we define that ${\bm{f}}_{2}$ is a candidate to be aligned to the reference ${\bm{f}}_{1}$ at threshold $\varepsilon\geq 0$ if
+For two flow linestrings $(f_{1},{\bm{f}}_{1})$ and $(f_{2},{\bm{f}}_{2})$ from a flow map $\mathcal{F}$, we define that ${\bm{f}}_{2}$ is a candidate to be aligned to the reference ${\bm{f}}_{1}$ at threshold $\varepsilon\geq 0$ if
 ${\bm{f}}_{2}\subseteq{\tt ST\_BUFFER}({\bm{f}}_{1},\varepsilon),\ f_{2}\leq f_{1}.$ (1)
-This relation is insensitive to any local complexities in ${\bm{f}}_{2}$ as long as they are all contained within the buffer zone around ${\bm{f}}_{1}$.
-The buffer zone we employ has flat edges, e.g.
-for the R package sf, this corresponds to ST_BUFFER(endCapStyle="FLAT"), so the buffer zone ends at the boundary points of ${\bm{f}}_{1}$.
-This avoids erroneously considering neighbouring segments of ${\bm{f}}_{1}$, which are connected to ${\bm{f}}_{1}$ as a part of a longer sequence of road segments, to be candidates to be aligned to ${\bm{f}}_{1}$.
-The condition on the flow values means that we place a higher priority on linestrings with higher flows.
-Since we can define a local reference linestring, a global road network graph is no longer required to align the candidate linestrings.
+The condition on the flow values $f_{2}\leq f_{1}$ means that we place a higher priority on the reference linestring ${\bm{f}}_{1}$ since its flow $f_{1}$ is larger than the flow $f_{2}$ on the candidate linestring ${\bm{f}}_{2}$.
+The buffer zone has flat edges, e.g.
+ST_BUFFER(endCapStyle="FLAT") in the R package sf (Pebesma, 2018), so the buffer zone ends at the boundary points of ${\bm{f}}_{1}$.
+This avoids erroneously considering neighbouring segments of ${\bm{f}}_{1}$, which are connected to ${\bm{f}}_{1}$ as a part of a longer sequence, to be candidates to be aligned to ${\bm{f}}_{1}$.
 Let ${\bm{f}}_{\rm ref}$ be a reference linestring from $\mathcal{F}$.
-The set of $m$ linestrings from \$\mathcal{F}\backslash\{{\bm{f}}_{\rm ref}\}$ which satisfy Equation (1) are the candidate linestrings $\mathcal{F}_{\rm cand}=\{{\bm{f}}_{{\rm cand},1},\dots,{\bm{f}}_{{\rm cand},n_{\mathcal{F}_{\rm cand}}}\}$, with the convention that if $n_{\mathcal{F}_{\rm cand}}=0$ then $\mathcal{F}_{\rm cand}$ is the empty set.
-We call our approach ‘line blending’, since we will blend the candidate linestrings onto the reference linestring.
-The inputs in Algorithm 4 are the reference linestring ${\bm{f}}_{\rm ref}$, the set of $m$ candidate linestrings $\mathcal{F}_{\rm cand}$, and the blend tolerance $\varepsilon$.
-The output are the modified reference and $m$ modified candidate flow linestrings, all with added interior points for accurate calculation of exactly equal linestring segments.
-In Step 1, we initialise a local network graph $\mathcal{N}^{*}$ with the reference linestring.
+The set of $n_{\mathcal{F}_{\rm cand}}$ linestrings from \$\mathcal{F}\backslash\{{\bm{f}}_{\rm ref}\}$ which satisfy Equation (1) are the candidate linestrings $\mathcal{F}_{\rm cand}=\{{\bm{f}}_{{\rm cand},1},\dots,{\bm{f}}_{{\rm cand},n_{\mathcal{F}_{\rm cand}}}\}$, with the convention that if $n_{\mathcal{F}_{\rm cand}}=0$ then $\mathcal{F}_{\rm cand}$ is the empty set.
+We call our approach “line blending”, since we blend the candidate linestrings onto the reference linestring.
+The inputs in Algorithm 4 are the reference linestring ${\bm{f}}_{\rm ref}$, the $n_{\mathcal{F}_{\rm cand}}$ candidate linestrings $\mathcal{F}_{\rm cand}$, and the blend tolerance $\varepsilon$.
+The output are the modified reference linestring and $n_{\mathcal{F}_{\rm cand}}$ modified candidate flow linestrings, all with added interior points for accurate calculation of exactly equal linestring segments.
+In Step 1, we initialize a local network graph $\mathcal{N}^{*}$ with the reference linestring.
 In Steps 2–3, we extract all points of the candidate linestrings, and use the network_blend function in the sfnetworks package (which we denote as ST_NETWORK_BLEND) to blend efficiently these points into the reference linestring (van der Meer et al., 2023).
-This network blending requires a blend threshold, which we set to be $\varepsilon$.
 ST_NETWORK_BLEND projects the candidate linestrings onto the reference linestring, and explicitly adds them to the network, thereby creating new edges in $\mathcal{N}^{*}$.
 The result is a local network graph with more, shorter edges and with nodes at the projected candidate points, and whose union is the reference linestring.
-In Steps 4–5, we extract the $m$ blended candidate linestrings with these added interior points by applying the shortest path search between the start and end point of each blended candidate linestring along the network graph $\mathcal{N}^{*}$.
-Step 6 is the equivalent for the blended reference linestring.
-Step 7 involves collating the blended candidate linestrings.
+In Steps 4–5, we extract the $n_{\mathcal{F}_{\rm cand}}$ blended candidate linestrings with these added interior points by applying the shortest path search between the start and end point of each blended candidate linestring along the network graph $\mathcal{N}^{*}$.
+Step 6 involves collating the blended candidate linestrings.
+Step 7 is the equivalent for the blended reference linestring.
 Algorithm 4 ST_LINEBLEND – Blend candidate linestrings onto reference linestring
 1:Input: ${\bm{f}}_{\rm ref}$ reference, $\mathcal{F}_{\rm cand}$ candidate flow linestrings, $\varepsilon$ blend tolerance
-2:Output: $({\bm{f}}^{*}_{\rm ref},\mathcal{F}^{*}_{\rm cand})$ blended reference and $m$ candidate flow linestrings
+2:Output: ${\bm{f}}^{*}_{\rm ref}$ blended reference, $\mathcal{F}^{*}_{\rm cand}$ blended candidate flow linestrings
 3:Initialise local network graph $\mathcal{N}^{*}$ with ${\bm{f}}_{\rm ref}$
@@
 7:     ${\bm{f}}^{*}_{{\rm cand},i}$ := shortest path from $\operatorname{Start}({\bm{f}}_{{\rm cand},i})$ to $\operatorname{End}({\bm{f}}_{{\rm cand},i})$ along network $\mathcal{N}^{*}$
-8:${\bm{f}}^{*}_{{\rm ref}}$ := shortest path from $\operatorname{Start}({\bm{f}}_{{\rm ref}})$ to $\operatorname{End}({\bm{f}}_{{\rm ref}})$ along network $\mathcal{N}^{*}$
-9:Collate $\mathcal{F}^{*}_{\rm cand}:=\{{\bm{f}}_{{\rm cand},1}^{*},\dots,{\bm{f}}_{{\rm cand},n_{\mathcal{F}_{\rm cand}}}^{*}\}$
+8:Collate $\mathcal{F}^{*}_{\rm cand}:=\{{\bm{f}}_{{\rm cand},1}^{*},\dots,{\bm{f}}_{{\rm cand},n_{\mathcal{F}_{\rm cand}}}^{*}\}$
+9:${\bm{f}}^{*}_{{\rm ref}}$ := shortest path from $\operatorname{Start}({\bm{f}}_{{\rm ref}})$ to $\operatorname{End}({\bm{f}}_{{\rm ref}})$ along network $\mathcal{N}^{*}$
 The result from ${\tt ST\_LINEBLEND}$ is the modified reference linestring and the projected candidate linestrings with added interior points.
-These added interior points resolve a limitation of the flow aggregation of ST_OVERLINE_PLANR (Morgan and Lovelace, 2021).
 If there are many candidate linestrings, then this may lead to many sub-segments in the projected linestrings, each with their own flow values.
-So we assign the weighted mean flow, weighted by the sub-segment length, to all sub-segments.
-This single flow value takes into account the contribution of each candidate linestring to the flow along the reference linestring.
-Moreover, we take the rounded value of this weighted mean flow to expedite the aggregation computations and to reduce visual clutter of the flow map.
-In Figure 7(a), we illustrate this line blending with the linestrings $ABC$ with flow $f_{ABC}=7$ in blue, and $AD$ with flow $f_{AD}=2$ in orange.
+So we assign the rounded value of the weighted mean flow, weighted by the sub-segment length, to all sub-segments.
+In Figure 3(a), we illustrate this line blending with the flow linestrings $(7,ABC)$ in blue and $(2,AD)$ in orange.
 As these two linestrings do not share a common sub-segment, so flow aggregation does not modify them.
-Since $AD\subset{\tt ST\_BUFFER}(ABC,4~{}\mathrm{m})$ (the pale blue rectangle), and $f_{AD}<f_{ABC}$, then $ABC$ is the reference ${\bm{f}}_{\rm ref}$ and $AD$ is the candidate linestring ${\bm{f}}_{\rm cand}$, according to Equation (1).
+Since $AD\subset{\tt ST\_BUFFER}(ABC,4\,\mathrm{m})$ (the pale blue rectangle), and $2<7$, then $ABC$ is the reference ${\bm{f}}_{\rm ref}$ and $AD$ is the candidate linestring ${\bm{f}}_{\rm cand}$, according to Equation (1).
 Line blending involves projecting $AD$ to $ABC$, so $D$ is projected to $D^{\prime}$ (which lies exactly on the reference linestring), and $B$ is added explicitly to this projected linestring.
@@
 We take the rounded value of the weighted mean of these two flow linestrings.
-The result of ST_LINEBLEND with ST_OVERLINE_PLANR is a single linestring $(9,ABD^{\prime}C)$, as shown in Figure 7(b).
-Refer to caption(a) Refer to caption(b)
-Figure 7: Line blending removes small misalignments and leads to correct flow aggregation.
+The result is a single linestring $(9,ABD^{\prime}C)$, as shown in Figure 3(b).
+Refer to caption Refer to caption
+Figure 3: Line blending leads to correct flow aggregation.
 (a) Before line blending and flow aggregation.
@@
 Blended linestring is $(9,ABD^{\prime}C)$.
+After aggregation, the projected point $D^{\prime}$ in $ABD^{\prime}C$ appears to be redundant.
+Suppose that we have another linestring $(3,ABC)$ without $D^{\prime}$.
+Then ST_OVERLINE_PLANR does not aggregate $(9,ABD^{\prime}C)$ with $(3,ABC)$ even if $D^{\prime}$ is exactly coincident on the segment $BC$, since this algorithm does not explicitly add any missing intersection nodes prior to aggregation (Morgan and Lovelace, 2021).
+It outputs three linestrings $(12,AB),(9,BC),(3,BD^{\prime}C)$, rather than a single linestring $(12,ABD^{\prime}C)$.
+So we retain these intermediate points in the flow linestrings to assist in subsequent flow aggregation.
 More complex situations arise when other linestrings touch the candidate linestring, but are not candidates themselves for blending.
@@
 In Steps 1–8, we iterate over each candidate-touching linestring.
-In Steps 2–3, we compute the intersection points between the candidate-touching linestring and the boundary points of the candidate linestrings, and the respective distances.
+In Steps 2–3, we compute the intersections between the candidate-touching linestring and the boundary points of the candidate linestrings, and the respective distances.
 In Steps 4–7, if this intersection point is within $\varepsilon_{S}$ distance to the boundary points of ${\bm{f}}_{\rm ref}$, then we snap the candidate-touching linestring to the closest boundary point.
@@
 11:Collate $\mathcal{F}^{*}_{\rm ct}:=\{{\bm{f}}^{*}_{{\rm ct},1},\dots,{\bm{f}}^{*}_{{\rm ct},n_{\mathcal{F}_{\rm ct}}}\}$
-Figure 8(a) is an illustration of ST_SNAP_CAND_TOUCH with the reference $ABC$ with flow $f_{ABC}=7$ (blue), the candidate linestring $AD$ with flow $f_{AD}=2$ (orange), and the blending buffer zone with blend tolerance $\varepsilon=4$ m (light blue).
+Figure 4(a) is an illustration of ST_SNAP_CAND_TOUCH with the reference $(7,ABC)$ (blue), the candidate linestring $(2,AD)$ (orange), and the blending buffer zone with blend tolerance $\varepsilon=4$ m (light blue).
 The candidate-touching linestring ${\bm{f}}_{\rm ct}=(5,DE)$ (purple) touches the candidate $AD$ at $D$.
-The orange line is entirely within the pale blue buffer zone around the blue reference line, whereas the purple line extends outside of the buffer zone and so is not a candidate for blending to the reference linestring.
-If we apply ST_LINEBLEND to blend the candidate linestring, then $D$ is projected to $D^{\prime}$ to lie on the reference linestring, i.e.
-${\bm{f}}_{\rm cand}^{*}=(2,ABD^{\prime})$.
-If we apply ST_SNAP_CAND_TOUCH to the candidate-touching linestring, then as the intersection point $D$ is less than $\varepsilon_{S}=4$ m from the boundary points of ${\bm{f}}_{\rm ref}$, it is projected to the boundary point $C$, i.e.
-${\bm{f}}^{*}_{\rm ct}=(5,CE)$.
-Thus line blending, candidate-touching snapping and flow aggregation yields that ${\bm{f}}^{*}_{\rm ct}=(5,CE)$ remains connected to ${\bm{f}}_{\rm ref}^{*}=(9,ABD^{\prime}C)$.
-Refer to caption(a) Refer to caption(b)
+The orange line is entirely within the pale blue buffer zone, whereas the purple line extends outside of the buffer zone and so is not a candidate for blending to the reference linestring.
+If we apply ST_LINEBLEND to blend the candidate linestring, then $D$ is projected to $D^{\prime}$ to lie on the reference linestring, i.e., ${\bm{f}}_{\rm cand}^{*}=(2,ABD^{\prime})$.
+If we apply ST_SNAP_CAND_TOUCH with snap tolerance $\varepsilon_{S}=4$ m to the candidate-touching linestring, then the intersection point $D$ is projected to the boundary point $C$, i.e., ${\bm{f}}^{*}_{\rm ct}=(5,CE)$.
+Thus line blending, candidate-touching snapping, and flow aggregation yields that ${\bm{f}}^{*}_{\rm ct}=(5,CE)$ remains connected to ${\bm{f}}_{\rm ref}^{*}=(9,ABD^{\prime}C)$.
+Refer to caption Refer to caption
 Figure 4: Snap candidate-touching linestrings to reference linestring after line blending, with blend tolerance $\varepsilon=4$ m, snap tolerance $\varepsilon_{S}=4$ m.
@@
 Reference linestring becomes $(9,ABD^{\prime}C)$, and candidate-touching linestring $(5,CE)$.
-So far we have only considered line blending where the reference linestring has a single flow.
-Due to the noisiness of the empirical trajectories, this is insufficient to reach a minimal flow map.
-So we allow the reference to be a sequence of $k$ connected edges, each with their own flow.
+We generalise the reference linestring to be a sequence of $k$ connected edges, each with their own flow.
 If we treat this edge sequence momentarily as a single linestring, with flow equal to the weighted mean flow, weighted by the length of the $k$ edges, then we can apply Equation (1) to search for potential candidate linestrings.
@@
 The generalisation to $k$-edge reference linestrings allows us to blend candidate linestrings which exceed the buffer zones of 1-edge reference linestrings.
-We have described the situation for blending candidate and candidate-touching linestrings to a single reference linestring (possibly composed of $k$ edges).
-The next step to determine the priority for line blending for a set of reference linestrings.
+We have described blending candidate and candidate-touching linestrings to a single reference linestring (possibly composed of $k$ edges).
+The next step is to determine the priority for line blending for a set of reference linestrings.
 We require that a reference linestring cannot be a candidate linestring to another reference linestring, and a candidate linestring is a candidate for one reference linestring only.
 These ensure that if two linestrings ${\bm{f}}_{2}\subseteq{\tt ST\_BUFFER}({\bm{f}}_{1},\varepsilon)$ and ${\bm{f}}_{1}\subseteq{\tt ST\_BUFFER}({\bm{f}}_{2},\varepsilon)$, then we select only one of them.
-The flow condition $f_{2}\leq f_{1}$ usually distinguishes the reference linestring, except for where $f_{1}=f_{2}$.
-In the situation of equal flow values, then we designate the longer linestring to be the reference.
+The flow condition $f_{2}\leq f_{1}$ usually distinguishes the reference linestring, except when $f_{1}=f_{2}$.
+In this case, then we designate the longer linestring to be the reference.
 The inputs to ST_LINEBLEND_PRIORITY in Algorithm 6 are the flow linestrings $\mathcal{F}$, the number of connected edges for the reference linestrings $k$, and the blend tolerance $\varepsilon$.
-The output is a set of non-overlapping reference linestrings $\mathcal{F}_{\rm ref}$, a set of unique candidate linestrings $\mathcal{F}_{\rm cand}$ and a set of (potentially repeated) candidate-touching $\mathcal{F}_{\rm ct}$ linestrings.
-In Steps 1–2, we set up a network graph from the linestrings $\mathcal{F}$, and then extract $\mathcal{K}$, all simple paths (i.e.
-all paths composed of connected, unique edges) with $k$-edges.
+The output is a set of non-overlapping reference linestrings $\mathcal{F}_{\rm ref}$, a set of unique candidate linestrings $\mathcal{F}_{\rm cand}$, and a set of (potentially repeated) candidate-touching $\mathcal{F}_{\rm ct}$ linestrings.
+In Steps 1–2, we set up a network graph from the linestrings $\mathcal{F}$, and then extract $\mathcal{K}$, all simple paths (i.e., all paths composed of connected, unique edges) with $k$-edges.
 In Steps 3–4, we compute the weighted mean of the $k$ flow values, weighted by the length of individual edges, and assign this to the entire $k$-edge path.
-The function ${\rm Edges}(\cdot)$ extracts the edges from a $k$-edge path.
+The function ${\rm Edge}(\cdot)$ extracts the component edges from a $k$-edge path.
 In Step 5, we sort the flow linestrings, in descending order of their flow values and length.
@@
 In Step 13, we search for candidate linestrings for ${\bm{k}}_{i}^{*}$, from those linestrings which are not already reference or candidate linestrings, according to Equation (1).
-In Steps 14–17, if there is at least one candidate linestring, then we update $\mathcal{F}_{{\rm ref}},\mathcal{F}_{\rm cand}$ and $\mathcal{F}_{\rm ct}$.
-In Step 18, we extract the candidate $\mathcal{F}_{\rm cand}$ and candidate-touching $\mathcal{F}_{\rm ct}$ linestring sets, since the reference linestring set is already computed as $\mathcal{F}_{\rm ref}$.
+In Steps 14–17, if there is at least one candidate linestring, then we update $\mathcal{F}_{{\rm ref}},\mathcal{F}_{\rm cand}$, and $\mathcal{F}_{\rm ct}$.
+In Step 18, we extract the candidate $\mathcal{F}_{\rm cand}$ and candidate-touching $\mathcal{F}_{\rm ct}$ linestring sets.
 Algorithm 6 ST_LINEBLEND_PRIORITY – Compute line blending priority
@@
 5:for $i:=1$ to $n_{\mathcal{K}}$ do
-6:     $f_{i}$ := weighted mean of $k$ flows, weight := len(Edges(${\bm{k}}_{i}$))
+6:     $f_{i}$ := weighted mean of $k$ flows, weight := len(Edge(${\bm{k}}_{i}$))
 7:Sort $\mathcal{K}$ by descending flow and combined length
@@
 12:for $i:=1$ to $n_{\mathcal{K}^{*}}$ do
-13:     if (${\rm Edges}({\bm{k}}_{i}^{*})\notin\mathcal{F}_{\rm cand}$then
+13:     if (${\rm Edge}({\bm{k}}_{i}^{*})\notin\mathcal{F}_{\rm cand}$then
 14:         Set reference linestring ${\bm{k}}_{{\rm ref}}:={\bm{k}}_{i}^{*}$
-15:         Search candidate linestrings $\mathcal{F}_{{\rm cand}}^{*}:=\{{\bm{f}}\in\mathcal{F}\backslash({\rm Edges}(\mathcal{F}_{\rm ref})\cup\{{\bm{k}}_{\rm ref}\}\cup\mathcal{F}_{\rm cand}):{\bm{f}}\subseteq$
-16:    ${\tt ST\_BUFFER}({\bm{k}}_{\rm ref},\varepsilon),f\leq f_{\rm ref}\}$
-17:         if $\mathcal{F}_{\rm cand}^{*}\neq\{\}$ then
-18:              Update $\mathcal{F}_{\rm ref}:=\mathcal{F}_{\rm ref}\cup\{{\bm{k}}_{\rm ref}\},\mathcal{F}_{\rm cand}:=\mathcal{F}_{\rm cand}\cup\mathcal{F}_{\rm cand}^{*}$
-19:              Search touches $\mathcal{F}_{{\rm ct}}:=\{{\bm{f}}\in\mathcal{F}\backslash\{{\rm Edges}(\mathcal{F}_{\rm ref})\ \cup\mathcal{F}_{\rm cand}\}:{\tt ST\_TOUCHES}(\mathcal{F}_{{\rm cand}}^{*},{\bm{f}})\}$
-20:              Update $\mathcal{Q}:=\mathcal{Q}\cup\{(\{{\bm{k}}_{\rm ref}\},\mathcal{F}_{\rm cand}^{*},\mathcal{F}_{\rm ct}^{*})\}$
-21:Extract $\mathcal{F}_{\rm cand}:=\{\mathcal{F}^{*}_{{\rm cand},1},\dots,\mathcal{F}^{*}_{{\rm cand},n_{\mathcal{Q}}}\},\mathcal{F}_{\rm ct}:=\{\mathcal{F}^{*}_{{\rm ct},1},\dots,\mathcal{F}^{*}_{{\rm ct},n_{\mathcal{Q}}}\}$ from $\mathcal{Q}$
-The local alignment of road segments, in Algorithm 7, is established by combining Algorithms 26.
-Its inputs are the flow linestrings $\mathcal{F}$, the split node type $S$, the number of connected edges for the reference linestrings $k$, the blend tolerance $\varepsilon$ and the snap tolerance $\varepsilon_{S}$.
+15:         Search $\mathcal{F}_{{\rm cand}}^{*}:=\{{\bm{f}}\in\mathcal{F}\backslash({\rm Edge}(\mathcal{F}_{\rm ref})\cup\{{\bm{k}}_{\rm ref}\}\cup\mathcal{F}_{\rm cand})\colon{\bm{f}}\subseteq{\tt ST\_BUFFER}({\bm{k}}_{\rm ref},\varepsilon)\}$
+16:         if $\mathcal{F}_{\rm cand}^{*}\neq\{\}$ then
+17:              Update $\mathcal{F}_{\rm ref}:=\mathcal{F}_{\rm ref}\cup\{{\bm{k}}_{\rm ref}\},\mathcal{F}_{\rm cand}:=\mathcal{F}_{\rm cand}\cup\mathcal{F}_{\rm cand}^{*}$
+18:              Search $\mathcal{F}_{{\rm ct}}:=\{{\bm{f}}\in\mathcal{F}\backslash\{{\rm Edge}(\mathcal{F}_{\rm ref})\cup\mathcal{F}_{\rm cand}\}\colon{\tt ST\_TOUCHES}(\mathcal{F}_{{\rm cand}}^{*},{\bm{f}})\}$
+19:              Update $\mathcal{Q}:=\mathcal{Q}\cup\{(\{{\bm{k}}_{\rm ref}\},\mathcal{F}_{\rm cand}^{*},\mathcal{F}_{\rm ct}^{*})\}$
+20:Extract $\mathcal{F}_{\rm cand}:=\{\mathcal{F}^{*}_{{\rm cand},1},\dots,\mathcal{F}^{*}_{{\rm cand},n_{\mathcal{Q}}}\},\mathcal{F}_{\rm ct}:=\{\mathcal{F}^{*}_{{\rm ct},1},\dots,\mathcal{F}^{*}_{{\rm ct},n_{\mathcal{Q}}}\}$ from $\mathcal{Q}$
+The local alignment of flow segments, in Algorithm 7, is established by combining Algorithms 26.
+The inputs to ST_OVERLINE_LINEBLEND are the flow linestrings $\mathcal{F}$, the split node type $S$, the number of connected edges $k$, the blend tolerance $\varepsilon$, and the snap tolerance $\varepsilon_{S}$.
 In Step 1, we node split the flow linestrings $\mathcal{F}$.
-Step 2, we node snap the flow linestrings.
+In Step 2, we node snap the flow linestrings.
 In Step 3, we set up the priority for the line blending, and in Step 4 we store all the linestrings that will not be modified by the line blending in $\mathcal{F}^{c}$.
@@
 In Steps 10–11, we collate the modified linestrings from Steps 5–9, with the unmodified linestrings $\mathcal{F}^{c}$ from Step 4, and sort them in descending flow and length.
-Algorithm 7 ST_OVERLINE_LINEBLEND – Locally align road segment flows using line blending
+Algorithm 7 ST_OVERLINE_LINEBLEND – Locally align segment flows using line blending
 1:Input: $\mathcal{F}$ flow linestrings, $S$ split node, $k$edges, $\varepsilon$ blend tolerance, $\varepsilon_{S}$ snap tolerance
@@
 4:$\mathcal{F}:={\tt ST\_SNAPNODE}(\mathcal{F},\varepsilon_{S})$
-5:$(\mathcal{F}_{\rm ref},\mathcal{F}_{\rm cand},\mathcal{F}_{\rm ct}):={\tt ST\_LINEBLEND\_PRIORITY}(\mathcal{F},k,\varepsilon)$
-6:$\Edgessubscriptrefsubscriptcandsubscriptct\mathcal{F}^{c}:=\mathcal{F}\backslash({\rm Edges}(\mathcal{F}_{\rm ref})\cup\mathcal{F}_{\rm cand}\cup\mathcal{F}_{\rm ct})$
+5:$\{\mathcal{F}_{\rm ref},\mathcal{F}_{\rm cand},\mathcal{F}_{\rm ct}\}:={\tt ST\_LINEBLEND\_PRIORITY}(\mathcal{F},k,\varepsilon)$
+6:$\Edgesubscriptrefsubscriptcandsubscriptct\mathcal{F}^{c}:=\mathcal{F}\backslash({\rm Edge}(\mathcal{F}_{\rm ref})\cup\mathcal{F}_{\rm cand}\cup\mathcal{F}_{\rm ct})$
 7:for $i:=1$ to $n_{\mathcal{F}_{\rm ref}}$ do
-8:     $({\bm{f}}_{{\rm ref},i}^{*},\mathcal{F}_{{\rm cand},i}^{*}):={\tt ST\_LINEBLEND}({\bm{f}}_{{\rm ref},i},\mathcal{F}_{{\rm cand},i},\varepsilon)$
+8:     $\{{\bm{f}}_{{\rm ref},i}^{*},\mathcal{F}_{{\rm cand},i}^{*}\}:={\tt ST\_LINEBLEND}({\bm{f}}_{{\rm ref},i},\mathcal{F}_{{\rm cand},i},\varepsilon)$
 9:     ${\bm{f}}_{{\rm ref},i}^{*}:={\tt ST\_OVERLINE\_PLANR}({\bm{f}}_{{\rm ref},i}^{*}\cup\mathcal{F}_{{\rm cand},i}^{*})$
-10:     $f^{*}_{{\rm ref},i}$ := weighted mean of ${\bm{f}}^{*}_{{\rm ref},i}$, weight $:=\mathrm{len}({\rm Edges}({\bm{f}}_{{\rm ref},i}^{*}))$
+10:     $f^{*}_{{\rm ref},i}$ := weighted mean of ${\bm{f}}^{*}_{{\rm ref},i}$, weight $:=\mathrm{len}({\rm Edge}({\bm{f}}_{{\rm ref},i}^{*}))$
 11:     $\mathcal{F}_{{\rm ct},i}^{*}:={\tt ST\_SNAP\_CAND\_TOUCH}({\bm{f}}_{{\rm ref},i}^{*},\mathcal{F}_{{\rm cand},i},\mathcal{F}_{{\rm ct},i},\varepsilon_{S})$
@@
 13:Sort $\mathcal{F}^{*}$ in descending order of flow and length
-We outline the overall workflow ST_OVERLINE to compute a minimal flow map, by combining pre- and post-processing with the iteration of local alignment.
-For Algorithm 8, the inputs are the map matched routes $\mathcal{M}$, the split node type $S$, the blend tolerance $\varepsilon$, the simplify tolerance $\varepsilon_{D}$, the snap tolerance $\varepsilon_{S}$, the number of connected edges in reference linestrings ${\bm{k}}$, and the maximum number of iterations $j_{\max}$.
-In Step 1, we apply ST_OVERLINE_PLANR to the map matched routes to produce an initial flow map $\mathcal{F}$.
-In Step 2, we apply ST_SPLITNODE to this flow map, because applying ST_SPLITNODE on a flow map is more robust than on the map matched routes.
+In Figure 5(a) are the flow segments without any line blending, where, for example, the 3 segments with flow 2, 152 and 163, should be blended into a single segment.
+In Figure 5(b) are the flow segments after line blending.
+These 3 separate segments are blended into a segment with flow 317 (pink), which is the sum of their individual flows.
+In contrast, the segments with flows 8, 12 and 16, forming part of the T-junction, are blended into a segment with a weighted mean of 15.
+Due to the action of ST_OVERLINE_LINEBLEND, the line blended flow segments in Figure 5(b) are an improvement over Figure 5(a).
+Refer to caption Refer to caption
+Figure 5: (a) Flow segments without line blending. (b) Flow segments with line blending. Label is traffic flow.
+To compute a locally aligned flow map, we combine Algorithms 17 into an initial and an iterative phase.
+The initial phase ST_OVERLINE_INITIAL is described in Algorithm 8.
+Its inputs are the map matched routes $\mathcal{R}$, the split node type $S$, and the simplify tolerance $\varepsilon_{D}$.
+In Step 1, we apply ST_OVERLINE_PLANR to the map matched routes to produce a flow map $\mathcal{F}$.
+In Step 2, we apply ST_SPLITNODE to this flow map.
 In Step 3, we employ the standard ST_SIMPLIFY with simplify tolerance $\varepsilon_{D}$.
@@
 These simplified linestrings usually lead to more overlapping segments, which assist the flow aggregation in Step 4.
-Steps 5–10 is the iteration of ST_OVERLINE_LINEBLEND for the $k$-edge reference linestrings.
-For each $k$ in ${\bm{k}}$, we iterate until the maximum number of iterations $j_{\max}$ is reached or two consecutive flow maps are identical.
-Within each iteration, we search for the $k$-edge reference linestrings, blend the candidate linestrings, and snap the candidate-touching linestrings, and compute the flow aggregation.
-In Steps 11–14 is some housekeeping in ST_OVERLINE_PRUNE, where we remove pseudo nodes, and replace the incident edges with a concatenated edge with the weighted mean flow, as well as some loops and leaf edges with low flow values.
-Algorithm 8 ST_OVERLINE – Compute locally aligned flow map from map matched routes
-1:Input: $\mathcal{M}$ map matched routes, $S$ split node, $\varepsilon_{D}$ simplify tolerance, ${\bm{k}}$edges, $j_{\max}$ max #iterations, $\varepsilon$ blend tolerance, $\varepsilon_{S}$ snap tolerance
-2:Output: $\mathcal{F}$ aligned road segment flows
-3:$\mathcal{F}:={\tt ST\_OVERLINE\_PLANR}(\mathcal{M})$;
+Algorithm 8 ST_OVERLINE_INITIAL – Compute initial locally aligned flow map
+1:Input: $\mathcal{R}$ map matched routes, $S$ split node, $\varepsilon_{D}$ simplify tolerance
+2:Output: $\mathcal{F}$ initial aligned flow map
+3:$\mathcal{F}:={\tt ST\_OVERLINE\_PLANR}(\mathcal{R})$
 4:$\mathcal{F}:={\tt ST\_SPLITNODE}(\mathcal{F},S)$
@@
 6:$\mathcal{F}:={\tt ST\_OVERLINE\_PLANR}(\mathcal{F})$
-7:for $k$ in ${\bm{k}}$ do
+With an initial flow map on hand, we proceed iteratively towards a locally aligned flow map in Algorithm 9.
+The inputs of ST_OVERLINE are the initial flow map $\mathcal{F}$, the split node type $S$, the blend tolerance $\varepsilon$, the snap tolerance $\varepsilon_{S}$, the number of connected edges ${\bm{k}}$, and the maximum number of iterations $j_{\max}$.
+Steps 1–6 is the iteration of ST_OVERLINE_LINEBLEND for the $k$-edge reference linestrings.
+For each $k$ in ${\bm{k}}$, we iterate until $j_{\max}$ is reached or two consecutive flow maps are identical.
+Within each iteration, we search for the $k$-edge reference linestrings, blend the candidate linestrings, and snap the candidate-touching linestrings, and compute the flow aggregation.
+In Steps 7–10 is some housekeeping in ST_OVERLINE_PRUNE, where we remove pseudo nodes, remove leaf edges with low flow values, and replace thin loops with non-looped linestrings.
+Algorithm 9 ST_OVERLINE – Compute iterated locally aligned flow map
+1:Input: $\mathcal{F}$ initial flow map, $S$ split node, ${\bm{k}}$edges, $j_{\max}$ max #iterations, $\varepsilon$ blend tolerance, $\varepsilon_{S}$ snap tolerance
+2:for $k$ in ${\bm{k}}$ do
+3:     $\mathcal{F}_{\rm prev}:=\{\}$; $j:=0$
+4:     while $j<j_{\max}$ and $\mathcal{F}_{\rm prev}\neq\mathcal{F}$ do
+5:         $\mathcal{F}_{\rm prev}:=\mathcal{F}$; $j:=j+1$
+6:         $\mathcal{F}:={\tt ST\_OVERLINE\_LINEBLEND}(\mathcal{F},k,\varepsilon,\varepsilon_{S})$
+7:         $\mathcal{F}:={\tt ST\_OVERLINE\_PLANR}(\mathcal{F})$
 8:$\mathcal{F}_{\rm prev}:=\{\}$; $j:=0$
@@
 10:     $\mathcal{F}_{\rm prev}:=\mathcal{F}$; $j:=j+1$
-11:         $\mathcal{F}:={\tt ST\_OVERLINE\_LINEBLEND}(\mathcal{F},k,\varepsilon,\varepsilon_{S})$
-12:         $\mathcal{F}:={\tt ST\_OVERLINE\_PLANR}(\mathcal{F})$
-13:$\mathcal{F}_{\rm prev}:=\{\}$; $j:=0$
-14:while $j<j_{\max}$ and $\mathcal{F}_{\rm prev}\neq\mathcal{F}$ do
-15:     $\mathcal{F}_{\rm prev}:=\mathcal{F}$; $j:=j+1$
-16:     $\mathcal{F}:={\tt ST\_OVERLINE\_PRUNE}(\mathcal{F})$
-4 Results
-In this section we compute a minimal flow map for the Hannover trajectories.
-From the complete set of 1183 trajectories, we keep 1177 trajectories with length greater than $100$ m.
-We input the 1177 trajectories into ST_ROUTE with $M$ as the map matching and $R$ the route finding APIs from the Valhalla routing engine, and ${\bm{n}}_{W}=3,13,23,33,43,63,83$ waypoints.
-We employ the dockerised version 3.4.0 of the Valhalla routing engine (GIS OPS, 2023).
-Of these input trajectories, 1147 yield a sufficiently high quality match, where the Hausdorff distance $d_{\operatorname{Haus}}(M^{*}(G),G)<100$ m, and the ratio $\operatorname{len}(M^{*}(G))/\operatorname{len}(G)<1.1$.
-We continue the analysis with these 1147 map matched routes.
-For all iterations, we set $j_{\max}=20$.
-For the line blending, we begin with 1 iteration of ST_OVERLINE (Steps 1–4) with simplify tolerance $\varepsilon_{D}=1~{}\mathrm{m}$.
-We follow with ST_OVERLINE (Steps 5–14) with blend and snap tolerances $\varepsilon=\varepsilon_{S}=4~{}\mathrm{m}$, #edges ${\bm{k}}=1,2$, and node split type $S=$ ‘subdivision’.
-These simplify, blend and snap tolerances were chosen heuristically as a trade-off between being sufficiently large to account for the noisiness of the trajectories and the map matching/route finding APIs, whilst not being too large to obscure separate road segments within region with a dense road network.
-We set ${\bm{k}}=1,2$ since the search for connected $k$-edges with $k>2$ is too computationally intensive for our setup due to the large number of road segments (13 495).
-We set $S=$ ‘subdivision’ node splitting, as it provides more stable line blending priority at this early stage.
-We follow with 1 iteration of ST_OVERLINE (Steps 5–14), with the same tuning parameter choices, except with $S=$ ‘unary’.
-‘Unary’ node splitting is usually applied after an iteration of ‘subdivision’ node splitting since the former can now add any missing intersection nodes without adversely affecting the line blending priority.
-We end with 2 iterations of ST_OVERLINE (Steps 5–14), with $\varepsilon=\varepsilon_{S}=5~{}\mathrm{m},{\bm{k}}=1,2,3,4,S=$ ‘unary’, since the larger values for $\varepsilon,\varepsilon_{S}$ capture any line blending missed at 4 m, and the connected 3-, 4-edge searches become computationally feasible with the lower number of road segments.
-These 5 iterations of ST_OVERLINE converge to a minimal flow map with 1 413 road segments, i.e.
-a 89.5% reduction of the 13 495 segments in the initial flow map.
-Figure 9 illustrates the results of these iterations.
-In Figure 9(a) are the trajectories (green circles) and their map matched routes resulting from ST_ROUTE (blue lines).
-Whilst the map matched routes no longer completely obscure the road network, they remain misaligned to each other and to the road network.
-So it is not possible to accurately estimate the traffic flow map directly from them.
-In Figure 9(b) is a flow map resulting from ST_OVERLINE (Steps 1–4) with $S=$ ‘subdivision’.
-This is similar to the flow map that would be obtained by following Morgan and Lovelace (2021) without rasterisation.
-Since each linestring is labelled by its flow value, the crowding of the labels indicates that this is unlikely to be a minimal flow map.
-In Figure 9(c), we complete the line blending iteration ST_OVERLINE (Steps 5–14) with $S=$ ‘subdivision’.
-The crowding of the flow value labels is reduced so we are progressing to a minimal flow map.
-In Figure 9(d), we carry out a further 3 line blending iterations ST_OVERLINE (Steps 5–14) with $S=$ ‘unary’ to arrive at a minimal flow map.
-Refer to caption(a) Refer to caption(b)
-Refer to caption(c) Refer to caption(d)
-Figure 9: Iterations to compute locally aligned flow map.
-(a) Empirical trajectories (green circles), map matched routes from ST_ROUTE (blue lines).
-(b) Initial flow map from ST_OVERLINE (1–4), $S=$ ‘subdivision’.
-(c) Intermediate flow map from 1 iteration of ST_OVERLINE (5–14), $S=$ ‘subdivision’.
-(d) Minimal flow map from 3 further iterations of ST_OVERLINE (5–14), $S=$ ‘unary’.
+11:     $\mathcal{F}:={\tt ST\_OVERLINE\_PRUNE}(\mathcal{F})$
+2.5 Workflow
+Our proposed workflow from an input of trajectory data to a locally aligned flow map is illustrated in Figure 6.
+We begin with the trajectories $\mathcal{G}$.
+If the measurement errors in these trajectories are small enough to be input directly into ST_OVERLINE, then we set $\mathcal{R}=\mathcal{G}$.
+Otherwise to obtain $\mathcal{R}$, we carry out map matching using ST_ROUTE for each number of waypoints ${\bm{n}}_{W}=3,13,23,33,43,63,83$ since these ${\bm{n}}_{W}$ cover a range of possible values which likely includes an approximately optimal value.
+For the flow map aggregation, we begin with an iteration of ST_OVERLINE_INIITAL with simplify tolerance $\varepsilon_{D}=1$ m to create an initial flow map.
+In Figure 6, the dotted rectangle surrounding ST_LINEBLEND_PLANR and ST_SNAPNODE indicates that these algorithms are inputs to ST_OVERLINE_INIITAL.
+For the subsequent steps, we set the maximum number of iterations $j_{\max}=20$, and the snap tolerance $\varepsilon_{S}=\varepsilon$.
+We call ST_OVERLINE with #edges ${\bm{k}}=1,2$, and node split type $S=$ “subdivision”.
+At this early stage, the search for connected $k$-edges with $k>2$ can be computationally intensive, and “subdivision” node splitting provides more stable line blending priority.
+The dotted rectangle indicates the algorithms which are inputs to ST_OVERLINE.
+We follow with a call to ST_OVERLINE with “unary” node splitting.
+The “unary” node splitting is usually applied after “subdivision” node splitting since the former can now add any missing intersection nodes without adversely affecting the line blending priority.
+After these two steps, the flow map has begun to stabilise, and 3-, 4-edge searches become computationally feasible with the lower number of flow segments.
+We iterate ST_OVERLINE with ${\bm{k}}=1,2,3,4,S=$ “unary”, until the flow map $\mathcal{F}$ converges.
+In this workflow, the only unspecified tuning parameter is the blend tolerance $\varepsilon$, whose properties we investigate in the next section.
+Figure 6: Workflow for locally aligned flow map. Input are trajectories $\mathcal{G}$, output is locally aligned flow map $\mathcal{F}$. Blend tolerance $\varepsilon$ is tuning parameter.
+3 Results and discussion
+3.1 Synthetic trajectories
+Despite the ubiquity of GNSS-enabled devices, there is a lack of gold standard empirical trajectories and flow maps, due to privacy concerns and other reasons.
+Our solution is to rely on noise-free synthetic trajectories which are exactly aligned to a theoretical road network (Anastasiou et al., 2022a).
+These authors generate synthetic trajectories from an origin-destination matrix, and make available a set of 1.5 million synthetic trajectories for the metropolitan area of Los Angeles, USA (Anastasiou et al., 2022b).
+For illustrative purposes, we focus on a sample of $n=100$ trajectories, in EPSG 6339 (NAD83 (2011)/UTM zone 10N) coordinates, in Culver City (located about 16 km from downtown Los Angeles) which pass near the intersection of Main Street and Venice Boulevard, as indicated by the solid yellow line in Figure 7.
+We add a random uniform noise of ±$\pm 5$ m to the noise-free trajectories to obtain the noisy trajectories, displayed as the green lines in Figures 7(a–b).
+This noise mimics the typical measurement error of 5 m observed in empirical trajectories from GNSS-enabled smartphones (Van Diggelen and Enge, 2015).
+The reference noise-free flow map is in Figure 7(c), and the rasterised flow map in Figure 7(d).
+Morgan and Lovelace (2021) do not recommend a value for tuning the rasterisation grid resolution, so we select $\varepsilon_{R}=2$ m subjectively so that the raster grid cells do not exceed the road widths.
+Refer to caption Refer to caption
+Refer to caption Refer to caption
+Figure 7: Synthetic Culver City trajectories $(n=100)$.
+(a) Noisy trajectories (green lines).
+(b) Zoom of noisy trajectories near intersection of Main St and Venice Bd (solid yellow line).
+(c) Reference noise-free flow map.
+(d) Rasterised flow map with $\varepsilon_{R}=2$ m.
+Label in flow maps is traffic flow.
+For our locally aligned flow maps, the free tuning parameter is the blend tolerance parameter $\varepsilon$.
+So we investigate the effect of $\varepsilon$ on the accuracy of flow aggregation.
+For the blend tolerance $\varepsilon=1$ m, the flow map in Figure 8(a) is unable to consistently blend the map matched routes correctly since 1 m is typically smaller than the remaining discrepancies after map matching in ST_ROUTE.
+For the blend tolerance $\varepsilon=5$ m, the central road segments labelled with flow 51 and 45 in Figure 8(b), which are close to 49 and 45 in the reference flow map.
+For the blend tolerance $\varepsilon=10$ m, the central road segments are also labelled with flow 50 and 45 in Figure 8(c), similar to with blend tolerance $\varepsilon=5$ m in Figure 8(b).
+Near the intersection of Venice Bd and Main St (solid yellow line), the flow segment (with flow 17) is poorly aligned to the road network, unlike the more accurate alignment in Figure 8(b).
+Refer to caption Refer to caption
+Figure 8: Effect of blend tolerance $\varepsilon$ on synthetic Culver City flow maps.
+(a) Flow map with $\varepsilon=1$ m.
+(b) Flow map with $\varepsilon=5$ m.
+(c) Flow map with $\varepsilon=10$ m.
 Label is traffic flow.
-Colour (purple to orange) of road segments is proportional to traffic flow.
-4.1 Validation
-A visual inspection of the minimal flow map in Figure 9(d) reveals good alignment to the OSM road network in general, though there remain some data artefacts elsewhere.
-For example, in Figure 10(a), the green trajectories give no indication of the blue wonky route output by ST_ROUTE, and in Figure 10(c), the trajectories do not involve a loop, though ST_ROUTE contains a loop.
-Since our local alignment does not use an external road network to correct these wonky or extraneous loops in the map matched routes, they are propagated to the flow maps in Figure 10(a, d).
-Overall, these type of data artefacts are small and infrequent, and are associated with road segments with low flow.
-Refer to caption(a) Refer to caption(b)
-Refer to caption(a) Refer to caption(b)
-Figure 10: Data artefacts in locally aligned road segment flows.
-(a–b) Wonky road segments.
-(c–d) Extraneous road segments.
-(a, c) Trajectories (green circles), map matched routes (blue lines).
-(b, d) Colour (purple to orange) of road segments is proportional to traffic flow.
-For a more quantitative validation of the accuracy of our proposed flow map, we require a gold standard reference flow map.
-The experimental design of the Hannover trajectories is to serve as a reference data set for learning turning rules at road junctions, rather than as a reference flow map (Zourlidou et al., 2022).
-So we have to compute a proxy reference flow map.
-For this, we first compute line transects $\mathcal{T}_{i}=\{{\bm{t}}_{i,1},\dots,{\bm{t}}_{i,n_{\mathcal{T}_{i}}}\}={\tt ST\_TRANSECT}({\bm{f}}_{i},\varepsilon_{T},\delta_{T})$ for a road linestring ${\bm{f}}_{i}\in\mathcal{F}$, where a line transect is an orthogonal line segment (of length $2\varepsilon_{T}$ m) to ${\bm{f}}_{i}$.
-If $\operatorname{len}({\bm{f}}_{i})>\delta_{T}$ m, then these transects are placed at every $\delta_{T}$ m of ${\bm{f}}_{i}$, or if $\operatorname{len}({\bm{f}}_{i})\leq\delta_{T}$ m then the single transect is subtended from the centroid of ${\bm{f}}_{i}$.
-The empirical proxy flows are the number of intersection points between the map matched routes from $\mathcal{M}$ which intersect the line transects $f_{i,j}^{\mathrm{emp}}=\#\{{\tt ST\_INTERSECTION}(\mathcal{M},{\bm{t}}_{i,j}):{\bm{t}}_{i,j}\in\mathcal{T}_{i}\},j=1,\dots,n_{\mathcal{T}_{i}}$, and the mean discrepancy over all line transects of ${\bm{f}}_{i}$ is $\bar{f}_{i}^{\mathrm{emp}}=\mathrm{mean}\{f_{i,j}^{\mathrm{emp}},j=1,\dots,n_{\mathcal{T}_{i}}\}$.
-Our error measure of ${\bm{f}}_{i}$ is the absolute difference $\operatorname{Err}_{i,j}=|f_{i}-\bar{f}_{i}^{\mathrm{emp}}|$ for all line transects ${\bm{t}}_{j},j=1,\dots,n_{\mathcal{T}_{i}}$ which intersect ${\bm{f}}_{i}$.
-We also compute the relative error $\operatorname{RErr}_{i,j}=\operatorname{Err}_{i,j}/f_{i}$, which is well-defined since $f_{i}\geq 1$ for all ${\bm{f}}_{i}\in\mathcal{F}$.
-With the tuning parameters tolerance $\varepsilon_{T}=5$ m and $\delta_{T}=50$ m, then we compute 7 380 line transects.
-The colour (purple to orange) of the line transects is proportional to the absolute error.
-In Figure 11(a), the vast majority of the error values are low (purple), with only a few high error values (orange), in/near the solid black rectangle.
-The zoom of the black rectangle is given in Figure 11(b), along with the road segment flows in blue.
-The road segment ${\bm{f}}_{11}$ (thick blue segment) with the estimated flow is $f_{11}=238$, and the empirical proxy flow $\bar{f}_{11}^{\mathrm{emp}}=254$, has a high absolute error, i.e.
-$\operatorname{Err}_{11}=16,\operatorname{RErr}_{11}=0.07$.
-The road segment ${\bm{f}}_{331}$ with $f_{331}=25,\bar{f}_{331}^{\mathrm{emp}}=12$, has a higher relative error, i.e.
-$\operatorname{Err}_{331}=13,\operatorname{RErr}_{331}=0.52$.
-These errors are due to that these road segments, at an earlier iteration of the flow map, comprised shorter sub-segments with higher flows due to the presence of the more than $100$ trajectories passing near the intersection of ${\bm{f}}_{11}$ and ${\bm{f}}_{331}$.
-Since the nodes of these sub-segments are pseudo nodes, they were removed, and the concatenated sub-segments assigned the single weighted mean flow = 25.
-Refer to caption(a) Refer to caption(b)
-Figure 11: Validation of line transects of locally aligned flow map.
-(a) Neighbourhood level.
-(b) Zoom of black rectangle.
-Line transects computed with tolerance $\varepsilon_{T}=5$ m, at every $\delta_{T}=50$ m.
-Label is absolute error (Err).
-Colour (purple to orange) of line transects is proportional to absolute error.
-Width of road segments (blue) is proportional to traffic flow.
-To supplement the visual examination of these flow estimation errors, Figure 12 shows the bivariate (Err, RErr) histogram plot of the errors.
-The vast majority (6518 or 88.3%) in the orange hexagonal bin have zero absolute and relative error, and only 39 (0.53%) have Err $>4$, and 364 (4.9%) have RErr $>0.1$, and only 16 (0.21%) have Err $>4$ and RErr $>0.1$ which comprise most of the purple bins.
-Figure 12: Histogram plot of absolute and relative errors of line transects of locally aligned flow map.
-Label in hexagonal bin is number of line transects.
-Vertical dotted line is Err = 4, horizontal dotted line is RErr = 0.1.
-This demonstrates that our flow map is mostly accurate as it usually controls the estimation error.
-By accuracy, we mean by how close the estimated flows are to the empirical proxy flows.
-However accuracy is insufficient on its own: if we take the extreme situation of a flow map with a single road segment with zero error, then this has the highest accuracy, but we have no estimated flows outside of this single segment.
-So we require that the spatial coverage of the flow map be also high.
-We cannot resolve this question of spatial coverage unambiguously since we do not have a gold standard flow map, though we can verify that all $n=1147$ map matched routes intersect at least one line transect from the estimated flow map.
-Whilst this calculation does not exclude that there can be some regions of some map matched routes are without nearby line transects, we can further verify with a visual inspection that these regions tend to be small in area and/or comprise a low number of routes.
-So we claim that our minimal flow map has high levels of accuracy and spatial coverage.
-To conclude the validation of our flow maps, the flow map at the city level is illustrated in Figure 13(a).
-The orange segments with high flows are apparent, whereas these high flows were not apparent from the scatter plot of the trajectories in Figure 1(a).
-The desire line map (or spider diagram) is in Figure 13(b) and represent the traffic flows between the origin/destination hub nodes (black solid circles) These hub nodes are the centroids from a hierarchical clustering, similar to that in ST_SNAPNODE, of the boundary points of trajectories with the cutoff at 5 000 m.
-The straight lines indicate the trajectories whose origin/destination are associated with different hub nodes, and the circles indicate the trajectories whose origin/destination are associated with the same hub node.
-The desire line map is in effect a low resolution map of straight-line flows between the hub nodes only, whereas the high resolution map shows the traffic flows on all road segments.
-Refer to caption(a) Refer to caption(b)
-Figure 13: High and low resolution flow maps. (a) Locally aligned flow map. (b) Desire line map. Colour (purple to orange) and width of road segments is proportional to traffic flow.
-4.2 Software
-These analysis algorithms have been developed in R since the complexity of Algorithms 18 require a mix of advanced geospatial and statistical methods.
-As a compiled language, R can have slower execution times.
-There are two main computational bottlenecks.
-The first consists of the map matching/route finding in Algorithm 1.
-The Valhalla routing engine APIs are available as a web-based service (e.g.
-https://valhalla1.openstreetmap.de) or as a local dockerised image (e.g.
-https://github.com/gis-ops/docker-valhalla).
-We use the latter as it allows for faster computation since these local API requests are not sent to a remote web-based server, and can be parallelised on a stand-alone machine.
-We conduct a small study of the execution times based on 10 replicates of ST_ROUTE on 10 randomly selected trajectories on an Intel i5 Quad core 3.10 GHz machine running Ubuntu 22.04 and R 4.4.0.
-Executing ST_ROUTE with a local Valhalla API is around 7.1 times faster than the web-based API, and a parallelisation (with 3 workers) is around 1.8 times faster than a serial computation.
-This is less than 3 because the dockerised image is not optimised for simultaneous API calls.
-Combining these together, a parallelised local API achieves around a 12.6 fold speed improvement in comparison to a serial web-based API.
-The second bottleneck is the line blending in Algorithm 7.
-Based on the execution times based on 10 replicates of ST_OVERLINE on a subset of 433 trajectories, 3-worker parallelisation is around 1.8 times faster than a serial computation.
-This is less than 3 because only the repeated application of line blending (Algorithm 4) is parallelised, whilst the line blending priority (Algorithm 6) remains a serial computation.
-These speed factors are intended to be illustrative, since execution times, involving remote web servers APIs and parallelisation, are difficult to predict on different internet connections and machines.
-We tentatively claim that a local Valhalla API reduces the execution time by an order of magnitude, whereas parallelisation reduces it almost linearly by the number of workers.
-Our long term goal is to integrate fast implementations of these local alignment algorithms into an industry standard GIS, so GIS practitioners will be able to efficiently create flow maps from empirical trajectories in a familiar environment.
-Since this is under development, in the mean time, we provide a geopackage and QGIS project with the input trajectories, map matched routes, iterated flow maps and desire lines, as listed in Table 1.
-The interested reader is able to interactively explore in QGIS the added value of our proposed high resolution minimal flow map flowmap4, in comparison to the input trajectories traj, the map matched routes route, the desire line flow map flowmap_desire, and the flow map computed according to a leading alternative without rasterization Morgan and Lovelace (2021) which is similar to flowmap0.
-Layer Description $n$
-traj Empirical trajectories 1 147
-route Map matched routes ST_ROUTE 1 147
-flowmap0 Flow map ST_OVERLINE(1–4), $\varepsilon_{D}=1~{},S=$ ‘subdivision’ 13 495
-flowmap1 Flow map ST_OVERLINE(5–14), ${\bm{k}}=1,2,\varepsilon=\varepsilon_{S}=4,S=$ ‘subdivision’ 2 437
-flowmap2 Flow map ST_OVERLINE(5–14), ${\bm{k}}=1,2,\varepsilon=\varepsilon_{S}=4,S=$ ‘unary’ 1 953
-flowmap3 Flow map ST_OVERLINE(5–14), ${\bm{k}}=1,2,3,4,\varepsilon=\varepsilon_{S}=5,S=$ ‘unary’ 1 867
-flowmap4 Flow map ST_OVERLINE(5–14), ${\bm{k}}=1,2,3,4,\varepsilon=\varepsilon_{S}=5,S=$ ‘unary’ 1 413
-flowmap_desire Desire lines ST_DESIRELINE, $\varepsilon=5000$ 41
-Table 1: Geospatial layers in geopackage. The first column is layer name, the second is description, and the third is number of geospatial features $n$.
+To complement the above heuristic comparisons, we define a quantitative discrepancy between a flow map $\mathcal{F}$ and the reference flow map $\mathcal{F}_{\rm ref}$.
+First, we compute the transect points for each linestring ${\bm{f}}$ from $\mathcal{F}$: if $\operatorname{len}({\bm{f}})>\varepsilon_{T}$ m, then a point is placed at every $\varepsilon_{T}$ m of ${\bm{f}}$; or if $\operatorname{len}({\bm{f}})\leq\varepsilon_{T}$ m, then a single point is placed at the centroid of ${\bm{f}}$.
+So $\mathcal{T}=\{{\bm{t}}_{1},\dots,{\bm{t}}_{n_{\mathcal{T}}}\}$ is a set of $n_{\mathcal{T}}$ transect points.
+For each ${\bm{t}}_{i},i=1,\dots,n_{\mathcal{T}}$, we find the closest linestring in $\mathcal{F}_{\rm ref}$, that is, ${\bm{f}}_{{\rm ref},i}^{*}=\operatorname{argmin}_{{\bm{f}}_{\rm ref}\in\mathcal{F}_{\rm ref}}\mathrm{dist}({\bm{t}}_{i},{\bm{f}}_{\rm ref})$.
+The flow errors are $\operatorname{Err}_{i}=|f_{i}-f_{{\rm ref},i}^{*}|$ for the transect points on ${\bm{t}}_{i}\in\mathcal{T}\subseteq\mathcal{F}$.
+We reverse the order of the comparison to obtain the flow errors $\operatorname{Err}_{j}=|f_{{\rm ref},j}-f_{j}^{*}|$ for the transect points on ${\bm{t}}_{j}\in\mathcal{T}_{\rm ref}\subseteq\mathcal{F}_{\rm ref}$, and concatenate these two sets of flow errors.
+Refer to caption Refer to caption
+Refer to caption Refer to caption
+Figure 9: Flow errors for synthetic Culver City flow maps.
+(a) Flow map with $\varepsilon=1$ m.
+(b) Flow map with $\varepsilon=5$ m.
+(c) Flow map with $\varepsilon=10$ m.
+(d) Rasterised flow map with $\varepsilon_{R}=2$ m.
+Flow segments are coloured lines, transect points are grey circles, flow errors are labels.
+In Figure 9(a), we display the transect points with $\varepsilon_{T}=50$ m, as the grey circles, superposed over the flow map $\mathcal{F}$.
+Figures 9(a–c) are the flow errors for the locally aligned flow maps with blend tolerance $\varepsilon=1,5,10$ m, and Figure 9(d) is for the rasterised flow map.
+The $\varepsilon=1$ m flow map and the rasterised flow map lead to many high error values.
+Whereas for the other two locally aligned flow maps, the flow errors are similarly low.
+We repeat the flow map error calculations for $N=10$ re-samples of $n=100$ Culver City trajectories.
+The mean and SD of the flow errors for the locally aligned flow maps are $3.061\pm 0.565$ ($\varepsilon=1$ m), $0.446\pm 0.047$ ($\varepsilon=5$ m), $0.477\pm 0.048$ ($\varepsilon=10$ m), and for the rasterised flow map $1.514\pm 0.201$ ($\varepsilon_{R}=2$ m).
+Setting $\varepsilon=1$ m leads to under-blending since $\varepsilon$ is smaller than the misalignments in the map matched routes, and setting $\varepsilon=10$ m leads to over-blending since $\varepsilon$ is larger than the distance between separate road segments.
+The $\varepsilon_{R}=2$ m rasterised flow map is similarly under-blended like the $\varepsilon=1$ m flow map.
+So we tentatively recommend a value of about $\varepsilon=5$ m for the blend tolerance parameter since it leads to optimal line blending.
+3.2 Empirical trajectories
+We return to the empirical Hannover trajectories.
+From the 1183 trajectories, we keep 1177 trajectories with length greater than $100$ m, and follow the workflow in Figure 6.
+We map match these 1177 trajectories with ST_ROUTE, and we continue with 1152 of them which have a sufficiently high quality match, where the Hausdorff distance $\operatorname{dist}_{\operatorname{Haus}}(R(G),G)<100$ m, and the ratio $\operatorname{len}(R(G))/\operatorname{len}(G)<1.1$, for a trajectory $G$ and a map matched route $R(G)$.
+That is, we discard 25 trajectories or a 2.02% failure rate for the Valhalla map matching, which is consistent with previous studies (Saki and Hagen, 2022).
+After some trial and error, we set the blend tolerance of $\varepsilon=4$ m, which is 1 m less than the 5 m tentatively recommended from the synthetic trajectories computations above.
+The workflow leads to a locally aligned flow map displayed in Figure 10(a).
+The corresponding rasterised flow map is in Figure 10(b).
+At the city level, both of these flow maps visualise well the traffic flows.
+Though if we zoom into Klusreide (solid yellow circle), then we observe that in Figure 10(c) the locally aligned map displays the aggregated flows on linear road segments in comparison to the pixelated road segments of the rasterised flow map in Figure 10(d).
+Whilst we do not have a reference flow map for these empirical trajectories, like we had for the synthetic Los Angeles trajectories, we are still able to compute a proxy flows based on empirical counts of the map matched routes with orthogonal line transects from the flow map road segments.
+Along Klusriede, the locally aligned flow map is composed to two road segments (in pink) with 0 and 2 flow errors) corresponding to the proxy flows 312 and 315.
+For the rasterised flow map, some grid cells in the pink region have flow error 2.
+Though rasterisation some times leads to double counting and thus to much larger flow errors, e.g.
+in the orange region near the T-junction, the estimated flow is 643 whereas the proxy flow is 315, so the flow error is 328.
+The locally aligned map does not suffer from these rasterisation effects, so its flow aggregation is more accurate.
+Refer to caption Refer to caption
+Refer to caption Refer to caption
+Figure 10: Hannover flow maps.
+(a) Locally aligned flow map at city level.
+(b) Rasterised flow map at city level.
+Solid yellow circle is Klusriede.
+(c) Locally aligned flow map at road segment level.
+(d) Rasterised flow map at road segment level.
+Colour scale is traffic flow, label is traffic flow error.
+Our future goal is the integration of the local alignment algorithms into an industry standard web GIS.
+In the mean time, we make available interactive web maps at https://mvstat.net/traj which allow the user to freely explore our proposed multi-scale traffic flow maps from the individual road to the city-wide level.
+3.3 Computation
+We report the execution times for the flow map computations for the synthetic and empirical trajectories.
+Since it is difficult to reliably reproduce execution times on different systems, we provide our execution times as a guide only.
+These are obtained on a stand-alone set-up running Ubuntu 24.04 with 16Gb RAM and 4 × Intel i5 CPU @ 3.10 GHz.
+We deploy a local dockerised Valhalla routing engine API (https://github.com/gis-ops/docker-valhalla) (GIS OPS, 2023) rather than relying on a web-based service.
+Since the local map matching API requests can be efficiently parallelised on a stand-alone machine, we are confident that ST_ROUTE is suitably optimised.
+For $n=100$ synthetic Culver City trajectories, ST_ROUTE executes in 188 s.
+For $n=1000$ synthetic trajectories, it executes in 2233 s, which is 10 times longer for a 10-fold increase in $n$.
+So it is straightforward to infer a precise execution time of ST_ROUTE for large sample sizes from a small sample size.
+The execution time for ST_OVERLINE increases as $n$ increases, but not in the same uniform manner as ST_ROUTE as the former also depends on the value of $\varepsilon$.
+For $n=100$, the execution times for ST_OVERLINE are 1777 s ($\varepsilon=1$ m), 1049 s ($\varepsilon=5$ m), 880 s ($\varepsilon=10$ m); and for $n=1000$, they are 15 457 s ($\varepsilon=1$ m), 6002 s ($\varepsilon=5$ m), 3897 s ($\varepsilon=10$ m).
+So it is difficult to infer precisely the execution times for ST_OVERLINE for larger sample sizes and smaller blend tolerances from smaller $n$ and larger $\varepsilon$.
+This is because the line blending (ST_LINEBLEND) is parallelised, but the line blending priority (ST_LINEBLEND_PRIORITY) remains a serial computation.
+With an optimal value of the blend tolerance $\varepsilon=5$ m, the total execution time of the workflow for the $n=100$ Culver City trajectories is 1238 s or $\sim$21 mins, and for $n=1000$ is 8236 s or $\sim$137 mins.
+The total execution time for the $n=1183$ Hannover trajectories with $\varepsilon=4$ m is 5545 s or $\sim$92 mins.
+Thus locally aligned flow maps are currently suitable for offline analyses of small and medium sized data sets.
+Due to time and computing constraints, (a) data sets with a larger number of trajectories (like the full set of 1.5 million synthetic Los Angeles trajectories) and (b) online computations were out of scope.
+For these to be computationally feasible, we require additional implementation improvements in the local alignment algorithms, which we leave for future development.
 4 Conclusion
-We have introduced novel analysis algorithms to compute a flow map from empirical trajectories.
-Our starting point is to focus on aligning segments of the map matched routes rather than the complete routes.
-We define a spatial relation to set up local reference road segments, which allows us to align other nearby road segments to this local reference segment.
-This local alignment is the key innovation to computing a minimal flow map that is aligned to the underlying road network.
-We presented solid evidence for the high level of spatial resolution, accuracy and coverage for our proposed minimal flow map.
-Since it accurately shows the traffic flow on all road segments at all scales, it provides increased added value in comparison to the empirical trajectories, to the low resolution desire lines map, and to existing high resolution flow map methodologies.
+We have introduced an iterative workflow which takes noisy GNSS vehicle trajectories as input and gives an accurate traffic flow map as output.
+Local alignment of flow segments is the key innovation.
+To carry out this local alignment, we defined a spatial relation to set up local reference flow segments, which then allowed us to align other nearby flow segments to this local reference segment.
+We presented solid evidence for the high level of spatial resolution and accuracy for our proposed locally aligned flow map for both synthetic and empirical trajectories.
+We demonstrated that they effectively reduce the noise of the input trajectories, and accurately aggregate the traffic flows on all road segments, and so they are a valuable addition to the toolkit for multi-scale traffic flow analysis.

### v6

@@
 Figure 1: Trajectory data in Hannover, Germany.
-(a) Empirical trajectories (green circles).
-Klusreide road segment (solid yellow circle).
-(b) Zoom of trajectories and map matched routes (blue lines), with road network (solid black lines), near Klusreide.
-(d) Zoom of 2 m $\times$ 2 m rasterised flow aggregation.
-Colour scale is traffic flow in grid cells.
-If we zoom in on these trajectory points around Klusriede (solid yellow circle), a residential street near the A2 autobahn, in Figure 1(b), then we observe that the points deviate from the road network (solid black lines).
+(a) Empirical trajectories (green circles), Loebensteinstraße (solid cyan circle).
+(b) Zoom of trajectories and map matched routes (blue lines).
+(c) Zoom of ideal flow aggregation.
+(d) Zoom of existing flow aggregation.
+(e) Zoom of 2 m $\times$ 2 m rasterised flow aggregation.
+Colour scale/label is traffic flow in road segments/grid cells.
+If we zoom in on these trajectory points around Loebensteinstraße (solid cyan circle), formerly known as Hindenburgstraße, is a residential street near the A2 autobahn, then we observe that in Figure 1(b) the points deviate from the road network.
 These noisy trajectories are unsuitable for direct aggregation into a flow map, so the usual first step is to apply a map matching procedure (Quddus et al., 2007, Chao et al., 2020).
@@
 Nonetheless the misalignment is reduced to less than 5 m, which is, for many purposes, including most notably navigation, indeed a suitable result.
-For these inexactly aligned trajectories, applying the state-of-the-art traffic flow aggregation API (Morgan and Lovelace, 2021) leads to an inaccurate flow map.
-The authors resolve this by rasterising the flow map to a raster grid: and within each raster grid cell, the flow is the sum of the flows of all trajectories which intersect the grid cell.
-This rasterised flow map is displayed in Figure 1(c).
-The colour scale for the flow ranges from 100 (violet) to 300 (pink) to 600 (orange).
+The ideal flow aggregation from the empirical trajectories is in Figure 1(c).
+The colour scale for the flow ranges from 0 (violet) to 150 (pink) to 300 (yellow).
+Applying the state-of-the-art traffic flow aggregation API (Morgan and Lovelace, 2021) to the inexactly aligned trajectories leads to an inaccurate flow map in Figure 1(d).
+The authors resolve this by rasterising the flow map to a raster grid: within each raster grid cell, the flow is the sum of the flows of all trajectories which intersect the grid cell.
+This rasterised flow map is displayed in Figure 1(e).
 The 2 m $\times$ 2 m raster grid obscures the road network at this scale.
@@
 If we require a higher resolution, then we must recalculate the flow map on a raster grid with a higher resolution.
-Our goal is to produce a traffic flow map from map matched routes, which can be utilised at any scale without recalculation, from the city level to the individual road level.
+Our goal is to produce a traffic flow map from empirical trajectories which is aligned to the road network and which can be utilised at any scale without recalculation, from the city level to the individual road level.
 In our proposed local alignment approach, we infer which road segments should serve as a local reference, and then proceed to align other nearby road segments to it.
-To accomplish this, we introduce several novel algorithms which employ a mix of advanced statistical and geospatial methods.
-These include “node snapping” where the nearby boundary points of road segments are combined via statistical clustering, and “line blending” where road segments, which are near each other but do not share overlapping sub-segments, are aligned to maximise their overlapping sub-segments.
+To accomplish this, we introduce a novel algorithm “line blending” where road segments, which are near each other but do not share overlapping sub-segments, are aligned to maximise their overlapping sub-segments.
+Line blending is a “fuzzy” aggregation method which calculates accurate flow aggregation of noisy trajectories.
+This is in contrast to the current state-of-the-art flow aggregation methods which are accurate only for noise-free trajectories.
 We combine these novel algorithms into an iterative workflow which outputs a locally aligned flow map.
@@
 2 Local alignment of route segments for flow aggregation
-We represent the road network by a graph $\mathcal{N}=\mathcal{N}(\mathcal{E},\mathcal{V})$, where the $n_{\mathcal{E}}$ edges =$\mathcal{E}=\{{\bm{e}}_{1},\dots,{\bm{e}}_{n_{\mathcal{E}}}\}$ are the road segments and the $n_{\mathcal{V}}$ nodes/vertices $\mathcal{V}=\{{\bm{v}}_{1},\dots,{\bm{v}}_{n_{\mathcal{V}}}\}$ indicate that different road segments are accessible to/from each other at this node point.
+We represent the road network by a graph $\mathcal{N}=\mathcal{N}(\mathcal{E},\mathcal{V})$, where the $n_{\mathcal{E}}$ edges $\mathcal{E}=\{{\bm{e}}_{1},\dots,{\bm{e}}_{n_{\mathcal{E}}}\}$ are the road segments and the $n_{\mathcal{V}}$ nodes/vertices $\mathcal{V}=\{{\bm{v}}_{1},\dots,{\bm{v}}_{n_{\mathcal{V}}}\}$ indicate that different road segments are accessible to/from each other at this node point.
 We set $\mathcal{N}$ to be the OpenStreetMap (OSM) road network.
-Each path in the network is composed of a connected sequence of edges, and we refer to this as a “linestring” geometry, following the Open Geospatial Consortium (OGC) terminology (OGC, 2010).
+Each path in the network is composed of a connected sequence of edges, and we refer to this as a “linestring” geometry.
 We denote a single trajectory $G=\{{\bm{g}}_{1},\dots,{\bm{g}}_{n_{G}}\}$ as a temporally ordered sequence of $n_{G}$ points ${\bm{g}}_{i},i=1,\dots,n_{G}$.
-This is a “multipoint” geometry (OGC, 2010).
+This is a “multipoint” geometry.
 The points of a trajectory $G$ are not necessarily coincident with the road network $\mathcal{N}$.
@@
 We focus on the Valhalla routing engine (https://valhalla.github.io/valhalla), which includes its highly recommended map matching API (Saki and Hagen, 2022).
-To this off-the-shelf Valhalla map matching, we apply a post hoc modification to improve its suitability as input for flow map aggregation, as outlined in Algorithm 1.
-The inputs of ST_ROUTE are the trajectory $G$, the map matching API $M$, the route finding API $R$, and the number of waypoints ${\bm{n}}_{W}$ for the route finding.
-These waypoints guide the route finding API to pass near them.
-Since we do not have an a priori single optimal value for the number of waypoints, we consider a range of $w$ values ${\bm{n}}_{W}=(n_{W,1},\dots,n_{W,w})$.
-In Step 1, we compute the initial map matched route $M(G)$ from the trajectory $G$ by calling the map matching API $M$.
-This initial map matched route $M(G)=\{{\bm{m}}_{1},\dots,{\bm{m}}_{n_{M}}\}$ is a linestring with $n_{M}$ edges, with $n_{M}+1$ points.
-In Steps 2–5, we loop over the number of waypoints in ${\bm{n}}_{W}$.
-In Steps 3–4, we take a sample of $n_{W,i}$ waypoints, where the first waypoint ${\bm{w}}_{1}$ is the start point of ${\bm{m}}_{1}$ and the $n_{W,i}$th waypoint is the end point of ${\bm{m}}_{n_{M}}$, and the intermediate waypoints ${\bm{w}}_{2},\dots,{\bm{w}}_{n_{W,i}-1}$ are sampled from the start points of the edges ${\bm{m}}_{2},\dots,{\bm{m}}_{n_{M}}$.
-In Step 5, we call the route finding API $R$ with the waypoints ${\bm{w}}_{1},\dots,{\bm{w}}_{n_{W,i}}$.
-The result is the map matched route $R_{i}(G)=R({\bm{w}}_{1},\dots,{\bm{w}}_{n_{W,i}})$.
-In Step 6, we select the route with the smallest dynamic time warping (DTW) normalised distance between the routes $R_{i}(G)$ and the trajectory $G$, for $i=1,\dots w$.
-The DTW distance is based on the lengths of all the distortions of $R_{i}(G)$ to achieve maximal alignment between $R_{i}(G)$ and $G$ (Sakoe and Chiba, 1978, Giorgino, 2009).
-Algorithm 1 ST_ROUTE – Map matched route
-1:Input: $G$ trajectory, $M$ map matching API, $R$ route finding API, ${\bm{n}}_{W}$waypoints
-2:Output: $R(G)$ map matched route
-3:Compute initial map matched route $M(G)$ from empirical trajectory $G$
-4:for $i:=1$ to $w$ do
-5:     Set $n_{W}:=n_{W,i}$
-6:     Sample waypoints ${\bm{w}}_{1},\dots,{\bm{w}}_{n_{W}}$ from initial map matched $M(G)$
-7:     Compute map matched routes $R_{i}(G):=R({\bm{w}}_{1},\dots,{\bm{w}}_{n_{W}})$
-8:Select minimal route $R(G):=\operatorname{argmin}_{i\in\{1,\dots,w\}}\,\operatorname{DTW}(R_{i}(G),G)$
-We set $R$ to be the Valhalla Odin turn-by-turn route finding API to be consistent with our choice of $M$ as the Valhalla Meili map matching API.
-We set the number of waypoints as ${\bm{n}}_{W}=3,13,23,33,43,63,83$ since these ${\bm{n}}_{W}$ cover a range of possible values which likely includes an approximately optimal value.
-The map matched Hannover trajectories are displayed as the blue lines in Figure 1.
+The inputs to ST_ROUTE are the trajectory $G$ and the map matching API $M$.
+For readability, we defer the detailed description of the algorithm (and all subsequent algorithms) to the Appendix.
 2.2 Flow aggregation
-To the road network graph $\mathcal{N}=\mathcal{N}(\mathcal{E},\mathcal{V})$, we add the traffic flows on each of the network edges to form a flow map =$\mathcal{F}=\{(f,{\bm{f}}):{\bm{f}}\in\mathcal{N},f>0\}=\{(f_{1},{\bm{f}}_{1}),\dots,(f_{n_{\mathcal{F}}},{\bm{f}}_{n_{\mathcal{F}}})\}$ composed of $n_{\mathcal{F}}$ flows, where each ${\bm{f}}_{i}$ is composed of edges in $\mathcal{E}$, with traffic flow $f_{i}$, $i=1,\dots,n_{\mathcal{F}}$.
+To the road network graph $\mathcal{N}=\mathcal{N}(\mathcal{E},\mathcal{V})$, we add the traffic flows on each of the network edges to form a flow map $\mathcal{F}=\{(f,{\bm{f}}):{\bm{f}}\in\mathcal{N},f>0\}=\{(f_{1},{\bm{f}}_{1}),\dots,(f_{n_{\mathcal{F}}},{\bm{f}}_{n_{\mathcal{F}}})\}$ composed of $n_{\mathcal{F}}$ flows, where each ${\bm{f}}_{i}$ is composed of edges in $\mathcal{E}$, with traffic flow $f_{i}$, $i=1,\dots,n_{\mathcal{F}}$.
 Our goal is to estimate the traffic flows $f_{1},\dots,f_{n_{\mathcal{F}}}$ and the flow network graph $\mathcal{F}$.
 For exactly aligned trajectories, to compute the flow aggregation, we deploy the overline function in the R package stplanr, which we refer to as ST_OVERLINE_PLANR (Lovelace and Ellison, 2018).
-That is, starting with the map matched routes =$\mathcal{R}=\{R(G_{1}),\dots,R(G_{n})\}$, the flow map is $\mathcal{F}={\tt ST\_OVERLINE\_PLANR}(\mathcal{R})$.
-To compute an accurate flow map from noisy, empirical trajectories relies on resolving the crucial problem of how to aggregate similar, but not exactly overlapping, road segments.
-Many solutions have been offered, such as edge bundling (Zhou et al., 2013) and rasterisation (Wood et al., 2010, Morgan and Lovelace, 2021).
+That is, starting with the map matched routes $\mathcal{M}=\{M(G_{1}),\dots,M(G_{n})\}$, the flow map is $\mathcal{F}={\tt ST\_OVERLINE\_PLANR}(\mathcal{M})$.
+To compute an accurate flow map from noisy trajectories relies on resolving the crucial problem of how to aggregate similar, but not exactly overlapping, road segments.
+Historically most solutions have focused on eliminating the discrepancies in the map matching so that flow aggregation then becomes a straightforward procedure, drawing on the many map matching algorithms available.
+Despite this research effort, map matching algorithms which yield error-free output are limited to special cases, and none have the same availability and geographical coverage as the open source, global Valhalla map matching.
+Following the recommended best practice, we carry out some preprocessing to simplify the map matched routes to improve the flow aggregation.
+We highlight two existing methods, which we call node snapping and node splitting.
+Node snapping is where the boundary points of the traffic flow linestrings are snapped to each other.
+Since the former are also nodes of the flow map, this gives the name to the algorithm.
+We focus on snapping these nodes since the linestring misalignments are in part caused by the existence of nodes which are close to each other but not exactly equal.
+Since we are searching for points which are close to together, then this is well-suited to statistical clustering.
+There are many statistical clustering algorithms available, and we focus on hierarchical clustering.
+To resolve the computational bottleneck of hierarchical clustering, we approximate the 1-pass complete linkage clustering by a nested 2-pass clustering.
+We begin with an efficient single linkage clustering in the R package fastcluster (Müllner, 2013).
+Since single linkage can result in chain-like clusters, we compute a subsequent complete linkage clustering to break these potential chains.
+In this nested approach, the complete linkage distance matrix is calculated only within each single linkage cluster, and so it is less likely to reach computational limits.
+This approximation is required since fastcluster does not implement an efficient complete linkage clustering.
+The inputs to Algorithm 2 (ST_SNAPNODE) are the flow linestrings $\mathcal{F}$ and the snap tolerance $\varepsilon_{S}$.
+The output are the node snapped linestrings $\mathcal{F}^{*}$.
+As its name suggests, all node points of $\mathcal{F}$ within $\varepsilon_{S}$ distance of each other are snapped together.
+In comparison to the unsnapped linestrings $\mathcal{F}$, there are fewer snapped linestrings $\mathcal{F}^{*}$ which have higher flow values, are more interconnected, and share more exactly overlapping sub-segments.
+Node snapping is focused near the boundary points of the flow linestrings, so it does not address misalignments far from the boundary points.
+Moreover there remain some linestrings whose intersection is not correctly computed by ST_OVERLINE_PLANR.
+The solution to both of these problems is the explicit addition of the missing intersection nodes to these linestring, and then split these linestrings into simple linestring segments at these added nodes.
+This is known as node splitting.
+The inputs to Algorithm 3 (ST_SPLITNODE) are the flow linestrings $\mathcal{F}$ and the node splitting type $S$.
+Our method deploys for $S=$ “subdivision” to_spatial_subdivision in the sfnetworks package (van der Meer et al., 2021), and for $S=$ “unary” geos_unary_union in the geos package (Dunnington and Pebesma, 2020).
+The first option tends to find fewer intersection nodes and is similar to the intersection computation performed by ST_OVERLINE_PLANR.
+The second option finds more missing intersection nodes, and leads to more comprehensive flow aggregation.
+In Figure 2(a) are the map matched routes without any node splitting or node snapping.
+There are missing intersection nodes, and there are intersection nodes (solid black circles) which are close to each other.
+In Figure 2(b) is the flow aggregation after node splitting ($S=$ “unary”) and then node snapping (with snap tolerance $\varepsilon_{S}=4$ m).
+The result is that there are fewer flow segments with higher flow values.
+Due to the combined action of ST_SPLITNODE and ST_SNAPNODE, Figure 2(b) is closer to the ideal flow map topology than Figure 2(a).
+Figure 2: Unsplit/unsnapped and node split/node snapped routes.
+(a) Unsplit and unsnapped routes.
+(b) Node split and node snapped, with snap tolerance $\varepsilon_{S}=4$ m and with flow aggregation.
+Colour scale/label is traffic flow in road segments.
+Whilst the preprocessing does improve the flow aggregation, there remain some discrepancies, since Figure 2(b) is not quite the ideal flow map.
+To tackle directly to the problem of eliminating the discrepancies during flow aggregation, heuristic solutions have most likely been investigated, though few have been published so there is scarce literature on this problem.
+Among the few published solutions include edge bundling (Zhou et al., 2013) and rasterisation (Wood et al., 2010, Morgan and Lovelace, 2021).
 Edge bundling consists of clustering trajectory linestrings and replacing all cluster members with a single representative linestring.
@@
 By local alignment, we mean that we align segments of the routes, rather than the complete routes.
-Our proposal is divided into several algorithms, which we develop within the R statistical analysis environment due to its integrated access to advanced statistical and geospatial analysis methods.
-2.3 Node snapping and node splitting
-The proposed algorithm is ST_SNAPNODE, where the boundary points of the traffic flow linestrings are snapped to each other.
-Since the former are also nodes of the flow map, this gives the name to the algorithm.
-We focus on snapping these nodes since the linestring misalignments are in part caused by the existence of nodes which are close to each other but not exactly equal.
-Since we are searching for points which are close to together, then this is well-suited to statistical clustering.
-There are many statistical clustering algorithms available, and we focus on hierarchical clustering.
-To resolve the computational bottleneck of hierarchical clustering, we approximate the 1-pass complete linkage clustering by a nested 2-pass clustering.
-We begin with an efficient single linkage clustering in the R package fastcluster (Müllner, 2013).
-Since single linkage can result in chain-like clusters, we compute a subsequent complete linkage clustering to break these potential chains.
-In this nested approach, the complete linkage distance matrix is calculated only within each single linkage cluster, and so it is less likely to reach computational limits.
-Algorithm 2 is a description of ST_SNAPNODE.
-The inputs are the $n_{\mathcal{F}}$ flow linestrings =$\mathcal{F}=\{(f_{1},{\bm{f}}_{1}),\dots,(f_{n_{\mathcal{F}}},{\bm{f}}_{n_{\mathcal{F}}})\}$ and the snap tolerance $\varepsilon_{S}$.
-As its name suggests, all node points within $\varepsilon_{S}$ distance of each other are snapped together.
+Our proposal is developed within the R statistical analysis environment due to its integrated access to advanced statistical and geospatial analysis methods.
+2.3 Line blending to align similar linestrings to local reference linestring
+So far we have focused on improving the alignment of linestrings by resolving inconsistencies at their intersections.
+We now focus on aligning linestrings more generally.
+For this, we require a comparison relationship to establish an order of alignment of nearby linestrings.
+For two flow linestrings $(f_{1},{\bm{f}}_{1})$ and $(f_{2},{\bm{f}}_{2})$ from a flow map $\mathcal{F}$, we define that ${\bm{f}}_{2}$ is a candidate to be aligned to the reference ${\bm{f}}_{1}$ at threshold $\varepsilon\geq 0$ if
+${\bm{f}}_{2}\subseteq{\tt ST\_BUFFER}({\bm{f}}_{1},\varepsilon),\ f_{2}\leq f_{1}.$ (1)
+The condition on the flow values $f_{2}\leq f_{1}$ means that we place a higher priority on the reference linestring ${\bm{f}}_{1}$ since its flow $f_{1}$ is larger than the flow $f_{2}$ on the candidate linestring ${\bm{f}}_{2}$.
+The buffer zone has flat edges, e.g.
+ST_BUFFER(endCapStyle="FLAT") in the R package sf (Pebesma, 2018), so the buffer zone ends at the boundary points of ${\bm{f}}_{1}$.
+This avoids erroneously considering neighbouring segments of ${\bm{f}}_{1}$, which are connected to ${\bm{f}}_{1}$ as a part of a longer sequence, to be candidates to be aligned to ${\bm{f}}_{1}$.
+[…]
+The inputs in Algorithm 4 (ST_LINEBLEND) are the reference linestring ${\bm{f}}_{\rm ref}$, the candidate linestrings $\mathcal{F}_{\rm cand}$, and the blend tolerance $\varepsilon$.
+The output are the modified reference linestring ${\bm{f}}^{*}_{\rm ref}$ and projected candidate flow linestrings $\mathcal{F}^{*}_{\rm cand}$, all with added interior points for accurate calculation of exactly equal linestring segments.
+If there are many candidate linestrings, then this may lead to many sub-segments in the projected linestrings, each with their own flow values.
+So we assign the rounded value of the weighted mean flow, weighted by the sub-segment length, to all sub-segments.
+In Figure 3(a), we illustrate this line blending with the flow linestrings $(7,AB)$ in blue and $(2,AC)$ in orange.
+Since $AD\subset{\tt ST\_BUFFER}(AB,4\,\mathrm{m})$ (pale blue rectangle), and $2<7$, then $AC$ is the reference ${\bm{f}}_{\rm ref}$ and $AD$ is the candidate linestring ${\bm{f}}_{\rm cand}$, according to Equation (1).
+Line blending involves projecting $AD$ to $AC$, so $C$ is projected to $C^{\prime}$ (which lies exactly on the reference linestring).
+We also add $C^{\prime}$ to the reference linestring.
+The reference linestring becomes ${\bm{f}}^{*}_{\rm ref}=(7,AC^{\prime}B)$, and the candidate ${\bm{f}}^{*}_{\rm cand}=(2,AC^{\prime})$.
+Now ${\bm{f}}^{*}_{\rm ref},{\bm{f}}^{*}_{\rm cand}$ share the sub-segment $AC^{\prime}$ exactly, and ST_OVERLINE_PLANR gives the aggregated flows as $(9,AC^{\prime})$ and $(7,C^{\prime}B)$.
+We take the rounded value of the weighted mean of these two flow linestrings.
+The result is a single linestring $(9,AC^{\prime}B)$, as shown in Figure 3(b).
+Line blending is required to aggregate correctly the linestrings $(7,AC)$ and $(2,AD)$ since they do not share a common sub-segment.
+Figure 3: Line blending leads to correct flow aggregation.
+(a) Before line blending and flow aggregation.
+Reference linestring $(7,ABC)$ in blue, candidate linestring $(2,AD)$ in orange.
+(b) After blending candidate into reference linestring, with blend tolerance $\varepsilon=4$ m, and flow aggregation.
+Blended linestring is $(9,AC^{\prime}B)$.
+More complex situations arise when other linestrings touch the candidate linestring, but are not candidates themselves for blending.
+Since line blending projects the candidates to the reference linestring, then we have to also project these other linestrings to avoid leaving a gap in the updated flow map.
+The inputs to Algorithm 5 (ST_SNAP_CAND_TOUCH) are the reference linestring ${\bm{f}}_{\rm ref}$, the candidate-touching linestrings $\mathcal{F}_{\rm ct}$, and the snap tolerance $\varepsilon_{S}$.
+The output are the snapped candidate-touching flow linestrings $\mathcal{F}^{*}_{\rm ct}$, which maintain connectivity between $\mathcal{F}^{*}_{\rm ct}$ and ${\bm{f}}_{\rm ref}$.
+Figure 4(a) is an illustration of ST_SNAP_CAND_TOUCH with the reference $(7,AB)$ (blue), the candidate linestring $(2,AD)$ (orange), and the blending buffer zone with blend tolerance $\varepsilon=4$ m (pale blue rectangle).
+The candidate-touching linestring ${\bm{f}}_{\rm ct}=(5,CD)$ (purple) touches the candidate $AC$ at $C$.
+The orange line is entirely within the pale blue buffer zone, whereas the purple line extends outside of the buffer zone and so is not a candidate for blending to the reference linestring.
+If we apply ST_LINEBLEND to blend the candidate linestring, then $C$ is projected to $C^{\prime}$ to lie on the reference linestring, i.e., ${\bm{f}}_{\rm cand}^{*}=(2,AC^{\prime})$.
+If we apply ST_SNAP_CAND_TOUCH with snap tolerance $\varepsilon_{S}=4$ m to the candidate-touching linestring, then the intersection point $C$ is projected to the boundary point $B$, i.e., ${\bm{f}}^{*}_{\rm ct}=(5,BD)$.
+Thus line blending, candidate-touching snapping, and flow aggregation yields that ${\bm{f}}^{*}_{\rm ct}=(5,BD)$ remains connected to ${\bm{f}}_{\rm ref}^{*}=(9,AC^{\prime}B)$.
+Figure 4: Snap candidate-touching linestrings to reference linestring after line blending.
+(a) Before line blending and snapping.
+Reference linestring $(7,AC)$ in blue, candidate linestring $(2,AC)$ orange, and candidate touching linestring $(5,CD)$ purple.
+(b) After line blending (blend tolerance $\varepsilon=4$ m) and snapping (snap tolerance $\varepsilon_{S}=4$ m).
+Reference linestring becomes $(9,AC^{\prime}B)$, and candidate-touching linestring $(5,BD)$.
+To illustrate line blending for the Hannover map matched routes, Figure 5(a) is the flow segments without any line blending, and Figure 5(b) is the flow segments after line blending.
+Due to the action of ST_OVERLINE_LINEBLEND, each road segment in the line blended flow map Figure 5(b) corresponds unambiguously to a road segment in the OSM road network, and thus it approaches the ideal flow map.
+This is an improvement over the multiple road segments in the unblended flow map in Figure 5(a), and the pixelated road segments in the rasterised flow map in Figure 1(e).
+Figure 5: Hannover routes line blending. (a) Flow segments without line blending. (b) Flow segments with line blending. Colour scale/label is traffic flow.
+2.4 Workflow
+Line blending for accurate flow aggregation at a single intersection in a road network is demonstrated in Figure 5.
+To extend line blending to cover the entire road network, we require two further generalisations.
+First, we generalise the reference linestring to be a sequence of $k$ connected edges, each with their own flow.
+If we treat this edge sequence momentarily as a single linestring, with flow equal to the weighted mean flow, weighted by the length of the $k$ edges, then we can apply Equation (1) to search for potential candidate linestrings.
+Due to computational limitations, we retain that candidates be single edges with single flow values.
+The generalisation to $k$-edge reference linestrings allows us to blend candidate linestrings which exceed the buffer zones of 1-edge reference linestrings.
+Second, we determine the priority for line blending for a set of reference linestrings.
+We require that a reference linestring cannot be a candidate linestring to another reference linestring, and a candidate linestring is a candidate for one reference linestring only.
+These ensure that if two linestrings ${\bm{f}}_{2}\subseteq{\tt ST\_BUFFER}({\bm{f}}_{1},\varepsilon)$ and ${\bm{f}}_{1}\subseteq{\tt ST\_BUFFER}({\bm{f}}_{2},\varepsilon)$, then we select only one of them.
+The flow condition $f_{2}\leq f_{1}$ usually distinguishes the reference linestring, except when $f_{1}=f_{2}$.
+In this case, then we designate the longer linestring to be the reference.
+The inputs to Algorithm 6 (ST_LINEBLEND_PRIORITY) are the flow linestrings $\mathcal{F}$, the number of connected edges for the reference linestrings $k$, and the blend tolerance $\varepsilon$.
+The output is a set of non-overlapping reference linestrings $\mathcal{F}_{\rm ref}$, a set of unique candidate linestrings $\mathcal{F}_{\rm cand}$, and a set of (potentially repeated) candidate-touching $\mathcal{F}_{\rm ct}$ linestrings.
+By combining Algorithms 26, we obtain an algorithm for local alignment of flow segments, in Algorithm 7 (ST_OVERLINE_LINEBLEND).
+The inputs to ST_OVERLINE_LINEBLEND are the flow linestrings $\mathcal{F}$, the blend tolerance $\varepsilon$, and the snap tolerance $\varepsilon_{S}$.
+The output are the $\mathcal{F}^{*}$ aggregated aligned flow linestrings.
+To compute a locally aligned flow map, we combine Algorithms 17 into an initial and an iterative phase.
+The inputs to Algorithm 8 (ST_OVERLINE_INITIAL) are the map matched routes $\mathcal{M}$ and the split node type $S$.
+The output is an initial aligned flow map.
+With an initial flow map on hand, we proceed iteratively towards a locally aligned flow map.
+The inputs to this iterative step Algorithm 9 (ST_OVERLINE) are the initial flow map $\mathcal{F}$, the split node type $S$, the blend tolerance $\varepsilon$, the snap tolerance $\varepsilon_{S}$, the number of connected edges $\bm{k}$}, and the maximum number of iterations $j_{\max}$.
+After a suitable number of iterations, the output is a locally aligned flow map.
+Our proposed workflow from an input of trajectory data to a locally aligned flow map is illustrated in Figure 6.
+We begin with the trajectories $\mathcal{G}$.
+If the measurement errors in these trajectories are small enough to be input directly into ST_OVERLINE, then we set $\mathcal{M}=\mathcal{G}$.
+Otherwise we carry out map matching with ST_ROUTE.
+For the flow map aggregation, we begin with an iteration of ST_OVERLINE_INIITAL to create an initial flow map.
+In Figure 6, the dotted rectangle surrounding ST_LINEBLEND_PLANR and ST_SNAPNODE indicates that these algorithms are inputs to ST_OVERLINE_INIITAL.
+For the subsequent steps, we set the maximum number of iterations $j_{\max}=20$, and the snap tolerance $\varepsilon_{S}=\varepsilon$.
+We call ST_OVERLINE with #edges ${\bm{k}}=1,2$, and node split type $S=$ “subdivision”.
+At this early stage, the search for connected $k$-edges with $k>2$ can be computationally intensive, and “subdivision” node splitting provides more stable line blending priority.
+The dotted rectangle indicates the algorithms which are inputs to ST_OVERLINE.
+We follow with a call to ST_OVERLINE with “unary” node splitting.
+The “unary” node splitting is usually applied after “subdivision” node splitting since the former can now add any missing intersection nodes without adversely affecting the line blending priority.
+After these two steps, the flow map has begun to stabilise, and 3-, 4-edge searches become computationally feasible with the lower number of flow segments.
+We iterate ST_OVERLINE with ${\bm{k}}=1,2,3,4,S=$ “unary”, until the flow map $\mathcal{F}$ converges.
+In this workflow, the only unspecified tuning parameter is the blend tolerance $\varepsilon$, whose properties we investigate in the next section.
+Figure 6: Workflow for locally aligned flow map. Input are trajectories $\mathcal{G}$, output is locally aligned flow map $\mathcal{F}$. Blend tolerance $\varepsilon$ is tuning parameter.
+3 Results and discussion
+3.1 Synthetic trajectories
+Despite the ubiquity of GNSS-enabled devices, there is a lack of gold standard empirical trajectories and flow maps, due to privacy concerns and other reasons.
+Our solution is to rely on noise-free synthetic trajectories which are exactly aligned to a theoretical road network (Anastasiou et al., 2022a).
+These authors generate synthetic trajectories from an origin-destination matrix, and make available a set of 1.5 million synthetic trajectories for the metropolitan area of Los Angeles, USA (Anastasiou et al., 2022b).
+For illustrative purposes, we focus on a sample of $n=100$ trajectories, in EPSG 6339 (NAD83 (2011)/UTM zone 10N) coordinates, in Culver City (located about 16 km from downtown Los Angeles) which pass near the intersection of Main Street and Venice Boulevard, as indicated by the solid cyan line in Figure 7.
+We add a random uniform noise of $\pm 5$ m to the noise-free trajectories to obtain the noisy trajectories, displayed as the green lines in Figures 7(a–b).
+This noise mimics the typical measurement error of 5 m observed in empirical trajectories from GNSS-enabled smartphones (Van Diggelen and Enge, 2015).
+The reference noise-free flow map is in Figure 7(c), and the rasterised flow map in Figure 7(d).
+Morgan and Lovelace (2021) do not recommend a value for tuning the rasterisation grid resolution, so we select $\varepsilon_{R}=2$ m subjectively so that the raster grid cells do not exceed the road widths.
+Figure 7: Synthetic Culver City trajectories $(n=100)$.
+(a) Noisy trajectories (green lines).
+(b) Zoom of noisy trajectories near intersection of Main St and Venice Bd (solid cyan line).
+(c) Reference noise-free flow map.
+(d) Rasterised flow map with $\varepsilon_{R}=2$ m.
+Colour scale/label is traffic flow.
+For our locally aligned flow maps, the free tuning parameter is the blend tolerance parameter $\varepsilon$.
+So we investigate the effect of $\varepsilon$ on the accuracy of flow aggregation.
+For the blend tolerance $\varepsilon=0$ m, this means that no line blending is performed, so we compute an unblended flow map in Figure 8(a) which comprises many small road segments with low flow values.
+For the blend tolerance $\varepsilon=1,5$ m, the flow map in Figure 8(b–c), the flow maps are similar.
+Though we observe that $\varepsilon=1$ m leads to an unblended road segment with flow 1 (in the upper north west corner)m, since 1 m is smaller than the remaining trajectory discrepancies.
+For the blend tolerance $\varepsilon=10$ m, near the intersection of Venice Bd and Main St (solid cyan line), the flow segment (with flow 16) is poorly aligned to the road network, unlike the more accurate alignment in Figures 8(b–c).
+Figure 8: Effect of blend tolerance $\varepsilon$ on synthetic Culver City flow maps.
+(a) Unblended flow map with $\varepsilon=0$ m.
+(b) Flow map with $\varepsilon=1$ m.
+(c) Flow map with $\varepsilon=5$ m.
+(d) Flow map with $\varepsilon=10$ m.
+Colour scale/label is traffic flow.
+We observe the usual sensitivity trade-off in the evolution of the values of the blend tolerance parameter.
+If this value is small (e.g.
+0 or 1 m in Figures 8(a–b)), then the action of line blending is limited to a small radius.
+The input trajectories are insufficiently processed, so the output flow map contains too many road segments.
+Whilst each discrepancy to the true road network is small (low bias), whenever the input trajectories are perturbed slightly, this would lead to a different flow map (high variability).
+If the blend tolerance is large (e.g.
+10 m in Figure 8(d)), the action of line blending is extended to a large radius.
+When the blend tolerance is larger than the distance between two distinct neighbouring road segments, then input trajectories which belong to different segments on the true road network can be erroneously blended together, and the output flow map contains an estimated road segment with a large discrepancy from any true road segment.
+Whilst the discrepancy is large (high bias), whenever the input trajectories are perturbed slightly, this would lead to a similar flow map (low variability).
+If the blend tolerance is between these two extremes (e.g.
+5 m in Figure 8(c)), then we achieve a trade-off between bias and variability.
+Therefore we set the blend tolerance to be sufficiently large so that it simplifies the flow map topology which approaches the true road network, and also not be too large so that it oversimplifies the flow map topology which deviates from the true road network.
+To complement the above heuristic comparisons, we define a quantitative discrepancy between a flow map $\mathcal{F}$ and the reference flow map $\mathcal{F}_{\rm ref}$.
+First, we compute the transect points for each linestring $\bm{f}$} from $\mathcal{F}$: if $\operatorname{len}({\bm{f}})>\varepsilon_{T}$ m, then a point is placed at every $\varepsilon_{T}$ m of $\bm{f}$}; or if $\operatorname{len}({\bm{f}})\leq\varepsilon_{T}$ m, then a single point is placed at the centroid of $\bm{f}$}.
+So $\mathcal{T}=\{{\bm{t}}_{1},\dots,{\bm{t}}_{n_{\mathcal{T}}}\}$ is a set of $n_{\mathcal{T}}$ transect points.
+For each ${\bm{t}}_{i},i=1,\dots,n_{\mathcal{T}}$, we find the closest linestring in $\mathcal{F}_{\rm ref}$, that is, ${\bm{f}}_{{\rm ref},i}^{*}=\operatorname{argmin}_{{\bm{f}}_{\rm ref}\in\mathcal{F}_{\rm ref}}\mathrm{dist}({\bm{t}}_{i},{\bm{f}}_{\rm ref})$.
+The flow errors are $\operatorname{Err}_{i}=|f_{i}-f_{{\rm ref},i}^{*}|$ for the transect points on ${\bm{t}}_{i}\in\mathcal{T}\subseteq\mathcal{F}$.
+We reverse the order of the comparison to obtain the flow errors $\operatorname{Err}_{j}=|f_{{\rm ref},j}-f_{j}^{*}|$ for the transect points on ${\bm{t}}_{j}\in\mathcal{T}_{\rm ref}\subseteq\mathcal{F}_{\rm ref}$, and concatenate these two sets of flow errors.
+Figure 9: Flow errors for synthetic Culver City flow maps.
+(a) Unblended flow map with $\varepsilon=0$ m.
+(b) Flow map with $\varepsilon=1$ m.
+(c) Flow map with $\varepsilon=5$ m.
+(d) Flow map with $\varepsilon=10$ m.
+(e) Rasterised flow map with $\varepsilon_{R}=2$ m.
+Colour scale is traffic flow, label is traffic flow error.
+In Figure 9, we display the transect points with $\varepsilon_{T}=50$ m, as the grey circles, superposed over the flow map $\mathcal{F}$.
+Figure 9(a) is the unblended flow map, Figures 9(b–d) are the flow errors for the locally aligned flow maps with blend tolerance $\varepsilon=1,5,10$ m, and Figure 9(e) is the rasterised flow map.
+The rasterised flow map lead to many high error values.
+Whereas for the locally aligned flow maps, the flow errors are low, especially for $\varepsilon=1,5$ m flow maps.
+We repeat the flow map error calculations for $N=10$ re-samples of $n=100$ Culver City trajectories.
+The mean and SD of the flow errors for the unblended flow map is $0.386\pm 0.048$ ($\varepsilon=0$ m), for the locally aligned flow maps $0.351\pm 0.055$ ($\varepsilon=1$ m), $0.350\pm 0.056$ ($\varepsilon=5$ m), $0.374\pm 0.061$ ($\varepsilon=10$ m), and for the rasterised flow map $0.438\pm 0.035$ ($\varepsilon_{R}=2$ m).
+Setting $\varepsilon=0$ m leads to under-blending since $\varepsilon$ is smaller than the misalignments in the map matched routes, and setting $\varepsilon=10$ m leads to over-blending since $\varepsilon$ is larger than the distance between separate road segments.
+The $\varepsilon_{R}=2$ m rasterised flow map is similarly under-blended like the $\varepsilon=0$ m flow map.
+So we tentatively recommend a value between 1 and 5 m for the blend tolerance parameter $\varepsilon$ since it leads to optimal line blending.
+This optimal value is a trade-off between the magnitude of the misalignments in the empirical trajectories and the road density of the underlying road network.
+3.2 Empirical trajectories
+We return to the empirical Hannover trajectories.
+From the 1183 trajectories, we keep 1181 trajectories with length greater than 0 m, and follow the workflow in Figure 6.
+We map match these 1181 trajectories with ST_ROUTE, and we continue with 1176 of them which have a sufficiently high quality match, where the Hausdorff distance $\operatorname{dist}_{\operatorname{Haus}}(R(G),G)<100$ m, and the ratio $\operatorname{len}(R(G))/\operatorname{len}(G)<1.1$, for a trajectory $G$ and a map matched route $R(G)$.
+That is, we discard 5 trajectories or a 0.42% failure rate for the Valhalla map matching, which is consistent with previous studies (Saki and Hagen, 2022).
+After some trial and error, we set the blend tolerance of $\varepsilon=4$ m, which is consistent with the tentative recommendation of $\varepsilon$ between 1 and 5 m from the synthetic trajectories computations above.
+The workflow leads to a locally aligned flow map displayed in Figure 10(a).
+The corresponding unblended flow map is Figure 10(b) and the rasterised flow map is Figure 10(c).
+At the city level, all these flow maps visualise well the traffic flows.
+Figure 10: Hannover flow maps at city level. (a) Locally aligned flow map. (b) Unblended flow map. (c) Rasterised flow map. Solid cyan circle is Loebensteinstraße. Colour scale is traffic flow.
+If we zoom into Loebensteinstraße (solid cyan circle), then we observe that in Figure 11(a) the locally aligned map displays the aggregated flows on single linear road segments in comparison to the multiple road segments of the unblended flow map in Figure 11(b), and the pixelated road segments of the rasterised flow map in Figure 11(c).
+Whilst we do not have a reference flow map for these empirical trajectories, like we had for the synthetic Los Angeles trajectories, we are still able to compute a proxy flows based on empirical counts of the map matched routes with orthogonal line transects from the flow map road segments.
+Along Loebensteinstraße, the locally aligned flow map is composed to two road segments (in pink) with 0 and 2 flow errors corresponding to the proxy flows 70 and 58.
+For the unblended and rasterised maps, the flow errors vary from 0 to over 100.
+Rasterisation some times leads to double counting and thus to larger flow errors, e.g.
+in the orange region near the intersection, the estimated flow is 365 whereas the proxy flow is 224, so the flow error is 141.
+The locally aligned map does not suffer from these rasterisation effects or from an insufficiently simplified topology, so its flow aggregation is more accurate.
+Figure 11: Hannover flow maps at road segment level.
+(a) Locally aligned flow map.
+(b) Unblended flow map.
+(c) Rasterised flow map.
+Solid cyan circle is Loebensteinstraße.
+Colour scale is traffic flow, label is traffic flow error.
+For our purposes, the synthetic trajectories in Los Angeles, and the empirical trajectories collected from a single driver in Hannover, can be considered to represent the complete driver population.
+So the resulting flow maps describe the complete traffic flows.
+This is not the case for the usual crowd-sourced trajectory data which are drawn from a highly biased sample of the complete driver population, and so the resulting flow map is a highly biased estimate of the complete traffic flow.
+We leave it to future work to apply the necessary adjustments to compute an unbiased estimate of the complete traffic flow map (Iqbal et al., 2014, Miller et al., 2020).
+3.3 Computation
+We report the execution times for the flow map computations for the synthetic and empirical trajectories.
+Since it is difficult to reliably reproduce execution times on different systems, we provide our execution times as a guide only.
+These are obtained on a stand-alone set-up running Ubuntu 24.04 with 16Gb RAM and 4 × Intel i5 CPU @ 3.10 GHz.
+We deploy a local dockerised Valhalla routing engine API (https://github.com/gis-ops/docker-valhalla) (GIS OPS, 2023) rather than relying on a web-based service.
+Since the local map matching API requests can be efficiently parallelised on a stand-alone machine, we are confident that ST_ROUTE is suitably optimised.
+For $n=100$ synthetic Culver City trajectories, ST_ROUTE executes in 18 s.
+For $n=1000$ synthetic trajectories, it executes in 158 s, which is about 10 times longer for a 10-fold increase in $n$.
+So it is straightforward to infer a precise execution time of ST_ROUTE for large sample sizes from a small sample size.
+The execution time for ST_OVERLINE increases as $n$ increases, but not in the same uniform manner as ST_ROUTE as the former also depends on the value of $\varepsilon$.
+For $n=100$, the execution times for ST_OVERLINE are 94 s ($\varepsilon=1$ m), 85 s ($\varepsilon=5$ m), 99 s ($\varepsilon=10$ m); and for $n=1000$, they are 530 s ($\varepsilon=1$ m), 358 s ($\varepsilon=5$ m), 409 s ($\varepsilon=10$ m).
+So it is difficult to infer precisely the execution times for ST_OVERLINE for larger sample sizes and smaller blend tolerances from smaller $n$ and larger $\varepsilon$.
+This is because the line blending (ST_LINEBLEND) is only partially parallelised.
+With a blend tolerance $\varepsilon=5$ m, the total execution time of the workflow for the $n=100$ Culver City trajectories is 103 s or $\sim$2 mins, and for $n=1000$ is 243 s or $\sim$4 mins.
+The total execution time for the $n=1183$ Hannover trajectories with $\varepsilon=4$ m is 1514 s or $\sim$25 mins, reflecting the more complex structure of empirical trajectories compared to synthetic trajectories.
+Thus locally aligned flow maps are currently suitable for offline analyses of small and medium sized data sets.
+Due to time and computing constraints, (a) data sets with a larger number of trajectories (like the full set of 1.5 million synthetic Los Angeles trajectories) and (b) online computations were out of scope.
+For these to be computationally feasible, we require additional implementation improvements in the local alignment algorithms, which we leave for future development.
+Our other future goal is the integration of the local alignment algorithms into an industry standard web GIS.
+In the mean time, we make available interactive web maps at https://mvstat.net/traj which allow the user to freely explore our proposed multi-scale traffic flow maps from the individual road to the city-wide level.
+4 Conclusion
+We have introduced an iterative workflow which takes noisy GNSS vehicle trajectories as input and gives an accurate traffic flow map as output.
+Local alignment of flow segments is the key innovation.
+To carry out this local alignment, we defined a spatial relation to set up local reference flow segments, which then allowed us to align other nearby flow segments to this local reference segment.
+We presented solid evidence for the high level of spatial resolution and accuracy for our proposed locally aligned flow map for both synthetic and empirical trajectories.
+We demonstrated that they effectively reduce the noise of the input trajectories, and accurately aggregate the traffic flows on all road segments, and so they are a valuable addition to the toolkit for multi-scale traffic flow analysis.
+Appendix A Details of algorithms
+A.1 Map matched route
+The inputs to ST_ROUTE are the trajectory $G$ and the map matching API $M$.
+The output is a map matched route $M(G)$.
+In Step 1, we compute the map matched route $M(G)$ from the trajectory $G$ by calling the map matching API $M$.
+This map matched route $M(G)=\{{\bm{m}}_{1},\dots,{\bm{m}}_{n_{M}}\}$ is a linestring with $n_{M}$ edges, with $n_{M}+1$ points.
+Algorithm 1 ST_ROUTE – Map matched route
+1:Input: $G$ trajectory, $M$ map matching API
+2:Output: $M(G)$ map matched route
+3:Compute map matched route $M(G)$ from empirical trajectory $G$
+A.2 Snap node clustering of linestrings
+The inputs to ST_SNAPNODE are the $n_{\mathcal{F}}$ flow linestrings $\mathcal{F}=\{(f_{1},{\bm{f}}_{1}),\dots,(f_{n_{\mathcal{F}}},{\bm{f}}_{n_{\mathcal{F}}})\}$ and the snap tolerance $\varepsilon_{S}$.
+The output are the snapped linestrings $\mathcal{F}^{*}$.
 In Step 1, we extract the boundary points of all flow linestrings.
@@
 In Steps 13–16, within each of these $C^{*}$ clusters, for all points of the corresponding flow linestrings which are closer than $\varepsilon_{S}$ distance to the centroid, we snap them to the centroid.
-In Steps 17–18, we collate and sort all the $n_{\mathcal{F}^{*}}$ snapped linestrings =$\mathcal{F}^{*}=\{(f_{1}^{*},{\bm{f}}_{1}^{*}),\dots,(f_{n_{\mathcal{F}^{*}}}^{*},{\bm{f}}_{n_{\mathcal{F}^{*}}}^{*})\}$.
-In comparison to the unsnapped linestrings $\mathcal{F}$, there are fewer snapped linestrings $\mathcal{F}^{*}$ which have higher flow values, are more interconnected, and share more exactly overlapping sub-segments.
+In Steps 17–18, we collate and sort all the $n_{\mathcal{F}^{*}}$ snapped linestrings $\mathcal{F}^{*}=\{(f_{1}^{*},{\bm{f}}_{1}^{*}),\dots,(f_{n_{\mathcal{F}^{*}}}^{*},{\bm{f}}_{n_{\mathcal{F}^{*}}}^{*})\}$.
 Algorithm 2 ST_SNAPNODE – Snap node clustering of linestrings
@@
 18:  $\mathcal{F}_{i}^{*}$ := rejoin snapped and unsnapped points into linestrings
-19:Collate snapped linestrings $\{\mathcal{F}_{1}^{*},\dots\mathcal{F}_{C^{\prime*}}^{*}\}$ into :=$\mathcal{F}^{*}:=\{(f_{1}^{*},{\bm{f}}_{1}^{*}),\dots,(f_{n_{\mathcal{F}^{*}}}^{*},{\bm{f}}_{n_{\mathcal{F}^{*}}}^{*})\}$
+19:Collate snapped linestrings $\{\mathcal{F}_{1}^{*},\dots\mathcal{F}_{C^{\prime*}}^{*}\}$ into $\mathcal{F}^{*}:=\{(f_{1}^{*},{\bm{f}}_{1}^{*}),\dots,(f_{n_{\mathcal{F}^{*}}}^{*},{\bm{f}}_{n_{\mathcal{F}^{*}}}^{*})\}$
 20:Sort $\mathcal{F}^{*}$ by descending flow and length
-Node snapping is focused near the boundary points of the flow linestrings, so it does not address misalignments far from the boundary points.
-Moreover there remain some linestrings whose intersection is not correctly computed by ST_OVERLINE_PLANR.
-The solution to both of these problems is the explicit addition of the missing intersection nodes to these linestrings.
-ST_SPLITNODE is the collation of a couple of algorithms which explicitly add the nodes at the intersections in the interior of linestrings, and splits these linestrings into simple linestring segments at these added nodes.
-This gives the name to this method.
-The inputs to Algorithm 3 are the flow linestrings $\mathcal{F}$ and the node splitting type $S$.
-Our method deploys for $S=$ “subdivision” in Steps 1–3, to_spatial_subdivision in the sfnetworks package (van der Meer et al., 2023), and for $S=$ “unary” in Steps 4–5, geos_unary_union in the geos package (Dunnington and Pebesma, 2023).
-The first option tends to find fewer intersection nodes and is similar to the intersection computation performed by ST_OVERLINE_PLANR.
-The second option finds more missing intersection nodes, and leads to more comprehensive flow aggregation.
-We require both types of node splitting to compute a locally aligned flow map.
+A.3 Split nodes at interior intersections of linestrings
+The inputs to ST_SPLITNODE are the flow linestrings $\mathcal{F}$ and the node splitting type $S$.
+The output are the split linestrings with added nodes $\mathcal{F}^{*}$.
+We deploy, for $S=$ “subdivision” in Steps 1–3, to_spatial_subdivision in the sfnetworks package (van der Meer et al., 2021), and for $S=$ “unary” in Steps 4–5, geos_unary_union in the geos package (Dunnington and Pebesma, 2020).
 Algorithm 3 ST_SPLITNODE – Split nodes at interior intersections of linestrings
@@
 4:if $S==$ “subdivision” then
-5:     $\mathcal{F}^{*}:=$ sfnetworks::to_spatial_subdivision$(\mathcal{N}^{*})$
+[…]
 6:else if $S==$ “unary” then
-7:     $\mathcal{F}^{*}:=$ geos::geos_unary_union$(\mathcal{N}^{*})$
-In Figure 2(a) are the flow segments without any node splitting or node snapping.
-There are missing intersection nodes, and there are intersection nodes which are close to each other.
-In Figure 2(b) are the flow segments after node splitting ($S=$ “unary”) and then node snapping (with snap tolerance $\varepsilon_{S}=4$ m).
-The result is that there are fewer flow segments with higher flow values.
-Due to the combined action of ST_SPLITNODE and ST_SNAPNODE, Figure 2(b) is an improvement over Figure 2(a).
-Refer to caption Refer to caption
-Figure 2: Unsplit/unsnapped and node split/node snapped flow segments, with snap tolerance $\varepsilon_{S}=4$ m. (a) Unsplit and unsnapped. (b) Node split and node snapped.
-2.4 Line blending to align similar linestrings with local reference
-So far we have focused on improving the alignment of linestrings by resolving inconsistencies at their intersections.
-We now focus on aligning linestrings more generally.
-For this, we require a comparison relationship to establish an order of alignment of nearby linestrings.
-For two flow linestrings $(f_{1},{\bm{f}}_{1})$ and $(f_{2},{\bm{f}}_{2})$ from a flow map $\mathcal{F}$, we define that ${\bm{f}}_{2}$ is a candidate to be aligned to the reference ${\bm{f}}_{1}$ at threshold $\varepsilon\geq 0$ if
-${\bm{f}}_{2}\subseteq{\tt ST\_BUFFER}({\bm{f}}_{1},\varepsilon),\ f_{2}\leq f_{1}.$ (1)
-The condition on the flow values $f_{2}\leq f_{1}$ means that we place a higher priority on the reference linestring ${\bm{f}}_{1}$ since its flow $f_{1}$ is larger than the flow $f_{2}$ on the candidate linestring ${\bm{f}}_{2}$.
-The buffer zone has flat edges, e.g.
-ST_BUFFER(endCapStyle="FLAT") in the R package sf (Pebesma, 2018), so the buffer zone ends at the boundary points of ${\bm{f}}_{1}$.
-This avoids erroneously considering neighbouring segments of ${\bm{f}}_{1}$, which are connected to ${\bm{f}}_{1}$ as a part of a longer sequence, to be candidates to be aligned to ${\bm{f}}_{1}$.
-Let ${\bm{f}}_{\rm ref}$ be a reference linestring from $\mathcal{F}$.
-The set of $n_{\mathcal{F}_{\rm cand}}$ linestrings from \$\mathcal{F}\backslash\{{\bm{f}}_{\rm ref}\}$ which satisfy Equation (1) are the candidate linestrings $\mathcal{F}_{\rm cand}=\{{\bm{f}}_{{\rm cand},1},\dots,{\bm{f}}_{{\rm cand},n_{\mathcal{F}_{\rm cand}}}\}$, with the convention that if $n_{\mathcal{F}_{\rm cand}}=0$ then $\mathcal{F}_{\rm cand}$ is the empty set.
-We call our approach “line blending”, since we blend the candidate linestrings onto the reference linestring.
-The inputs in Algorithm 4 are the reference linestring ${\bm{f}}_{\rm ref}$, the $n_{\mathcal{F}_{\rm cand}}$ candidate linestrings $\mathcal{F}_{\rm cand}$, and the blend tolerance $\varepsilon$.
+[…]
+A.4 Blend candidate linestrings onto reference linestring
+The inputs to ST_LINEBLEND are the reference linestring ${\bm{f}}_{\rm ref}$, the $n_{\mathcal{F}_{\rm cand}}$ candidate linestrings $\mathcal{F}_{\rm cand}$, and the blend tolerance $\varepsilon$.
 The output are the modified reference linestring and $n_{\mathcal{F}_{\rm cand}}$ modified candidate flow linestrings, all with added interior points for accurate calculation of exactly equal linestring segments.
 In Step 1, we initialize a local network graph $\mathcal{N}^{*}$ with the reference linestring.
-In Steps 2–3, we extract all points of the candidate linestrings, and use the network_blend function in the sfnetworks package (which we denote as ST_NETWORK_BLEND) to blend efficiently these points into the reference linestring (van der Meer et al., 2023).
+In Steps 2–3, we extract all points of the candidate linestrings, and use the network_blend function in the sfnetworks package (which we denote as ST_NETWORK_BLEND) to blend efficiently these points into the reference linestring (van der Meer et al., 2021).
 ST_NETWORK_BLEND projects the candidate linestrings onto the reference linestring, and explicitly adds them to the network, thereby creating new edges in $\mathcal{N}^{*}$.
@@
 8:Collate $\mathcal{F}^{*}_{\rm cand}:=\{{\bm{f}}_{{\rm cand},1}^{*},\dots,{\bm{f}}_{{\rm cand},n_{\mathcal{F}_{\rm cand}}}^{*}\}$
-9:${\bm{f}}^{*}_{{\rm ref}}$ := shortest path from $\operatorname{Start}({\bm{f}}_{{\rm ref}})$ to $\operatorname{End}({\bm{f}}_{{\rm ref}})$ along network $\mathcal{N}^{*}$
-The result from ${\tt ST\_LINEBLEND}$ is the modified reference linestring and the projected candidate linestrings with added interior points.
-If there are many candidate linestrings, then this may lead to many sub-segments in the projected linestrings, each with their own flow values.
-So we assign the rounded value of the weighted mean flow, weighted by the sub-segment length, to all sub-segments.
-In Figure 3(a), we illustrate this line blending with the flow linestrings $(7,ABC)$ in blue and $(2,AD)$ in orange.
-As these two linestrings do not share a common sub-segment, so flow aggregation does not modify them.
-Since $AD\subset{\tt ST\_BUFFER}(ABC,4\,\mathrm{m})$ (the pale blue rectangle), and $2<7$, then $ABC$ is the reference ${\bm{f}}_{\rm ref}$ and $AD$ is the candidate linestring ${\bm{f}}_{\rm cand}$, according to Equation (1).
-Line blending involves projecting $AD$ to $ABC$, so $D$ is projected to $D^{\prime}$ (which lies exactly on the reference linestring), and $B$ is added explicitly to this projected linestring.
-We also add $D^{\prime}$ to the reference linestring.
-The reference linestring becomes ${\bm{f}}^{*}_{\rm ref}=(7,ABD^{\prime}C)$, and the candidate ${\bm{f}}^{*}_{\rm cand}=(2,ABD^{\prime})$.
-Now ${\bm{f}}^{*}_{\rm ref},{\bm{f}}^{*}_{\rm cand}$ share the sub-segment $ABD^{\prime}$ exactly, and ST_OVERLINE_PLANR gives the aggregated flows as $(9,ABD^{\prime})$ and $(2,D^{\prime}C)$.
-We take the rounded value of the weighted mean of these two flow linestrings.
-The result is a single linestring $(9,ABD^{\prime}C)$, as shown in Figure 3(b).
-Refer to caption Refer to caption
-Figure 3: Line blending leads to correct flow aggregation.
-(a) Before line blending and flow aggregation.
-Reference linestring $(7,ABC)$ in blue, candidate linestring $(2,AD)$ in orange.
-(b) After blending candidate into reference linestring, with tolerance $\varepsilon=4$ m, and flow aggregation.
-Blended linestring is $(9,ABD^{\prime}C)$.
-After aggregation, the projected point $D^{\prime}$ in $ABD^{\prime}C$ appears to be redundant.
-Suppose that we have another linestring $(3,ABC)$ without $D^{\prime}$.
-Then ST_OVERLINE_PLANR does not aggregate $(9,ABD^{\prime}C)$ with $(3,ABC)$ even if $D^{\prime}$ is exactly coincident on the segment $BC$, since this algorithm does not explicitly add any missing intersection nodes prior to aggregation (Morgan and Lovelace, 2021).
-It outputs three linestrings $(12,AB),(9,BC),(3,BD^{\prime}C)$, rather than a single linestring $(12,ABD^{\prime}C)$.
-So we retain these intermediate points in the flow linestrings to assist in subsequent flow aggregation.
-More complex situations arise when other linestrings touch the candidate linestring, but are not candidates themselves for blending.
-Since line blending projects the candidates to the reference linestring, then we have to also project these other linestrings to avoid leaving a gap in the updated flow map.
-This procedure ST_SNAP_CAND_TOUCH is outlined in Algorithm 5.
-The inputs are the reference linestring ${\bm{f}}_{\rm ref}$, the $n_{\mathcal{F}_{\rm cand}}$ candidate linestrings $\mathcal{F}_{\rm cand}$, the $n_{\mathcal{F}_{\rm ct}}$ candidate-touching linestrings $\mathcal{F}_{\rm ct}$, and the snap tolerance $\varepsilon_{S}$.
+[…]
+A.5 Snap candidate-touching linestrings onto reference linestring
+The inputs to ST_SNAP_CAND_TOUCH are the reference linestring ${\bm{f}}_{\rm ref}$, the $n_{\mathcal{F}_{\rm cand}}$ candidate linestrings $\mathcal{F}_{\rm cand}$, the $n_{\mathcal{F}_{\rm ct}}$ candidate-touching linestrings $\mathcal{F}_{\rm ct}$, and the snap tolerance $\varepsilon_{S}$.
+The output are $\mathcal{F}^{*}_{\rm ct}$ the snapped candidate-touching flow linestrings.
 In Steps 1–8, we iterate over each candidate-touching linestring.
@@
 In Step 9, we collate these snapped linestrings.
-Algorithm 5 ST_SNAP_CAND_TOUCH – Snap candidate-touching linestrings onto reference
+Algorithm 5 ST_SNAP_CAND_TOUCH – Snap candidate-touching linestrings onto reference linestring
 1:Input: ${\bm{f}}_{\rm ref}$ reference, $\mathcal{F}_{\rm cand}$ candidate, $\mathcal{F}_{\rm ct}$ candidate-touching flow linestrings, $\varepsilon_{S}$ snap tolerance
@@
 11:Collate $\mathcal{F}^{*}_{\rm ct}:=\{{\bm{f}}^{*}_{{\rm ct},1},\dots,{\bm{f}}^{*}_{{\rm ct},n_{\mathcal{F}_{\rm ct}}}\}$
-Figure 4(a) is an illustration of ST_SNAP_CAND_TOUCH with the reference $(7,ABC)$ (blue), the candidate linestring $(2,AD)$ (orange), and the blending buffer zone with blend tolerance $\varepsilon=4$ m (light blue).
-The candidate-touching linestring ${\bm{f}}_{\rm ct}=(5,DE)$ (purple) touches the candidate $AD$ at $D$.
-The orange line is entirely within the pale blue buffer zone, whereas the purple line extends outside of the buffer zone and so is not a candidate for blending to the reference linestring.
-If we apply ST_LINEBLEND to blend the candidate linestring, then $D$ is projected to $D^{\prime}$ to lie on the reference linestring, i.e., ${\bm{f}}_{\rm cand}^{*}=(2,ABD^{\prime})$.
-If we apply ST_SNAP_CAND_TOUCH with snap tolerance $\varepsilon_{S}=4$ m to the candidate-touching linestring, then the intersection point $D$ is projected to the boundary point $C$, i.e., ${\bm{f}}^{*}_{\rm ct}=(5,CE)$.
-Thus line blending, candidate-touching snapping, and flow aggregation yields that ${\bm{f}}^{*}_{\rm ct}=(5,CE)$ remains connected to ${\bm{f}}_{\rm ref}^{*}=(9,ABD^{\prime}C)$.
-Refer to caption Refer to caption
-Figure 4: Snap candidate-touching linestrings to reference linestring after line blending, with blend tolerance $\varepsilon=4$ m, snap tolerance $\varepsilon_{S}=4$ m.
-(a) Before line blending and snapping.
-Reference linestring $(7,ABC)$ in blue, candidate linestring $(2,AD)$ orange, and candidate touching linestring $(5,DE)$ purple.
-(b) After line blending and snapping.
-Reference linestring becomes $(9,ABD^{\prime}C)$, and candidate-touching linestring $(5,CE)$.
-We generalise the reference linestring to be a sequence of $k$ connected edges, each with their own flow.
-If we treat this edge sequence momentarily as a single linestring, with flow equal to the weighted mean flow, weighted by the length of the $k$ edges, then we can apply Equation (1) to search for potential candidate linestrings.
-Due to computational limitations, we retain that candidates be single edges with single flow values.
-The generalisation to $k$-edge reference linestrings allows us to blend candidate linestrings which exceed the buffer zones of 1-edge reference linestrings.
-We have described blending candidate and candidate-touching linestrings to a single reference linestring (possibly composed of $k$ edges).
-The next step is to determine the priority for line blending for a set of reference linestrings.
-We require that a reference linestring cannot be a candidate linestring to another reference linestring, and a candidate linestring is a candidate for one reference linestring only.
-These ensure that if two linestrings ${\bm{f}}_{2}\subseteq{\tt ST\_BUFFER}({\bm{f}}_{1},\varepsilon)$ and ${\bm{f}}_{1}\subseteq{\tt ST\_BUFFER}({\bm{f}}_{2},\varepsilon)$, then we select only one of them.
-The flow condition $f_{2}\leq f_{1}$ usually distinguishes the reference linestring, except when $f_{1}=f_{2}$.
-In this case, then we designate the longer linestring to be the reference.
-The inputs to ST_LINEBLEND_PRIORITY in Algorithm 6 are the flow linestrings $\mathcal{F}$, the number of connected edges for the reference linestrings $k$, and the blend tolerance $\varepsilon$.
+A.6 Compute line blending priority
+The inputs to ST_LINEBLEND_PRIORITY are the flow linestrings $\mathcal{F}$, the number of connected edges for the reference linestrings $k$, and the blend tolerance $\varepsilon$.
 The output is a set of non-overlapping reference linestrings $\mathcal{F}_{\rm ref}$, a set of unique candidate linestrings $\mathcal{F}_{\rm cand}$, and a set of (potentially repeated) candidate-touching $\mathcal{F}_{\rm ct}$ linestrings.
@@
 Algorithm 6 ST_LINEBLEND_PRIORITY – Compute line blending priority
-1:Input: $\mathcal{F}$ flow linestrings, $k$edges, $\varepsilon$ blend tolerance
+1:Input: $\mathcal{F}$ flow linestrings, $k$ #edges, $\varepsilon$ blend tolerance
 2:Output: $(\mathcal{F}_{\rm ref},\mathcal{F}_{\rm cand},\mathcal{F}_{\rm ct})$ reference, candidate, candidate-touching linestrings
@@
 20:Extract $\mathcal{F}_{\rm cand}:=\{\mathcal{F}^{*}_{{\rm cand},1},\dots,\mathcal{F}^{*}_{{\rm cand},n_{\mathcal{Q}}}\},\mathcal{F}_{\rm ct}:=\{\mathcal{F}^{*}_{{\rm ct},1},\dots,\mathcal{F}^{*}_{{\rm ct},n_{\mathcal{Q}}}\}$ from $\mathcal{Q}$
-The local alignment of flow segments, in Algorithm 7, is established by combining Algorithms 26.
+A.7 Locally align segment flows using line blending
 The inputs to ST_OVERLINE_LINEBLEND are the flow linestrings $\mathcal{F}$, the split node type $S$, the number of connected edges $k$, the blend tolerance $\varepsilon$, and the snap tolerance $\varepsilon_{S}$.
+The output are the modified reference linestring ${\bm{f}}^{*}_{\rm ref}$ and projected candidate flow linestrings $\mathcal{F}^{*}_{\rm cand}$, all with added interior points.
 In Step 1, we node split the flow linestrings $\mathcal{F}$.
@@
 Algorithm 7 ST_OVERLINE_LINEBLEND – Locally align segment flows using line blending
-1:Input: $\mathcal{F}$ flow linestrings, $S$ split node, $k$edges, $\varepsilon$ blend tolerance, $\varepsilon_{S}$ snap tolerance
+1:Input: $\mathcal{F}$ flow linestrings, $S$ split node, $k$ #edges, $\varepsilon$ blend tolerance, $\varepsilon_{S}$ snap tolerance
 2:Output: $\mathcal{F}^{*}$ aggregated aligned flow linestrings
-3:$\mathcal{F}:={\tt ST\_SPLITNODE}(\mathcal{F},S)$
-4:$\mathcal{F}:={\tt ST\_SNAPNODE}(\mathcal{F},\varepsilon_{S})$
-5:$\{\mathcal{F}_{\rm ref},\mathcal{F}_{\rm cand},\mathcal{F}_{\rm ct}\}:={\tt ST\_LINEBLEND\_PRIORITY}(\mathcal{F},k,\varepsilon)$
-6:$\Edgesubscriptrefsubscriptcandsubscriptct\mathcal{F}^{c}:=\mathcal{F}\backslash({\rm Edge}(\mathcal{F}_{\rm ref})\cup\mathcal{F}_{\rm cand}\cup\mathcal{F}_{\rm ct})$
+[…]
+[…]
+[…]
+[…]
 7:for $i:=1$ to $n_{\mathcal{F}_{\rm ref}}$ do
@@
 9:  ${\bm{f}}_{{\rm ref},i}^{*}:={\tt ST\_OVERLINE\_PLANR}({\bm{f}}_{{\rm ref},i}^{*}\cup\mathcal{F}_{{\rm cand},i}^{*})$
-10:     $f^{*}_{{\rm ref},i}$ := weighted mean of ${\bm{f}}^{*}_{{\rm ref},i}$, weight $:=\mathrm{len}({\rm Edge}({\bm{f}}_{{\rm ref},i}^{*}))$
+10:  $f^{*}_{{\rm ref},i}$ := weighted mean of ${\bm{f}}^{*}_{{\rm ref},i}$, weight $\mathrm{len}({\rm Edge}({\bm{f}}_{{\rm ref},i}^{*}))$
 11:  $\mathcal{F}_{{\rm ct},i}^{*}:={\tt ST\_SNAP\_CAND\_TOUCH}({\bm{f}}_{{\rm ref},i}^{*},\mathcal{F}_{{\rm cand},i},\mathcal{F}_{{\rm ct},i},\varepsilon_{S})$
-12:$\mathcal{F}^{*}:=\mathcal{F}_{\rm ref}^{*}\cup\mathcal{F}_{\rm ct}^{*}\cup\mathcal{F}^{c}$
+[…]
 13:Sort $\mathcal{F}^{*}$ in descending order of flow and length
-In Figure 5(a) are the flow segments without any line blending, where, for example, the 3 segments with flow 2, 152 and 163, should be blended into a single segment.
-In Figure 5(b) are the flow segments after line blending.
-These 3 separate segments are blended into a segment with flow 317 (pink), which is the sum of their individual flows.
-In contrast, the segments with flows 8, 12 and 16, forming part of the T-junction, are blended into a segment with a weighted mean of 15.
-Due to the action of ST_OVERLINE_LINEBLEND, the line blended flow segments in Figure 5(b) are an improvement over Figure 5(a).
-Refer to caption Refer to caption
-Figure 5: (a) Flow segments without line blending. (b) Flow segments with line blending. Label is traffic flow.
-To compute a locally aligned flow map, we combine Algorithms 17 into an initial and an iterative phase.
-The initial phase ST_OVERLINE_INITIAL is described in Algorithm 8.
-Its inputs are the map matched routes $\mathcal{R}$, the split node type $S$, and the simplify tolerance $\varepsilon_{D}$.
+A.8 Compute initial locally aligned flow map
+The inputs to ST_OVERLINE_INITIAL are the map matched routes $\mathcal{M}$, the split node type $S$.
+The output is an initial aligned flow map.
 In Step 1, we apply ST_OVERLINE_PLANR to the map matched routes to produce a flow map $\mathcal{F}$.
 In Step 2, we apply ST_SPLITNODE to this flow map.
-In Step 3, we employ the standard ST_SIMPLIFY with simplify tolerance $\varepsilon_{D}$.
-These simplified linestrings are modified so that all modified segments are at most $\varepsilon_{D}$ distance from the unmodified segments (Ramer, 1972, Douglas and Peucker, 1973).
-These simplified linestrings usually lead to more overlapping segments, which assist the flow aggregation in Step 4.
+In Step 3, we apply the flow aggregation.
 Algorithm 8 ST_OVERLINE_INITIAL – Compute initial locally aligned flow map
-1:Input: $\mathcal{R}$ map matched routes, $S$ split node, $\varepsilon_{D}$ simplify tolerance
+1:Input: $\mathcal{M}$ map matched routes, $S$ split node, $\varepsilon_{D}$ simplify tolerance
 2:Output: $\mathcal{F}$ initial aligned flow map
-3:$\mathcal{F}:={\tt ST\_OVERLINE\_PLANR}(\mathcal{R})$
-4:$\mathcal{F}:={\tt ST\_SPLITNODE}(\mathcal{F},S)$
-5:$\mathcal{F}:={\tt ST\_SIMPLIFY}(\mathcal{F},\varepsilon_{D})$
-6:$\mathcal{F}:={\tt ST\_OVERLINE\_PLANR}(\mathcal{F})$
-With an initial flow map on hand, we proceed iteratively towards a locally aligned flow map in Algorithm 9.
+[…]
+[…]
+[…]
+A.9 Compute iterated locally aligned flow map
 The inputs of ST_OVERLINE are the initial flow map $\mathcal{F}$, the split node type $S$, the blend tolerance $\varepsilon$, the snap tolerance $\varepsilon_{S}$, the number of connected edges $\bm{k}$}, and the maximum number of iterations $j_{\max}$.
+The output is an iteration of a locally aligned flow map.
 Steps 1–6 is the iteration of ST_OVERLINE_LINEBLEND for the $k$-edge reference linestrings.
@@
 Algorithm 9 ST_OVERLINE – Compute iterated locally aligned flow map
-1:Input: $\mathcal{F}$ initial flow map, $S$ split node, ${\bm{k}}$edges, $j_{\max}$ max #iterations, $\varepsilon$ blend tolerance, $\varepsilon_{S}$ snap tolerance
+1:Input: $\mathcal{F}$ initial flow map, $S$ split node, $\bm{k}$} #edges, $j_{\max}$ max #iterations, $\varepsilon$ blend tolerance, $\varepsilon_{S}$ snap tolerance
 2:for $k$ in $\bm{k}$} do
@@
 7:   $\mathcal{F}:={\tt ST\_OVERLINE\_PLANR}(\mathcal{F})$
-8:$\mathcal{F}_{\rm prev}:=\{\}$; $j:=0$
+[…]
 9:while $j<j_{\max}$ and $\mathcal{F}_{\rm prev}\neq\mathcal{F}$ do
@@
 11:  $\mathcal{F}:={\tt ST\_OVERLINE\_PRUNE}(\mathcal{F})$
-2.5 Workflow
-Our proposed workflow from an input of trajectory data to a locally aligned flow map is illustrated in Figure 6.
-We begin with the trajectories $\mathcal{G}$.
-If the measurement errors in these trajectories are small enough to be input directly into ST_OVERLINE, then we set $\mathcal{R}=\mathcal{G}$.
-Otherwise to obtain $\mathcal{R}$, we carry out map matching using ST_ROUTE for each number of waypoints ${\bm{n}}_{W}=3,13,23,33,43,63,83$ since these ${\bm{n}}_{W}$ cover a range of possible values which likely includes an approximately optimal value.
-For the flow map aggregation, we begin with an iteration of ST_OVERLINE_INIITAL with simplify tolerance $\varepsilon_{D}=1$ m to create an initial flow map.
-In Figure 6, the dotted rectangle surrounding ST_LINEBLEND_PLANR and ST_SNAPNODE indicates that these algorithms are inputs to ST_OVERLINE_INIITAL.
-For the subsequent steps, we set the maximum number of iterations $j_{\max}=20$, and the snap tolerance $\varepsilon_{S}=\varepsilon$.
-We call ST_OVERLINE with #edges ${\bm{k}}=1,2$, and node split type $S=$ “subdivision”.
-At this early stage, the search for connected $k$-edges with $k>2$ can be computationally intensive, and “subdivision” node splitting provides more stable line blending priority.
-The dotted rectangle indicates the algorithms which are inputs to ST_OVERLINE.
-We follow with a call to ST_OVERLINE with “unary” node splitting.
-The “unary” node splitting is usually applied after “subdivision” node splitting since the former can now add any missing intersection nodes without adversely affecting the line blending priority.
-After these two steps, the flow map has begun to stabilise, and 3-, 4-edge searches become computationally feasible with the lower number of flow segments.
-We iterate ST_OVERLINE with ${\bm{k}}=1,2,3,4,S=$ “unary”, until the flow map $\mathcal{F}$ converges.
-In this workflow, the only unspecified tuning parameter is the blend tolerance $\varepsilon$, whose properties we investigate in the next section.
-Figure 6: Workflow for locally aligned flow map. Input are trajectories $\mathcal{G}$, output is locally aligned flow map $\mathcal{F}$. Blend tolerance $\varepsilon$ is tuning parameter.
-3 Results and discussion
-3.1 Synthetic trajectories
-Despite the ubiquity of GNSS-enabled devices, there is a lack of gold standard empirical trajectories and flow maps, due to privacy concerns and other reasons.
-Our solution is to rely on noise-free synthetic trajectories which are exactly aligned to a theoretical road network (Anastasiou et al., 2022a).
-These authors generate synthetic trajectories from an origin-destination matrix, and make available a set of 1.5 million synthetic trajectories for the metropolitan area of Los Angeles, USA (Anastasiou et al., 2022b).
-For illustrative purposes, we focus on a sample of $n=100$ trajectories, in EPSG 6339 (NAD83 (2011)/UTM zone 10N) coordinates, in Culver City (located about 16 km from downtown Los Angeles) which pass near the intersection of Main Street and Venice Boulevard, as indicated by the solid yellow line in Figure 7.
-We add a random uniform noise of ±$\pm 5$ m to the noise-free trajectories to obtain the noisy trajectories, displayed as the green lines in Figures 7(a–b).
-This noise mimics the typical measurement error of 5 m observed in empirical trajectories from GNSS-enabled smartphones (Van Diggelen and Enge, 2015).
-The reference noise-free flow map is in Figure 7(c), and the rasterised flow map in Figure 7(d).
-Morgan and Lovelace (2021) do not recommend a value for tuning the rasterisation grid resolution, so we select $\varepsilon_{R}=2$ m subjectively so that the raster grid cells do not exceed the road widths.
-Refer to caption Refer to caption
-Refer to caption Refer to caption
-Figure 7: Synthetic Culver City trajectories $(n=100)$.
-(a) Noisy trajectories (green lines).
-(b) Zoom of noisy trajectories near intersection of Main St and Venice Bd (solid yellow line).
-(c) Reference noise-free flow map.
-(d) Rasterised flow map with $\varepsilon_{R}=2$ m.
-Label in flow maps is traffic flow.
-For our locally aligned flow maps, the free tuning parameter is the blend tolerance parameter $\varepsilon$.
-So we investigate the effect of $\varepsilon$ on the accuracy of flow aggregation.
-For the blend tolerance $\varepsilon=1$ m, the flow map in Figure 8(a) is unable to consistently blend the map matched routes correctly since 1 m is typically smaller than the remaining discrepancies after map matching in ST_ROUTE.
-For the blend tolerance $\varepsilon=5$ m, the central road segments labelled with flow 51 and 45 in Figure 8(b), which are close to 49 and 45 in the reference flow map.
-For the blend tolerance $\varepsilon=10$ m, the central road segments are also labelled with flow 50 and 45 in Figure 8(c), similar to with blend tolerance $\varepsilon=5$ m in Figure 8(b).
-Near the intersection of Venice Bd and Main St (solid yellow line), the flow segment (with flow 17) is poorly aligned to the road network, unlike the more accurate alignment in Figure 8(b).
-Refer to caption Refer to caption
-Figure 8: Effect of blend tolerance $\varepsilon$ on synthetic Culver City flow maps.
-(a) Flow map with $\varepsilon=1$ m.
-(b) Flow map with $\varepsilon=5$ m.
-(c) Flow map with $\varepsilon=10$ m.
-Label is traffic flow.
-To complement the above heuristic comparisons, we define a quantitative discrepancy between a flow map $\mathcal{F}$ and the reference flow map $\mathcal{F}_{\rm ref}$.
-First, we compute the transect points for each linestring ${\bm{f}}$ from $\mathcal{F}$: if $\operatorname{len}({\bm{f}})>\varepsilon_{T}$ m, then a point is placed at every $\varepsilon_{T}$ m of ${\bm{f}}$; or if $\operatorname{len}({\bm{f}})\leq\varepsilon_{T}$ m, then a single point is placed at the centroid of ${\bm{f}}$.
-So $\mathcal{T}=\{{\bm{t}}_{1},\dots,{\bm{t}}_{n_{\mathcal{T}}}\}$ is a set of $n_{\mathcal{T}}$ transect points.
-For each ${\bm{t}}_{i},i=1,\dots,n_{\mathcal{T}}$, we find the closest linestring in $\mathcal{F}_{\rm ref}$, that is, ${\bm{f}}_{{\rm ref},i}^{*}=\operatorname{argmin}_{{\bm{f}}_{\rm ref}\in\mathcal{F}_{\rm ref}}\mathrm{dist}({\bm{t}}_{i},{\bm{f}}_{\rm ref})$.
-The flow errors are $\operatorname{Err}_{i}=|f_{i}-f_{{\rm ref},i}^{*}|$ for the transect points on ${\bm{t}}_{i}\in\mathcal{T}\subseteq\mathcal{F}$.
-We reverse the order of the comparison to obtain the flow errors $\operatorname{Err}_{j}=|f_{{\rm ref},j}-f_{j}^{*}|$ for the transect points on ${\bm{t}}_{j}\in\mathcal{T}_{\rm ref}\subseteq\mathcal{F}_{\rm ref}$, and concatenate these two sets of flow errors.
-Refer to caption Refer to caption
-Refer to caption Refer to caption
-Figure 9: Flow errors for synthetic Culver City flow maps.
-(a) Flow map with $\varepsilon=1$ m.
-(b) Flow map with $\varepsilon=5$ m.
-(c) Flow map with $\varepsilon=10$ m.
-(d) Rasterised flow map with $\varepsilon_{R}=2$ m.
-Flow segments are coloured lines, transect points are grey circles, flow errors are labels.
-In Figure 9(a), we display the transect points with $\varepsilon_{T}=50$ m, as the grey circles, superposed over the flow map $\mathcal{F}$.
-Figures 9(a–c) are the flow errors for the locally aligned flow maps with blend tolerance $\varepsilon=1,5,10$ m, and Figure 9(d) is for the rasterised flow map.
-The $\varepsilon=1$ m flow map and the rasterised flow map lead to many high error values.
-Whereas for the other two locally aligned flow maps, the flow errors are similarly low.
-We repeat the flow map error calculations for $N=10$ re-samples of $n=100$ Culver City trajectories.
-The mean and SD of the flow errors for the locally aligned flow maps are $3.061\pm 0.565$ ($\varepsilon=1$ m), $0.446\pm 0.047$ ($\varepsilon=5$ m), $0.477\pm 0.048$ ($\varepsilon=10$ m), and for the rasterised flow map $1.514\pm 0.201$ ($\varepsilon_{R}=2$ m).
-Setting $\varepsilon=1$ m leads to under-blending since $\varepsilon$ is smaller than the misalignments in the map matched routes, and setting $\varepsilon=10$ m leads to over-blending since $\varepsilon$ is larger than the distance between separate road segments.
-The $\varepsilon_{R}=2$ m rasterised flow map is similarly under-blended like the $\varepsilon=1$ m flow map.
-So we tentatively recommend a value of about $\varepsilon=5$ m for the blend tolerance parameter since it leads to optimal line blending.
-3.2 Empirical trajectories
-We return to the empirical Hannover trajectories.
-From the 1183 trajectories, we keep 1177 trajectories with length greater than $100$ m, and follow the workflow in Figure 6.
-We map match these 1177 trajectories with ST_ROUTE, and we continue with 1152 of them which have a sufficiently high quality match, where the Hausdorff distance $\operatorname{dist}_{\operatorname{Haus}}(R(G),G)<100$ m, and the ratio $\operatorname{len}(R(G))/\operatorname{len}(G)<1.1$, for a trajectory $G$ and a map matched route $R(G)$.
-That is, we discard 25 trajectories or a 2.02% failure rate for the Valhalla map matching, which is consistent with previous studies (Saki and Hagen, 2022).
-After some trial and error, we set the blend tolerance of $\varepsilon=4$ m, which is 1 m less than the 5 m tentatively recommended from the synthetic trajectories computations above.
-The workflow leads to a locally aligned flow map displayed in Figure 10(a).
-The corresponding rasterised flow map is in Figure 10(b).
-At the city level, both of these flow maps visualise well the traffic flows.
-Though if we zoom into Klusreide (solid yellow circle), then we observe that in Figure 10(c) the locally aligned map displays the aggregated flows on linear road segments in comparison to the pixelated road segments of the rasterised flow map in Figure 10(d).
-Whilst we do not have a reference flow map for these empirical trajectories, like we had for the synthetic Los Angeles trajectories, we are still able to compute a proxy flows based on empirical counts of the map matched routes with orthogonal line transects from the flow map road segments.
-Along Klusriede, the locally aligned flow map is composed to two road segments (in pink) with 0 and 2 flow errors) corresponding to the proxy flows 312 and 315.
-For the rasterised flow map, some grid cells in the pink region have flow error 2.
-Though rasterisation some times leads to double counting and thus to much larger flow errors, e.g.
-in the orange region near the T-junction, the estimated flow is 643 whereas the proxy flow is 315, so the flow error is 328.
-The locally aligned map does not suffer from these rasterisation effects, so its flow aggregation is more accurate.
-Refer to caption Refer to caption
-Refer to caption Refer to caption
-Figure 10: Hannover flow maps.
-(a) Locally aligned flow map at city level.
-(b) Rasterised flow map at city level.
-Solid yellow circle is Klusriede.
-(c) Locally aligned flow map at road segment level.
-(d) Rasterised flow map at road segment level.
-Colour scale is traffic flow, label is traffic flow error.
-Our future goal is the integration of the local alignment algorithms into an industry standard web GIS.
-In the mean time, we make available interactive web maps at https://mvstat.net/traj which allow the user to freely explore our proposed multi-scale traffic flow maps from the individual road to the city-wide level.
-3.3 Computation
-We report the execution times for the flow map computations for the synthetic and empirical trajectories.
-Since it is difficult to reliably reproduce execution times on different systems, we provide our execution times as a guide only.
-These are obtained on a stand-alone set-up running Ubuntu 24.04 with 16Gb RAM and 4 × Intel i5 CPU @ 3.10 GHz.
-We deploy a local dockerised Valhalla routing engine API (https://github.com/gis-ops/docker-valhalla) (GIS OPS, 2023) rather than relying on a web-based service.
-Since the local map matching API requests can be efficiently parallelised on a stand-alone machine, we are confident that ST_ROUTE is suitably optimised.
-For $n=100$ synthetic Culver City trajectories, ST_ROUTE executes in 188 s.
-For $n=1000$ synthetic trajectories, it executes in 2233 s, which is 10 times longer for a 10-fold increase in $n$.
-So it is straightforward to infer a precise execution time of ST_ROUTE for large sample sizes from a small sample size.
-The execution time for ST_OVERLINE increases as $n$ increases, but not in the same uniform manner as ST_ROUTE as the former also depends on the value of $\varepsilon$.
-For $n=100$, the execution times for ST_OVERLINE are 1777 s ($\varepsilon=1$ m), 1049 s ($\varepsilon=5$ m), 880 s ($\varepsilon=10$ m); and for $n=1000$, they are 15 457 s ($\varepsilon=1$ m), 6002 s ($\varepsilon=5$ m), 3897 s ($\varepsilon=10$ m).
-So it is difficult to infer precisely the execution times for ST_OVERLINE for larger sample sizes and smaller blend tolerances from smaller $n$ and larger $\varepsilon$.
-This is because the line blending (ST_LINEBLEND) is parallelised, but the line blending priority (ST_LINEBLEND_PRIORITY) remains a serial computation.
-With an optimal value of the blend tolerance $\varepsilon=5$ m, the total execution time of the workflow for the $n=100$ Culver City trajectories is 1238 s or $\sim$21 mins, and for $n=1000$ is 8236 s or $\sim$137 mins.
-The total execution time for the $n=1183$ Hannover trajectories with $\varepsilon=4$ m is 5545 s or $\sim$92 mins.
-Thus locally aligned flow maps are currently suitable for offline analyses of small and medium sized data sets.
-Due to time and computing constraints, (a) data sets with a larger number of trajectories (like the full set of 1.5 million synthetic Los Angeles trajectories) and (b) online computations were out of scope.
-For these to be computationally feasible, we require additional implementation improvements in the local alignment algorithms, which we leave for future development.
-4 Conclusion
-We have introduced an iterative workflow which takes noisy GNSS vehicle trajectories as input and gives an accurate traffic flow map as output.
-Local alignment of flow segments is the key innovation.
-To carry out this local alignment, we defined a spatial relation to set up local reference flow segments, which then allowed us to align other nearby flow segments to this local reference segment.
-We presented solid evidence for the high level of spatial resolution and accuracy for our proposed locally aligned flow map for both synthetic and empirical trajectories.
-We demonstrated that they effectively reduce the noise of the input trajectories, and accurately aggregate the traffic flows on all road segments, and so they are a valuable addition to the toolkit for multi-scale traffic flow analysis.
+Data availability statement
+The data that support the findings of this manuscript are openly available.
+The Hannover trajectories are hosted in the Leibniz University data depository https://doi.org/10.25835/9bidqxvl, and the Los Angeles trajectories in the Dryad depository https://doi.org/10.5061/dryad.4j0zpc8gf.
+Disclosure statement
+The authors report there are no competing interests to declare.